### v1

# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

Abstract
Migration has been a core topic in German political debate, from millions of expellees post World War II over labor migration to refugee movements in the recent past. Studying political speech regarding such wide-ranging phenomena in depth traditionally required extensive manual annotations, limiting the scope of analysis to small subsets of the data. Large language models (LLMs) have the potential to partially automate even complex annotation tasks. We provide an extensive evaluation of a multiple LLMs in annotating (anti-)solidarity subtypes in German parliamentary debates compared to a large set of thousands of human reference annotations (gathered over a year). We evaluate the influence of model size, prompting differences, fine-tuning, historical versus contemporary data; and we investigate systematic errors. Beyond methodological evaluation, we also interpret the resulting annotations from a social science lense, gaining deeper insight into (anti-)solidarity trends towards migrants in the German post-World War II period and recent past. Our data reveals a high degree of migrant-directed solidarity in the postwar period, as well as a strong trend towards anti-solidarity in the German parliament since 2015, motivating further research. These findings highlight the promise of LLMs for political text analysis and the importance of migration debates in Germany, where demographic decline and labor shortages coexist with rising polarization.
keywords:
solidarity, large language models, annotation, political discourse
$\cup@datesfalse$
Aida Kostikova]aida.kostikova@uni-bielefeld.de
1 Introduction
Migration is a defining feature of the 21st century, explicitly recognized in the 2030 Agenda for Sustainable Development wilkinson2022climate (105, 1). As the worldbank2023 (106) stresses, global economic imbalances, demographic divergence, and climate change make migration increasingly unavoidable, especially from the Global South to the Global North almulhim2024climate (3, 55, 69). However, if governed responsibly, migration can also drive prosperity and contribute to achieving the Sustainable Development Goals worldbank2023 (106, 38, 1).
Within Europe, Germany occupies a particularly important position as a major migration destination eurostat2024migration (32). As the EU’s largest economy, it offers labor market opportunities, institutional support, and stability bertoli2016crisis (10). At the same time, demographic decline and skilled-labor shortages make sustained immigration economically beneficial smith2025demographic (27, 29, 4). Migration has also profoundly shaped Germany’s political development, from the resettlement of millions of expellees after WWII, to the labor migration of the 1950s–1970s, to refugee movements in the past decade kaya2017inclusion (52, 36). This raises important questions: how does German politics approach migration, and to what extent does it express solidarity – or anti-solidarity – toward migrants?
Answering such questions requires systematic analysis of political discourse across long time spans, traditionally relying on manual textual analysis and qualitative interpretation, which remain powerful tools randour2020twenty (78). However, such methods are labor-intensive bhattacherjee2012social (11) and thus often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized babalola2021current (7). Computational social science (CSS) emerged to address the scale and complexity of modern data, yet it often struggles with theoretical alignment and interpretive depth lazer2024future (62). Large language models (LLMs) promise to bridge this gap by enabling automated, in-depth analysis of large-scale text. While LLMs have shown strong performance on tasks like ideological stance detection ziems2024can (109), it remains an open question whether they can handle more complex, theory-driven sociopolitical annotation tasks. We provide a partial answer: If accompanied by substantial validation and careful interpretation, LLMs can approximate the performance of trained human annotators and generate annotations that help quantify large-scale trends of complex constructs. In particular, we apply LLMs to annotate solidarity and anti-solidarity towards women and migrants in German parliamentary speeches, compare LLM annotations to thousands of human reference annotations, and interpret the (anti-)solidarity trends between 1949-1969 and 2009-2025 in light of German political history.
Figure 1: Overview of the annotation scheme adapted from thijssen2012mechanical (97). Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none. Solidarity and anti-solidarity instances are then further categorized into subtypes based on the underlying rationale: group-based, exchange-based, compassionate, and empathic. The example sentence in the figure illustrates exchange-based anti-solidarity.
We focus on solidarity as a particularly demanding test case for such an approach. While frequently invoked in political discourse, solidarity has often been treated as a symbolic notion rather than an operational concept kozlowska2024solidarity (58), which has limited large-scale empirical research. To address this gap, we approach solidarity as a socially meaningful category that can be systematically studied in political text. For this purpose, we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97), which provides a structured and annotatable framework. It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure 1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification). Germany offers a particularly suitable context for such a study: solidarity is deeply rooted in its political vocabulary fiegle2003solidarite (34, 94, 84), and its multi-party system enables us to trace how meanings of (anti-)solidarity vary across parties and over time.
In a prior conference paper, we already performed a first analysis of LLM annotations of (anti-)solidarity towards women and migrants on this data set kostikova-fine (57). In this paper, we 1) evaluate a much wider range of LLMs, 2) systematically evaluate the effect of prompts, model size, and historic (versus contemporary) data, 3) include a much larger data set of human annotations, 4) validate robustness by comparing the time series of model-predicted (anti-)solidarity shares from the selected model (Llama-3.3-70B) and an independent model (GPT-4), and 5) most importantly, we interpret the (anti-)solidarity trends towards migrants over time to derive novel insights for political science. In particular, we find a) that different groups of migrants are associated with different types of (anti-)solidarity in German parliamentary debates during the post-World War II period; and b) that anti-solidarity towards migrants has dramatically increased in the German parliament over the past 10 years, indicating a shift to the right (especially in the liberal and conservative parties) on this issue that motivates further research.
1.1 Related Work
Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; and (ii) the use of LLMs as annotation tools in political and social science contexts. While prior studies have engaged with each of these dimensions individually, few have combined them to enable large-scale, longitudinal analysis of solidarity in political speech. We briefly review these literatures below.
Solidarity in Political Discourse and NLP
Sociological theories of solidarity often begin with Durkheim’s classical distinction between mechanical and organic forms, which describe different bases for social cohesion. While influential, this binary model has been criticized for its structuralism and linear view of social change. thijssen2012mechanical (97) expands on Durkheim by integrating recognition-based theories such as those developed by honneth1996struggle (45, 46) and proposing a fourfold typology with group-based, compassionate, exchange-based, and empathic subtypes, which can be applied to both historical and contemporary political discourse. thijssen2022s (98) apply this typology to Belgian party manifestos, showing how different parties emphasize different solidarity frames in line with broader ideological divides. We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs. While some studies have examined solidarity longitudinally using interviews or survey data schonweitz2024solidarity (89, 92), these approaches are not designed for large-scale text analysis or fine-grained typological classification.
At the same time, solidarity has received relatively little attention in computational text analysis, where more application-driven concerns such as hate speech and misinformation detection have dominated the field. Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements weber2024social (103, 28), the COVID-19 pandemic kalaloi2021mediated (51), and natural disasters santhanam2019stand (88). These works typically use supervised models to detect solidarity or anti-solidarity in large-scale Twitter data and track how such expressions evolve over short timeframes. While they demonstrate the feasibility of automated solidarity detection, they focus on specific, time-bound crises and treat solidarity as a binary sentiment-like category rather than a differentiated moral frame. In contrast, our work applies a theory-driven typology to political speech, enabling fine-grained analysis of solidarity in parliamentary discourse.
LLMs as Annotation Tools in Political and Social Science.
LLMs are increasingly used in political and social science research. Most applications focus on standard classification tasks such as political ideology detection (heseltine2024large, 43, e.g.); sentiment analysis lashitew2025corporate (61, 37); stance detection (cao2024can, 19, e.g.); and misinformation detection (huang2025unmasking, 47, e.g.). For broader surveys of LLM applications, benefits, and challenges in social and political science, see thapa2025large (95, 66, 109). However, conclusions about LLM performance in these fields vary. Some studies report that LLMs show near-human performance rytting2023towards (87) or even outperform humans, as well as traditional models, on short texts and tasks like ideology classification le2025positioning (63, 43). In contrast, several studies show that smaller fine-tuned models like BERT still outperform LLMs, especially when labeled data is available wang2024selecting (102, 18) or when tasks require expert-defined taxonomies and contextual reasoning ziems2024can (109). Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories. For example, erhard2025popbert (30) develop PopBERT, a fine-tuned transformer trained to detect populist rhetoric and ideological dimensions in German Bundestag speeches using a custom annotation scheme. card2022computational (20) analyze 140 years of U.S. political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, identify rhetorical frames, and detect dehumanizing metaphors, demonstrating the utility of fine-tuned LMs for large-scale, theory-aligned analysis. Related work has applied fine-tuned models to pragmatic and rhetorical categories in German parliamentary debates, such as speaker attribution, speech acts, and opinion role labelling reinig2024politics (83, 81, 80). In contrast to fine-tuned approaches, ranjit2024oath (79) use general-purpose LLMs (e.g., GPT-4, Flan-T5) to apply a custom typology of attitudes toward homelessness, developed from framing theory and grounded theory. Their evaluation focuses on aggregate trends and correlations with external structural indicators, using LLMs to annotate a large corpus of tweets at scale and over time.
Our work differs in several important ways. Unlike erhard2025popbert (30) and card2022computational (20), we use state-of-the-art LLMs and avoid the need for supervised data sets and fine-tuning. Unlike ranjit2024oath (79), who also use prompted LLMs, we adopt a typology grounded in established sociological theory rather than one developed inductively for a specific domain. We apply this typology to more formal political speech rather than short, informal social media texts, and evaluate model performance through systematic comparison with human annotations. Finally, our analysis spans over 150 years of parliamentary debate, encompassing two case studies, historical and contemporary, and enabling fine-grained, longitudinal insight into political discourse.
2 Methodology
2.1 Data
We build on the DeuParl dataset introduced in DeuParl (101), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022. We extend this corpus to include protocols up to June 2025. We follow the preprocessing steps from DeuParl (101) and include metadata such as date, session number, and legislative period.
We reuse the keyword sets (see a full list of keywords in A) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in kostikova-fine (57), consisting of 32 and 18 terms respectively. Based on these keywords, we extract 63k migrant-related and 138k women-related sentences (instances) from DeuParl, additionally including three sentences before and after each instance for context. We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) provided by poradaautomatische (77). See B for details. Fig. 2 shows the number of instances over time, as well as their share relative to all sentences in the DeuParl dataset. Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015, coinciding with refugee crisis debates. Notably, migrant mentions also surpass women-related discourse in the immediate post-war years.
Figure 2: Absolute number of instances (left) and their share relative to all DeuParl sentences (right) for the Woman and Migrant categories.
2.2 Data Annotation
To validate our models and enable further qualitative analysis, we build on the human-annotated dataset introduced in kostikova-fine (57), which comprises 2,864 instances (1,437 for migrants, 1,427 for women) from the DeuParl corpus. [^1] Annotation was carried out by five trained student assistants through an iterative process, achieving Cohen’s $\kappa$ of 0.62 (high level) and 0.42 (fine-grained). Final labels in the dataset are assigned by majority vote. Each instance is labeled according to a fine-grained solidarity framing scheme that distinguishes four high-level stance categories – solidarity (willingness to share resources, following Lahusen.2018 (60) and TwitterDataset (50)), anti-solidarity (restriction or exclusion of groups), mixed (both supportive and opposing elements), and none (neutral or unrelated content). Instances labeled as solidarity or anti-solidarity are further assigned one of four frame types adapted from thijssen2022s (98): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance in contributions), and empathic (respect for or rejection of social diversity).
We extend the evaluation to a newly annotated set of 756 instances for Migrant with more historical coverage (before 1950). For this set, annotations were conducted by 4 annotators following the same procedure as in kostikova-fine (57) Statistics on the human-annotated dataset are presented in Fig. 3. Overall, the label distribution differs notably by target group: woman-related instances are predominantly labeled as solidarity (74.1%), while migrant-related instances show a more diverse pattern, with lower solidarity (51.6%) and a higher share of anti-solidarity (21.3%).
Figure 3: Distribution of instances in the human annotated dataset across time and target groups. Total indicates number of all annotated samples for the given target group, including the None label. See Table 2 for the actual numbers of instances in the human annotated dataset.
2.3 Models and Experiments
We evaluate a diverse suite of models, ranging from smaller supervised models (BERT, fine-tuned SBERT) to proprietary LLMs (GPT-4) and recent open-weight instruction-tuned models. The final list includes:
  1. (i)
    smaller pre-trained LMs like BERT and SBERT;
  2. (ii)
    closed-source LLMs like gpt-4-1106-preview;
  3. (iii)
    open-weight LLMs such as Llama-3-8B-It, Llama-3.3-70B-It (and fine-tuned Llama-3-70B-It), Gemma-2-9b-It, Gemma-2-27b-It, and Mistral-2-Large.
Data and Metrics
We adopt the 70/15/15 train/dev/test split strategy as in kostikova-fine (57), generating three random splits of approximately 2,000 training, 400 development, and 430 test instances for each target group. These are used to train and evaluate supervised models. For inference-only experiments, we use only the test sets: Test 1 (original) and Test 2 (newly added, 756 instances), the latter offering broader historical coverage (1860s–1950s). We report Macro F1 to account for class imbalance, along with per-class F1 scores for both high-level and fine-grained classification tasks. [^2] High-level classification refers to the four stance categories: solidarity, anti-solidarity, mixed, and none. Fine-grained classification includes ten labels: the four subtypes of solidarity, the four subtypes of anti-solidarity, plus mixed and none.
Models
We use manually designed prompts for each target group. Each prompt consists of a classification instruction concatenated with the parliamentary instance, i.e., the keyword sentence plus its three-sentence left/right context as defined in 2.1. Note that, although speaker and party metadata are available in the corpus, this information is removed from the model input. The classification instruction includes (i) chain-of-thought reasoning in zero- and few-shot settings (via think step by step); (ii) label definitions; and (iii) a two-step classification format: high-level first, followed by subtype for solidarity or anti-solidarity prediction. [^3] For GPT-4, both steps are combined in a single prompt (see Figure 21 for Migrant and Figure 22 in the Appendix for Woman). For open-weight models, we instead use two prompts: one for high-level classification and, conditional on that output, a second for subtype selection (see Figure 24 for Migrant and Figure 27 in the Appendix for Woman). In few-shot mode, prompts include one representative example per label.
We apply these prompts to closed-source models like gpt-4-1106-preview, and to open-weight LLMs including Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it (full precision) in inference-only setting. Larger models (Llama-3.3-70B-Instruct and Mistral-Large-Instruct-2407) are run with $Q4_{0}$ quantization. Additionally, we fine-tune Llama-3-8B-It and Llama-3-70B-It using a three-model pipeline (high-level and two subtype classifiers), applying instruction-style fine-tuning with true labels paired with GPT-4-generated explanations as training targets.
We use BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details for all models used in our experiments are provided in C.
3 Results
In this section, we evaluate models against a human-annotated dataset to determine reliability for social scientific analysis of solidarity frames in parliamentary texts. Our aim is twofold: (i) to identify a model whose predictions approximate human-level annotation quality across both high-level and fine-grained stance categories, and (ii) to ensure the selected model is practically feasible to deploy at scale. Specifically, we ask whether recent open-weight models can match the performance of GPT-4, which was previously found to be highly reliable (though costly) kostikova-fine (57), through techniques such as few-shot prompting or lightweight fine-tuning.
3.1 Model Performance Overview
Table 1 reports macro F1 scores for each model across both inference-based and fine-tuned approaches. We benchmark these results against a human upper bound, estimated using a leave-one-out (LOO) procedure, [^4] For each annotator, we exclude their labels, compute a consensus label from the remaining annotators, and calculate the macro F1 score between the excluded annotator and the consensus. We then average these scores across all annotators. and report performance for both high-level and fine-grained classification tasks, with the latter shown in parentheses. Our observations are structured around two dimensions: (i) overall model performance, (ii) task and label-related challenges.
Model Method Woman Migrant (T1) Migrant (T2)
GPT-4 Zero-shot 0.37 (0.60) 0.42 (0.73) 0.29 (0.57)
Few-shot 0.37 (0.54) 0.43 (0.63) 0.29 (0.51)
Mistral-2-Large Zero-shot 0.20 (0.50) 0.27 (0.56) 0.24 (0.45)
Few-shot 0.26 (0.51) 0.38 (0.57) 0.26 (0.44)
Llama-3.3-70B Zero-shot 0.23 (0.50) 0.32 (0.54) 0.25 (0.47)
Few-shot 0.30 (0.50) 0.37 (0.65) 0.28 (0.53)
Llama-3-70B Fine-tuned 0.26 (0.49) 0.27 (0.65) 0.21 (0.47)
Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.17 (0.40)
Few-shot 0.14 (0.35) 0.19 (0.34) 0.15 (0.40)
Fine-tuned 0.16 (0.38) 0.20 (0.40) 0.13 (0.27)
Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.25 (0.45)
Few-shot 0.25 (0.55) 0.33 (0.48) 0.26 (0.46)
Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.22 (0.46)
Few-shot 0.23 (0.44) 0.31 (0.54) 0.24 (0.46)
BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.23 (0.45)
SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.18 (0.36)
Human upper bound 0.34 (0.62) 0.48 (0.75) 0.35 (0.57)
Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks. ?Migrant (T1)? uses the original benchmark; ?Migrant (T2)? is newly introduced in this paper. Green cells indicate the best-performing score for each target group, based on a model’s best setting (e.g., zero-shot or few-shot); yellow cells indicate the best-performing score from a different model (i.e., the second-best model overall).
  1. (i)
    Overall model performance.
    • GPT-4 achieves the highest scores, but Llama-3.3-70B is a strong alternative. Across all target groups (Woman and both Migrant test sets), GPT-4 (zero-shot) consistently shows the top performance, coming closest to the human upper bound in both high-level and fine-grained classification. Llama-3.3-70B is the best-performing open model, showing meaningful gains from few-shot prompting (in particular in the high-level classification).
    • Model size correlates with performance, but not perfectly. Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure 4 for the relationship between model size and performance). Mid-sized models follow, and smaller models perform worst. However, this relationship is not strictly linear: for example, Llama3-8B-it underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B-it.
      Figure 4: Model size vs. performance on high-level and fine-grained classification tasks in Migrant (T1). Each model is plotted with its best-performing configuration (zero- or few-shot). The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
    • Examples help mid-sized models more than small or very large ones. Few-shot prompting improves performance primarily for mid-sized models such as Gemma-9B-it, Gemma-27B-it, Mistral-2-Large, and Llama-3.3-70B-it. In contrast, smaller models like Llama-3-8B-it show no benefit, and GPT-4 sees minimal gains—likely because its zero-shot performance is already near-optimal. Fine-tuned variants (e.g., Llama-3.3-70B-
      it
      , Llama-3-8B-it) often show no improvement over inference-only models.
  2. (ii)
    Task and label-related challenges.
    • All models perform worse on the fine-grained task, underscoring its complexity: unlike the more intuitive high-level categories (e.g., solidarity, anti-solidarity), it requires alignment with annotation guidelines grounded in social theory, where even human annotators show limited agreement (reaching macro F1 score 0.34 for Woman and 0.48 for Migrant).
    • At the high level, solidarity is the easiest to detect (see Figure 5; full scores in Table 3 in the Appendix), especially in the Woman category, where even smaller models like BERT (0.77) and Gemma-9B (0.78) perform strongly. Anti-solidarity and mixed stance are the most challenging categories, with GPT-4 achieving the highest high-level anti-solidarity scores (0.64 for Migrant, 0.65 for Woman), but mid-size models (e.g., Mistral, Llama-3.3-70B, Gemma-27B, Gemma-9B) show moderate performance (with scores ranging from 0.52 to 0.54). Only GPT-4 and Llama-3.3-70B reliably handle mixed stances (e.g., 0.38–0.43 for Migrant, 0.30–0.37 for Woman).
    • On the fine-grained level, GPT-4, Llama-3.3-70B, and Gemma models offer the most balanced coverage across subtypes. GPT-4 stands out on exchange-based solidarity (0.56), group- and exchange-based anti-solidarity for Migrant (0.47 and 0.46), and on empathic and compassionate anti-solidarity for Woman (0.67 and 0.50). Among open models, Llama-3.3-70B performs particularly well on compassionate solidarity (0.51), and shows strong results on group- and exchange-based anti-solidarity (0.42 and 0.35). Interestingly, the Gemma models are able to distinguish the rare compassionate anti-solidarity subtype for Migrant with some success. Most other models struggle with fine-grained anti-solidarity, with scores close to zero.
      Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For the Migrant group, the two test sets (Test 1 and Test 2) are combined.
    • Performance drops on historical data. All models perform worse on the historical Migrant (T2) set, including the strongest ones. This reflects a broader pattern: model performance tends to decline on older data, as shown by macro F1 scores over time (see Figure 16)—so does agreement among human annotators (Cohen’s Kappa). However, while human trends are smooth and monotonic, model trajectories are less stable. Stronger models such as GPT-4, Mistral-2-Large, Llama-3.3-70B, and BERT show a general upward trend but exhibit clear dips in the 1930s and 1960s. In contrast, smaller models (e.g., Llama-3-8B, SBERT, Gemma-9B) show little to no temporal coherence.
In summary, GPT-4 remains the strongest model overall. While performance generally scales with model size, mid-sized models can approximate GPT-4 when given examples. Among open-weight alternatives, Llama-3.3-70B emerges as the most viable option, particularly for the Migrant category, delivering competitive results in both high-level and fine-grained classification, while offering a more cost-effective solution. We therefore adopt it for the subsequent social scientific analyses. However, since it still lags behind GPT-4 in capturing certain fine-grained subtypes, we further validate its reliability by comparing the temporal patterns produced by both models in 4.1.
3.2 Error Analysis
In this subsection, we examine the errors made by the two strongest models in our experiments, GPT-4 and Llama-3.3-70B, as determined by macro F1 scores. To identify where these models struggle, we compare their confusion matrices on the full Migrant test set (Figure 6 for high-level; see Figure 17 in the Appendix for confusion matrices for all models included in the study and Figure 18 in the Appendix for fine-grained label confusion matrices) with a human reference, constructed by averaging confusion matrices from four leave-one-out (LOO) comparisons among annotators.
Figure 6: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four LOO comparisons, where each annotator is compared to the consensus of the others.
Both models exhibit notable differences from human annotators. In particular, both models frequently annotate true Solidarity instances as None (108 instances for GPT-4, 128 for Llama) and true Anti-solidarity instances as None (59 for GPT-4, 60 for Llama), especially for common subtypes such as group-based and compassionate. These errors are far less frequent in human annotations (32 and 12, respectively), suggesting that models may more often underestimate both solidarity and anti-solidarity. Llama additionally shows a tendency to misclassify true Solidarity as Anti-solidarity (30 cases vs. 11 for GPT-4 and 13 for human annotators) and, to a lesser extent, true Anti-solidarity as Solidarity (18 vs. 4 for GPT-4 and 10 for humans). Some of these misclassifications arise when a speaker cites an anti-solidarity statement by another party, which Llama occasionally misinterprets as anti-solidarity (see Examples 9 and 11 in Table 14 in the Appendix; these are discussed in more detail in Section 4.3. However, manual inspection indicates that such cases are rare.
4 Temporal Trends in Solidarity Towards Migrants
In this section, we apply the best-performing open-weight model identified in 3, Llama-3.3-70B, to the full Migrant dataset of 63k instances to analyze trends across time. We focus on the Migrant dataset in particular because migration has long been one of the most contested issues in German politics, consistently framed through ?crisis narratives? that have shaped national identity and migration governance across decades frohlich2025migration (36). This makes it a compelling case for evaluating whether model-based annotations align with historical and political science literature. It is also reflected in our human-annotated dataset, where migrant-related instances display a more diverse label distribution, including substantially more anti-solidarity, compared to the predominantly solidarity-oriented discourse on women. More specifically, 4.1 compares the long-term (anti-)solidarity trends across human annotators and LLMs and thereby further validates the results of our models. In doing so, we also note two periods with notable shifts in solidarity, namely the post-war period and the recent past, which we cover in more detail in Section 4.2 and Section 4.3.
(a) GPT-4
(b) Llama-3.3-70B
Figure 7: Solidarity trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Values are normalized within each decade; percentages may not total 100% because instances classified as ?None? are not shown (though included in the calculation). The period 1933–1949 (grey shaded area) is excluded due to limited data availability during and immediately after the Nazi dictatorship.
4.1 Trends with GPT-4 vs with Llama-3.3-70B
(a) GPT-4
(b) Llama-3.3-70B
Figure 8: (Anti-)solidarity frames trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes; thus, the curves in both panels together sum to 100% per decade. The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
Figures 7 and 8 show the development of (anti-)solidarity for GPT-4 and Llama-3.3-70B on a high and fine-grained level across 155 years. The GPT-4 reference data is taken from kostikova-fine (57), based on a subset of 18.3k migrant-related instances (from an original 58k-instance dataset; see 2.1) spanning 1867–2022. While Llama-3.3-70B is applied to the full dataset (expanded to 63k instances through 2025) for the main analysis, we restrict it to the same 18.3k-instance subset for this comparison.
To validate the annotations of the models, we first compare them to each other. Comparing the trends between GPT-4 and Llama-3.3-70B, we observe a high correlation of 0.96 on the high level and between 0.77 and 0.95 across the more frequent fine-grained subtypes (see Table 4 in the Appendix). This indicates that the models show a high degree of agreement in temporal trends. [^5] Additionally, Llama-3.3-70B shows near-identical trends between the 18.3k subset and the full 63k dataset (e.g., $\rho$ = 0.996 for solidarity and anti-solidarity, and $\sim$0.97–0.98 for subtypes).
For reference, Figure 3 shows the high-level trends across time for human annotations ($N=2,084$). We observe similar relative distributions of (anti-)solidarity on a high level. Direct trend correlations between Llama-3.3-70B and human annotations are also substantial ($\rho$ = 0.84 for anti-solidarity, $\rho$ = 0.76 for solidarity, $\rho$ = 0.47 for mixed). In terms of Cohen’s $\kappa$ (computed on the 2,084 human annotated instances), GPT-4 slightly outperforms Llama-3.3-70B ($\kappa$ = 0.57 vs. 0.50 at the high level; 0.38 vs. 0.36 for subtypes), but differences are modest. Overall, GPT-4 and Llama-3.3-70B appear to produce similar results that agree to a moderate level with human annotations. Given that it is an open weight model that we can run on our own computing hardware at low further cost, we utilize Llama-3.3-70B for the subsequent analysis.
4.1.1 Trend Stability Across Keywords.
Before we move on to interpreting the observed trends derived from Llama-3.3-70B predictions, we verify that they are not artifacts of our keyword-based data collection (see 2.1). As Figures 19 and 20 in the Appendix show, individual keywords vary considerably in their temporal distribution and in the balance of solidarity vs. anti-solidarity. To rule out the possibility that our long-term trends merely reflect such keyword effects, we perform two robustness tests.
First, we conduct a stability test by randomly sampling 200 subsets of at least 5 keywords each and comparing the resulting high-level trends using Pearson’s correlation coefficient. Across all 19,900 pairwise comparisons, the trends are highly consistent (average correlation of 0.92 for solidarity, 0.81 for anti-solidarity, and 0.94 for mixed; with 25% quantiles of 0.90, 0.70, and 0.92, respectively). [^6] For more detail on keywords and the stability test, see A. Second, we restrict the analysis to the two most frequent and temporally widespread keywords, Ausländer (en.: foreigners) and Flüchtlinge (en.: refugees), which occur in every decade (19.7k and 12k mentions, respectively). Together, the resulting plots (Figure 23 in the Appendix) and the stability test results confirm that overall trends in solidarity and anti-solidarity are preserved both under this restricted vocabulary and across different keyword selections.
4.1.2 Notable Trends in 1867-2022.
Inspecting the trends in Figure 7, we observe four periods with notable shifts in solidarity: Between 1890-1920, in the 1930s, in the post-war period, and in the recent past after 2010.
Regarding the 1890-1920 period, we observe some disagreement between the models: GPT-4 shows an increase in solidarity between 1900 and 1920, while Llama-3.3-70B does not. We also note that the agreement between human annotators and between models and annotators is lower for historic data, such that we do not consider this trend to be sufficiently reliable for further analysis.
The 1930s and 1940s data points are outliers, as the parliament was swiftly sidelined by the Nazi regime and the only 109 available samples from 1933 to 1949 mostly concern inter-state relations (mentioning, e.g., ?foreign creditors?) that fall into the ?None? category. Hence, we exclude these data points from further analysis, as well.
For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look. In the postwar years (Section 4.2), parliamentary discourse reveals high levels of solidarity and little anti-solidarity toward migrants, despite the severe economic and infrastructural strain of the time braun2014immigration (14). This raises our first guiding question: Why is there so much solidarity (and so little anti-solidarity), and what was it based on? In contrast, the recent period (Section 4.3) shows a clear rise in anti-solidarity and a decline in solidarity, beginning in the early 2010s. These shifts likely reflect the political impact of recent refugee movements and the growing influence of anti-immigration views, such as those promoted by the far-right party Alternative für Deutschland (AfD) (en.: Alternative for Germany), which has gained substantial electoral support in Germany since its founding in 2013. This leads to our second question: What are the drivers of anti-solidarity in the contemporary period?
4.2 Germany after the Second World War, 1949-1957
The Second World War set off the largest wave of forced migrations in European history, with tens of millions of people, both non-Germans and ethnic Germans, displaced or resettled across Central and Eastern Europe schulze2017commemoration (90). The populations displaced by the war were highly heterogeneous, including refugees, displaced persons (DPs), forced laborers, and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans. For those who settled in postwar Germany, integration posed significant political and societal challenges, shaped by regional disparities, economic hardship, and tensions with the local population chevalier2024forced (21).
This subsection focuses on the early postwar period (1949–1957), when the Federal Republic of Germany institutionalized solidarity as part of nation-building. This period is analytically important because solidarity was no longer merely a moral gesture, but a tool of governance, embedded in legal frameworks like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy kittel2020stiefkinder (54). Crucially, such measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups from the benefits provided by these laws. However, since solidarity in this period was neither static nor uniform, shaped by diverse migrant groups, competing political ideologies, and contested notions of deservingness ther1996integration (96), we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97) to guide our analysis. By distinguishing between identity-, need-, and contribution-based forms, this framework allows us to examine how solidarity was constructed and sustained amid institutional strain.
4.2.1 Overall solidarity distribution
Throughout the 1949–1957 period, solidarity dominated parliamentary discourse on postwar population movements. As shown in Figure 9, [^7] We use raw counts as they provide a more interpretable view of stance distribution during this period, given the overwhelming dominance of solidarity and the low frequency of categories like anti-solidarity. mentions of solidarity consistently outpaced both anti-solidarity and mixed statements, at times reaching ten times their frequency.
Figure 9: Distribution of (Anti-)solidarity mentions in Bundestag debates, 1949–1957 (raw counts). LAG refers to the Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG to the Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG to the Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
Mentions of solidarity closely track key legislative moments, particularly in the debates leading up to the passage of major laws. A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced. A second cluster from January to May 1952 reflects debates on the LAG, preceding its adoption in May (see an excerpt from such a debate in Example 1 in Table 10 in the Appendix). Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified rights and integration measures, see Example 2 for such a debate. Later peaks in September 1954 and June 1955 reflect continued debate over LAG implementation and budget issues. Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
Figure 10: Distribution of (Anti-)solidarity frame subtypes from 1949 to 1957 (raw counts). An inset plot shows the evolving composition of solidarity subtypes, normalized within each 6-month period. Values represent the proportional share of each subtype among all solidarity mentions during that period.
The prominence of solidarity in this period is unsurprising. Parliamentary debates used it not only to justify compensation, but also to navigate postwar fragmentation and competing claims to victimhood and belonging hughes1999shouldering (49). This multifaceted nature of solidarity is evident in the distribution of all four subtypes across key legislative moments, as shown in Figure 10. Group-based and compassionate frames were especially prominent, reflecting an emphasis on both shared national belonging and moral obligation toward displaced populations (see, e.g., Example 1 in Table 10, where solidarity is tied to the right to the homeland and framed as a legal entitlement). More striking is the shift in the balance between these two subtypes over time. As shown in the inset of Figure 10, group-based solidarity dominated in the early 1950s but declined after mid-1953, while compassionate steadily increased. This suggests a rhetorical transition: once legal belonging was formalized, most notably with the LAG in 1952 and BVFG in 1953, solidarity no longer needed to be framed primarily in terms of shared identity. The shift may also reflect demographic change within the displaced populations, as newer arrivals like Sowjetzonenflüchtlinge from the GDR and Aussiedler (en.: resettlers) from the Eastern Bloc countries were less easily framed as co-nationals boke1996fluchtlinge (13). This trend aligns with Thijssen’s reading of Durkheim: as group-based solidarity extends to more diverse groups, it gives way to need-based appeals thijssen2012mechanical (97) (see Example 3 in Table 10 in the Appendix).
These patterns show that solidarity was used in parliamentary debates to manage the challenges of postwar reconstruction, as reflected in key legislative reforms. While the claimant groups of the broadest measure LAG were highly diverse (bombed out, evacuated, expelled, financially damaged), their eligibility for compensation under the law was based on Volkszugehörigkeit (en.: German ethnicity), excluding many from formal recognition kittel2020stiefkinder (54, 73). The concept of Volksdeutsche was a fluid and constructed Nazi category whose ambiguous definition nonetheless had concrete and deadly implications bergen2005tenuousness (9). We therefore examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in parliamentary discourse in that period.
4.2.2 Solidarity Frames Across Keywords in 1949-1957
This analysis includes only migrant-related keywords that appear at least 200 times in combination with a solidarity, anti-solidarity subtype or mixed stance. Keywords dominated by none or with low total counts are excluded (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants), etc.); see Table 5 for full distributions.
Figure 11 shows the distribution of (anti-)solidarity frames across migrant-related keywords in 1949–1957. Across all groups, solidarity was far more common than anti-solidarity, with group-based and compassionate frames dominating. However, the distribution of subtypes varies between groups.
Vertriebene (en.: expellees) are most often associated with group-based solidarity, accounting for around half of all expressions with a stance, while also receiving a substantial share of compassionate solidarity (nearly 30%). This reflects their dual position: as key representatives of German victimhood (alongside prisoners of war) during and after the World War II moeller2001war (71), they were both integrated into West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice. Still, around 5–6% of stance expressions toward Heimatvertriebene fall into the anti-solidarity category, particularly group-based (solidarity limited to ?us?, the in-group). These expressed concerns that expellees strained local infrastructures and were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment kossert2008kalte (56). Perceptions of injustice were heightened by the uneven distribution of expellees across West German states braun2014employment (15), particularly in regions where their prominent presence in public institutions and politics sparked fears of Überfremdung (en.: overforeignization) and contributed to a sense of unequal burden-sharing levy2005memories (64) see Example 4 in Table 10.
Figure 11: Relative distribution of (Anti-)Solidarity subtypes across selected migrant-related keywords, 1949–1957. Bars represent the proportion of each subtype within each keyword group, thus all subtypes for a given group sum to 100%.
Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are primarily framed through compassionate solidarity (45–52%), followed by group-based solidarity (25–30%). The label Flüchtlinge initially functioned as a broad catch-all category in the late 1940s nachum2018semantics (74). With the 1953 BVFG, it was more clearly differentiated into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler. Yet their Germanness was not valued equally: while Vertriebene were framed as victims of expulsion and central to West German nation-building, the label Flüchtlinge retained post-war connotations of marginality and foreignness nachum2018semantics (74). This helps explain a higher share of compassionate solidarity, as well as the notable anti-solidarity (around 10%). Sowjetzonenflüchtlinge, in turn, were primarily framed as Cold War refugees fleeing communist oppression and thus not integrated into the nation building narrative like the Vertriebene, prompting more compassionate than group-based solidarity nachum2018semantics (74, 13) (see Example 5).
Ausländer (en.: foreigners) were the main targets of anti-solidarity. Not a single coherent group, they were a discursively constructed ?other?, including DPs, Eastern European border crossers, wartime laborers, and early foreign workers smith1994germany (93). Often contrasted with German groups like Vertriebene or Einheimische, they were framed as non-belonging, economically burdensome, and sometimes criminalized holian1945missing (44) (see Example 6). However, despite this exclusionary framing, solidarity with Ausländer, especially group-based, exceeded anti-solidarity. This might be partly due to the category’s fluidity, which included ethnic Germans not yet formally recognized as such ruhkopf2023institutionalisierte (86) in the early postwar years. Politicians like Lukaschek also linked DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community ruhkopf2023institutionalisierte (86) (Example 7). The second-highest solidarity subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy within the Western alliance ruhkopf2023institutionalisierte (86) (Example 8). The 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure, granted DPs civic and social rights largely to meet conditions for sovereignty and improve the country’s international standing holian1945missing (44, 86). At the same time, instances of genuine solidarity toward non-ethnic German foreigners, particularly in CDU and SPD, statements, show that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 7 and 9).
In sum, solidarity was expressed differently depending on the migrant group and was shaped by perceived ?deservingness?. Vertriebene were largely included as part of the national community, while others, especially Ausländer, were treated as outsiders and received comparatively less support. These differences reflected not only group characteristics but also the ideological strategies of political parties. The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.
4.2.3 Distribution across parties
Figure 12 shows the distribution of solidarity and anti-solidarity subtypes across parties for the period 1949–1957. While solidarity was the dominant stance overall, parties varied in how they framed it, driven by ideological and strategic choices. Some parties used solidarity to promote national integration, while others employed it to defend regional concerns. Broadly, we divide parties into three categories: mainstream governing parties, expellee advocacy parties, and smaller parties with ideological or regional agendas.
The main parties – governing CDU/CSU and FDP, and opposition SPD – supported the integration of expellees and refugees through legal and institutional measures, primarily expressing group-based or compassionate solidarity (see Figure 12). The CDU/CSU played a leading role in proposing the LAG, which framed solidarity as a national obligation toward fellow Germans, and even included expellee representatives in the party ranks hughes1999shouldering (49). The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers connor2018refugees (22), resulting in occasional anti-solidarity statements. The FDP, while formally aligned with CDU/CSU and SPD on key refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation connor2018refugees (22) (see Example 10 in Table 10, where FDP speakers argued that expellees should build their own homes through self-help and cooperative effort, with state support limited to loans and low-interest credit).
Expellee advocacy parties (GB/BHE and DP) [^8] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights) and Deutsche Partei (DP) (en.: German Party) explicitly positioned themselves as political representatives of displaced Germans. Following multiple successes in state elections, GB/BHE entered the federal government coalition in 1953, thereby exerting significant pressure on other parties to address the demands and concerns of expellees connor2018refugees (22). They framed solidarity primarily in terms of historical injustice and economic entitlement, calling for wealth redistribution, fair pensions, and affordable housing. Their rhetoric relied heavily on group-based solidarity, portraying expellees as full members of the national community who deserved compensation and full integration (see Example 11). Eventually, both major parties of the Federal Republic CDU and SPD managed to absorb the demands and concerns of the expellees, which made GB/BHE and DP less successful connor2018refugees (22, 54).
Regionalist parties such as the WAV and BP [^9] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized group-based solidarity, but restricted it to ethnic Germans. The WAV, a short-lived nationalist party, appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories kranenpohl1998zwischen (59) (Example 12). The BP, responding to Bavaria’s disproportionate refugee burden, called for collective responsibility in crisis, fairer redistribution and showed preference for groups like the Sudeten Germans braun2020101330 (16) (see Example 13,where the BP argued that Sudeten and Southeast Germans should be concentrated in Bavaria because they were ?closest to the Bavarian cultural sphere? and thus easier to integrate). Both parties also expressed the highest levels of anti-solidarity, either by excluding non-German migrants (WAV) connor2018refugees (22) or opposing specific groups such as Silesians (BP).
Figure 12: (Anti-)solidarity frames distribution across major parties in the Bundestag from 1949 to 1957, by election period. Percentages are normalized within each party, such that all subtype bars (solidarity and anti-solidarity) together sum to 100% per party. We include only those parties with a sufficient number of stanced instances $(N>100)$. The full label distribution across all parties is provided in Figure 6 in the Appendix.
The KPD [^10] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) framed displaced populations through a lens of socialist integration, emphasizing group-based solidarity to support their incorporation into the new social order schwartz2008vertriebene (91). This solidarity, however, was assimilationist rather than inclusive: the party avoided the term Vertriebene, in line with Soviet repression of expellee identity heidemeyer2007deutsche (42), and this was reflected in its minimal use of empathic solidarity. The KPD referred almost exclusively to ?refugees?, advocating land reform, compensation, and integration as part of a broader effort to portray them as victims of fascism and capitalism and to gain political support connor2018refugees (22) (see Example 14, where the KPD argued that refugee employment could not be solved under capitalism).
Some parties active in the early postwar years, such as the KPD, WAV, and BP, did not stay in parliament for long, but they played an important role in shaping political debates of the time. Rather than merely reflecting existing social tensions, as suggested by lipset1967cleavage (67), these parties actively contributed to defining what solidarity meant and who was considered deserving of support. As suggested by thijssen2022s (98), such framings were shaped by the political and social challenges of the period and reflected broader efforts to maintain or renegotiate social cohesion.
4.3 Recent German history 2009-2025
Between 2009-2015, migration into Germany increased from just under 800,000 people per year to over 2 mio. people per year gessler2023politicization (41) in the context of the war in Syria. The peak in migration in 2015 has since been associated with the label ?migration crisis? gessler2023politicization (41, 36). Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017. Another spike in migration occurred in 2022 due to the Russian invasion of Ukraine bamf2025 (33). Alongside these historical events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
4.3.1 Overall solidarity distribution
Fig. 13 shows the (anti-)solidarity trends according to Llama-3.3-70b annotations between 2009-2025. We observe a decrease in solidarity between 2014-2018, a somewhat stable period between 2018-2022, and another decrease in solidarity 2022-2025. The decrease in solidarity is associated with an almost corresponding increase in anti-solidarity In terms of (anti-)solidarity subtypes, we observe a peak in compassionate solidarity in 2014, followed by a sharp decrease. The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity. In terms of anti-solidarity, we observe a strong increase in group-based anti-solidarity, accounting for over 50% of all migrant-related statements that either expressed solidarity or anti-solidarity. Exchange-based anti-solidarity exhibited a small increase between 2015-2020 but dropped again with the beginning of the last parliamentary period in 2021.
Overall, the 2020s mark the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements. This key finding is not only visible in the data provided by Llama-3.3-70B but also in human annotations (Figure 2). To provide more insight into this finding, we contextualize it with recent Germany history and a deeper look into party distribution.
(a) Solidarity and anti-solidarity distribution
(b) Solidarity and anti-solidarity frames distribution
Figure 13: Fig. 13(a) shows the fraction of solidarity, anti-solidarity, and mixed stance towards migrants from 2009 to 2025 (instances classified as ?None? are not shown, though included in the calculation). Fig. 13(b) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes according to Llama-3.3-70B, where percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year).
The spike in compassionate solidarity 2014 coincides with a time of heightened public solidarity with Syrian refugees: 15.4% of German survey respondents indicated solidarity as the most important personal value in 2014, up from 12.3% in 2013 (DeLeur2023, 23, p. 38). Syrian refugees were commonly framed as needing help, motivating calls for a ?German welcome culture? Gebauer2023 (39) (see Example 15 in Table 14 in the Appendix). Over the course of 2015, a counter-narrative emerged, framing the influx of refugees as overburdening German capacities and as a ?refugee crisis? Gebauer2023 (39) (see Example 16). During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests and moved notably right, followed by substantial electoral gains in both state and national elections and the AfD entering the German parliament in 2017 Frey2020 (35, 39). Regarding the broader European context, DeLeur2023 (23) diagnose a ?solidarity fatigue? across Europe after 2015, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level. The counter-narratives mentioned by Gebauer2023 (39, 35, 23) match several notions of anti-solidarity according to thijssen2022s (98): refugees as an outgroup that stands in competition to the German population (group-based anti-solidarity), refugees as taking away resources and not contributing sufficiently (exchange-based anti-solidarity), and refugees as not deserving of compassion (compassionate anti-solidarity); see Example 17 where refugees are described as not contributing, straining resources, and threatening the welfare of the ?real? German population. We also note a general shift towards right-wing sentiments in the general German population Zick2023 (108), which may be associated with heightened anti-solidarity.
More recently, the Russian invasion in Ukraine and the subsequent migration of Ukrainian refugees into Germany triggered another spike in public solidarity, albeit much less prominent Weber2024 (104). In our data, solidarity increases slightly in 2022, followed by a continued, rapid decline. Bojarskich2025 (12) identified several factors promoting anti-solidarity towards Ukraine, such as conspiracy mentality and populist voting, but also fears about economic and societal consequences, indicating exchange-based (insufficient contribution) or group-based (competition between groups) anti-solidarity (Example 18). Additionally, anti-Ukrainian framing in parliamentary debates may have been reinforced by the AfD’s long-standing pro-Russian alignment arzheimer2023russia (5), including opposition to sanctions and calls to end German military aid to Ukraine.
In contrast to the postwar period, (anti-)solidarity shifts in the German parliament between 2009 and 2025 appear less attached to specific legislative initiatives. Because of Germany’s membership in the European Union, migration policy has become more complex and is negotiated between European institutions and member states geddes2020migration (40). Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics. As such, we do not look deeper into specific laws passed during this time frame but rather focus on the trends for specific parties.
Figure 14: Raw yearly counts of parliamentary statements by selected political parties, 2009–2025.
4.3.2 Distribution across parties
(a) AfD
(b) CDU/CSU
(c) SPD
(d) FDP
(e) Die Linke
(f) Die Grünen
Figure 15: Distribution of solidarity and anti-solidarity subtypes over time (2009–2025) for selected German parties. For each party, percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year). Note that AfD was first represented in the Bundestag in 2017, and the FDP was not represented from 2013 to 2017. Therefore, data points are absent for these parties during those respective periods.
Fig. 14 shows the raw numbers of migration-related statements per party and year, Fig. 15 shows the fine-grained solidarity trends.
The right-wing AfD
utters almost no expression of solidarity and the anti-solidarity subtypes remain relatively stable over time, with group-based anti-solidarity being consistently dominant (ca. 60% in 2018-2019 and ca. 75% in 2021-2025), although a notable number of samples with exchange-based solidarity (ca. 25%) occurs in the parliamentary period 2017-2021. Although the AfD is not the largest faction in parliament, they account for the largest share of anti-solidarity samples overall (72.5%). The sharp rise of anti-solidarity after 2017 (Figure 13(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure 14). This can be explained by the fact that many AfD speeches, across topics, contain references to migrants illegally crossing the border, criminality of migrants, or migrants imposing additional burden on the German society leaving less resources to address the actual topic of legislative debate (see Example 19 in Table 14 in the Appendix where an AfD member argues that refugees displace elderly people from care homes).
The center-right CDU/CSU and the liberal FDP
changed the most in our data. In the period 2009-2015, we observe a clear majority of solidarity samples, with a spike of compassionate solidarity for the CDU/CSU in 2014 (the FDP was not in parliament at that time). The main difference between both parties is more emphasis on exchange-based solidarity in the FDP, e.g. the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: skilled labor migration law) in 2023 (see Example 20, where the FDP highlights the law as enabling foreign workers to contribute to the German economy). However, for both parties, we observe a rise in group-based anti-solidarity, accounting for 80% of all samples in 2025. When inspecting the samples, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, which is classified as group-based anti-solidarity by Llama-3.3-70B (see Example 21 where a CDU/CSU speaker frames migration as a threat to political stability and an excessive burden on local infrastructure, with calls for deportations, faster procedures, and rejections at the border; and Example 22 where an FDP speaker portrays asylum as being misused for state-subsidized residence).
The drastic shift toward group-based anti-solidarity over the last three years is insufficiently explained by recent literature. Partially, it may be that politicians from CDU/CSU and FDP take over populist rhetoric from AfD politicians who entered parliament since then Esguerra2023Accommodation (31). However,Lewandowsky2022 (65) and kesting2018discourse (53) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality. We also note that the first shift toward group-based anti-solidarity already occurred before the AfD enters parliament in 2017. Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population Zick2023 (108) by appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD. Overall, we cannot establish a full explanation but encourage further research into this striking and rapid trend.
The center-left SPD
also exhibits a spike in compassionate solidarity in 2014 and a slow decline in solidarity towards more group-based anti-solidarity (ca. 20% in 2024). Part of this is explained by mis-classifications of Llama-3.3-70B because SPD speakers often cite anti-solidarity stances of other parties (especially AfD) and these citations are mis-classified as anti-solidarity by the speaker (see Example 23 where an SPD speaker cites AfD proposals putting foreigners under general suspicion in order to criticize them). However, there are also samples where SPD speakers refer positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 24).
The left Green and Linke parties
only display very few instances classified as anti-solidarity and almost all of them are misclassifications where Green or Linke speakers cite anti-solidarity statements or parliamentary motions by other parties (see Example 25 ). These samples are nonetheless interesting because they express a polarizing influence of the AfD entering parliament, leading left-wing parties to positioning themselves in contrast Atzpodien2022 (6). Overall, Die Grünen initially emphasized compassionate solidarity, but later shifted to group-based appeals stressing integration and rights. In the 2024/25 campaign, their rhetoric focused strongly on exchange-based arguments about migrants’ economic and societal contributions Gruene2025 (17) (see Example 26). Similarly, Die Linke began with predominantly compassionate appeals, framing migration as a humanitarian duty, but later emphasized migration as integral to society and advanced a detailed pro-immigration agenda, thereby increasing group-based solidarity. In the 2024/25 campaign, they moved further toward empathic solidarity, criticizing stereotyping and positioning themselves against exclusionary discourses DieLinke2021 (25, 26) (see Example 27). A special case is the BSW, which split off from the Linke party in 2024, partially due to explicit anti-migration positions. However, our data set contains only very few samples from this party (38 statements), so we cannot provide a reliable quantitative analysis.
5 Discussion and Conclusion
In this paper, we have evaluated LLMs on the complex task of annotating (anti-)solidarity in German parliamentary speeches and interpreted the annotated data to derive new insights into (anti-)solidarity towards migrants in the post-war period as well as recent past. In more detail, we provide both methodological and sociological insights:
Methodological perspective.
We find that instruction-tuned LLMs like Llama-3.3-70B and GPT-4 are effective tools for annotating solidarity frames in parliamentary debates, with open models offering a more cost-effective alternative to closed ones. In our experiments, Llama-3.3-70B performed competitively with GPT-4, especially in few-shot settings. Still, no model consistently outperformed GPT-4, and even GPT-4 did not fully attain the human upper bound. Existing research offers mixed evidence on the reliability of LLMs for annotation tasks: some highlight their limitations in handling subjective or expert-defined taxonomies ziems2024can (109), while others show they can substitute or outperform human coders heseltine2024large (43, 99), or suggest that smaller, fine-tuned models are more reliable in political science applications wang2024selecting (102). In our case, model performance generally improved with model size, but also varied by target group, historical period (with older data proving more difficult), and inference method. These findings highlight the absence of a one-size-fits-all solution: rather than relying on a single model, both model and method selection must be adjusted to the requirements of the specific task.
Moreover, we emphasize the need for validation at every stage of the analysis pipeline. Even keyword search as a first filtering step can introduce biases in the data or can aggregate data that concerns very different sub-populations (such as different groups of migrants in the German post-war period). Regarding the LLM annotation itself, we validated model predictions against a large data set of human annotations (aggregating a year of labor involving multiple annotators), test multiple inference configurations (including zero- and few-shot prompting), and validate long-term trends produced by our two best-performing models (Llama-3.3-70B and GPT-4) against human annotations. We also conducted error analyses to identify where and why models fail to make sure that outputs can be reliably interpreted. This aligns with recent work advocating for human-in-the-loop workflows and partial human validation when using LLMs for annotation tasks, to ensure output quality and interpretability pangakis2024keeping (75, 100).
Despite these efforts, it remains essential to acknowledge model-specific weaknesses when interpreting results. For example, we observed that Llama-3.3-70B occasionally misclassifies anti-solidarity, particularly in instances where criticism of anti-migrant rhetoric is sometimes mistaken for anti-solidarity framing itself. In such cases, models may overlook logical direction or target relevance and assign misleading labels despite otherwise coherent reasoning zhang2025mpvstance (107). One possible reason for this behavior could be the influence of complex political language or long prompt context, which may distract the model from the intended target huang2025breaking (48). Importantly, we could only reveal model errors because we checked the primary data before drawing conclusions.
Social science perspective.
Beyond the methodological contributions, our study also demonstrates how an LLM-based pipeline can support a theoretically informed quantitative analysis through two illustrative case studies: one historical (1949–1957) and one contemporary (2009–2025).
For the early post-war period (1949–1957), we demonstrated how solidarity was shaped and institutionalized in the context of national reconstruction. While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insight for historical and social science research. First, it enables us to move beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity, and shifting rhetorical strategies across different migrant groups. Second, our approach helps integrate migration histories that are often treated separately, such as expellees, DPs, and refugees, into a shared analytical framework, allowing for direct comparison of how solidarity was articulated across groups. This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis. This may be partially explained by solidarity fatigue DeLeur2023 (23) or may be part of a larger right-wing shift Zick2023 (108) on migration issues. While stricter migration stances across Europe have been analyzed before, our study is the first to quantify the shift from solidarity to anti-solidarity in German parliament. Our data also revealed that the shift is most striking in the liberal and conservative parties, motivating further research. This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives aktas2024rise (2), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists bayerlein2021chasing (8). Comparative survey evidence further shows that anti-immigrant attitudes remain a central driver of right-wing populist support across the continent messing2021anti (70).
This type of fine-grained, longitudinal analysis of solidarity would hardly be feasible at such scale using manual annotation alone which, in our project, cost over €29,000 for only 3,500 instances. By contrast, with access to computing resources (e.g., in our case, in-house hardware with NVIDIA A40 GPUs), open-weight models such as Llama-3.3-70B can annotate tens of thousands of instances at minimal additional cost. Even closed models like GPT-4 are far cheaper, with one third of our dataset annotated for about €500. While machine annotation is therefore not ?free?, it is vastly more scalable than human annotation and enables large-scale analyses that would otherwise be financially prohibitive. In this sense, our method bridges ?distant reading? approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded systematic interpretation of parliamentary speech while retaining full corpus coverage.
6 Limitations and Future Work
While our analysis was performed carefully to avoid methodological pitfalls, we nonetheless acknowledge remaining limitations of our study.
Human Label Variation:
Even after months of training, discussion and improvements of the annotation manual, we still observe substantial disagreement between human annotators, especially on the fine-grained level. This indicates that (anti-)solidarity is intrinsically a challenging concept to annotate, especially on parliamentary speeches, and that reasonable disagreement may persist (e.g. whether anti-solidarity should be regarded as group-based or exchange-based when competition between groups around limited resources is highlighted). For this investigation, we used the majority label, but future work should consider methods that explicitly account for human label variation and embrace it plank-2022-problem (76).
Political Bias:
It may be that the LLMs compared in this paper may be subject to political bias as investigated by several prior works motoki2024more (72, 85). In particular, it may be that a left-wing political bias leads to more anti-solidarity being annotated. However, we do not observe a systematic bias toward anti-solidarity in the error analysis (Fig. 6); instead, models often mislabel solidarity and anti-solidarity samples as ?None?. It may also be that the human annotators are biased towards certain labels, but this would need to affect all annotators equally because we do not observe a systematic disagreement towards one label, either. Finally, even if bias were present, this would not explain differences and trends in (anti-)solidarity (which are the main subject of our analysis) but rather the overall level of (anti-)solidarity.
Analysis of (Anti-)solidarity towards women:
In the same way we performed an analysis of (anti-)solidarity trends towards migrants, the (anti-)solidarity towards women could be analyzed. For example, recent work by Diabate2023 (24) suggests that there may be a shift to the left in Germany regarding gender, which would be an opposite trend compared to what we found regarding migration. Due to space and time constraints, such an analysis is not included in this work but appears highly promising for future work.
Further model variations:
We considered a range of models as well as different prompts, and fine-tuning (for Llama models and BERT models) but a vast range of possible model variations remains to be explored, such as ensembles of models, less quantized models, newer models, etc. It may well be that human annotation quality can be fully matched with more refined tools.
Temporal Reliability:
Both model performance (macro F1) and human agreement vary over time, with notable dips in some decades (e.g., the 1960s for Llama-3.3-70B; Fig. 16). This reflects the difficulty of annotating historical language, but overall trend correlations between GPT-4, Llama-3.3-70B, and human annotations remain high. Since our analysis emphasizes broad temporal trends rather than isolated data points, these fluctuations do not undermine the validity of our conclusions.
Appendix
Appendix A List of Keywords
For Frau (en.: woman) we use Frauen, Frau, Mutter, Mädchen, Mütter, Ehefrau, Müttern, Hausfrauen, Hausfrau, Ehefrauen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Großmutter, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
Since Frau is the German word for woman but also for Mrs./Ms., we only include occurrences of the word Frau that are not followed by a capitalized word (which are probably surnames).
For Migrant (en.: refugee) we use Flüchtlinge, Ausländer, Flüchtlingen, Zuwanderung, Vertriebenen, Ausländern, Asylbewerber, Migranten, Migration, Heimatvertriebenen, Aussiedler, Einwanderung, Ansiedler, Vertriebene, Zuwanderer, Asylbewerbern, Flüchtling, Heimatvertriebene, Sowjetzonenflüchtlinge, Aussiedlern, Einwanderer, Asylsuchenden, Asylsuchende, Bürgerkriegsflüchtlinge, Zuwanderern, Ansiedlern, Migrantinnen, Vertriebener, Emigranten, Kriegsflüchtlinge, Ausländerinnen, and Immigranten.
When running stability tests (see Subsubsection 4.1.1), we ensure that each sample includes at least 5 keywords, covers at least 10% of the dataset, and spans at least 75% of the timeline, such that enough data is present to create the plots. For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred. The trends are shown in the Fig. 19 in the Appendix.
Appendix B Parties
For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag. For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by poradaautomatische (77).
Each new intervention is marked in the XML by a <redner> element containing both the speaker’s name and their party. During preprocessing (XML parsing, transcript cleaning, and sentence segmentation), we convert these into standardized <SPEAKER> tags and extract the associated party information. During extraction of migrant- or women-related sentences, the most recent <SPEAKER> marker (with party) is assigned to each sentence until a new one appears.
List of the parties included in the dataset, along with the variations of their names or abbreviations as they have been recorded: AfD (Alternative for Germany); Die Linke (The Left) with variations such as PDS, Gruppe der PDS; Bündnis 90/Die Grünen (Alliance 90/The Greens); CDU/CSU (Christian Democratic Union/Christian Social Union); SPD (Social Democratic Party of Germany); FDP (Free Democratic Party); DP (German Party) with variations such as DP/DPB, DP/FVP, FVP; GB/BHE (All-German Bloc/League of Expellees and Deprived of Rights); KPD (Communist Party of Germany); BP (Bavarian Party); WAV (Economic Reconstruction Union).
Appendix C Models Inference and Training Details
Hardware and Inference Settings.
All inference runs were conducted on an NVIDIA A40 (80GB) with temperature set to 0.6 and top-p to 0.9. For Mistral-Large-Instruct-2407, we used an NVIDIA A100 to accommodate memory requirements.
Prompt Format.
Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure 24 for Migrant and in Figure 27 in the Appendix for Woman).
In few-shot settings, one representative example is included for each class label. GPT-4 receives all instructions in a single prompt (see the prompt text in Figure 21 for Migrant and Figure 22 in the Appendix for Woman). Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
Baseline fine-tuning details.
For our baseline, we use a BERT-based pipeline with 110M parameters, comprising a high-level category classifier and two subtype models for solidarity and anti-solidarity. All models share a similar architecture, processing inputs with a target (Frau or Migrant) and full text comprising the focus sentence and left and right context. We add a fully connected layer with softmax activation atop the pooled output of the BERT-based models, with 4 output units for each model. To counter class imbalance, minority classes are oversampled to parity with the majority class. We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer AdamW (68) and categorical crossentropy loss.
Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 reimers2019sentence (82). We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions. The model incorporates ReLU activation and dropout (0.1), and maps the embeddings to four and ten target classes for high-level and fine-grained tasks, respectively. To manage class imbalance, we use oversampling. The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
Fine-Tuned Llama Models
We fine-tune Llama-3-8B-It on a single A40 GPU. Llama-3-70B-It is trained across two A40s using PyTorch FSDP and QLoRA [^11] https://www.philschmid.de/fsdp-qlora-llama3. Training spans 3 epochs with a constant learning rate of 2e-4, AdamW optimizer, and max gradient norm of 0.3.
Label Woman Migrant Total per label
Group-based Solidarity 112 (7.8%) 270 (12.4%) 382 (10.6%)
Exchange-based Solidarity 54 (3.8%) 79 (3.6%) 133 (3.7%)
Empathic Solidarity 125 (8.7%) 32 (1.5%) 157 (4.3%)
Compassionate Solidarity 732 (51.0%) 692 (31.7%) 1424 (39.3%)
Solidarity (no subtype) 41 (2.9%) 55 (2.5%) 96 (2.7%)
Total Solidarity 1064 (74.1%) 1128 (51.6%) 2192 (60.5%)
Group-based Anti-solidarity 10 (0.7%) 288 (13.2%) 298 (8.2%)
Exchange-based Anti-solidarity 0 (0.0%) 55 (2.5%) 55 (1.5%)
Empathic Anti-solidarity 17 (1.2%) 6 (0.3%) 23 (0.6%)
Compassionate Anti-solidarity 7 (0.5%) 97 (4.4%) 104 (2.9%)
Anti-solidarity (no subtype) 5 (0.3%) 19 (0.9%) 24 (0.7%)
Total Anti-solidarity 39 (2.7%) 465 (21.3%) 504 (13.9%)
Mixed 60 (4.2%) 149 (6.8%) 209 (5.8%)
None 273 (19.0%) 442 (20.2%) 715 (19.8%)
Total 1436 (100.0%) 2184 (100.0%) 3620 (100.0%)
Table 2: Distribution of labels in the human-annotated dataset for Woman and Migrant instances. Percentages are relative to each column total.
Figure 16: Macro F1 scores over time for each model (best configuration), and average pairwise Cohen’s Kappa for human annotators. Evaluated on the full test set, including Frau and both test sets for Migrant, grouped by decade.
Figure 17: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
Figure 18: Fine-grained confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
Table 3: F1 scores per model for fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman. The top section contains fine-grained categories, and the bold rows show performance on high-level classification. Scores are chosen for the best performing configuration for the model in terms of Macro F1.
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.33 0.30 0.29 0.25 0.25 0.34 0.32 0.16
Exchange-based Solidarity 0.56 0.39 0.36 0.25 0.41 0.20 0.21 0.19
Empathic Solidarity 0.04 0.27 0.10 0.08 0.15 0.14 0.00 0.03
Compassionate Solidarity 0.48 0.42 0.51 0.22 0.45 0.43 0.51 0.41
Solidarity (high-level) 0.78 0.52 0.69 0.60 0.58 0.63 0.66 0.51
Group-based Anti-solidarity 0.47 0.25 0.42 0.32 0.42 0.41 0.28 0.20
Exchange-based Anti-solidarity 0.46 0.29 0.35 0.12 0.28 0.29 0.09 0.22
Empathic Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Compassionate Anti-solidarity 0.00 0.18 0.07 0.00 0.19 0.11 0.17 0.12
Anti-solidarity (high-level) 0.64 0.53 0.52 0.44 0.56 0.54 0.35 0.35
Mixed 0.38 0.26 0.43 0.10 0.17 0.19 0.24 0.15
None 0.67 0.60 0.63 0.56 0.57 0.58 0.59 0.51
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.28 0.24 0.25 0.00 0.20 0.20 0.24 0.08
Exchange-based Solidarity 0.31 0.00 0.00 0.00 0.22 0.30 0.35 0.15
Empathic Solidarity 0.27 0.36 0.41 0.24 0.33 0.41 0.22 0.14
Compassionate Solidarity 0.43 0.32 0.51 0.00 0.37 0.21 0.62 0.52
Solidarity (high-level) 0.85 0.67 0.79 0.51 0.75 0.70 0.77 0.78
Group-based Anti-solidarity 0.31 0.36 0.22 0.00 0.27 0.26 0.00 0.00
Exchange-based Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Empathic Anti-solidarity 0.67 0.00 0.25 0.07 0.00 0.17 0.00 0.00
Compassionate Anti-solidarity 0.50 0.33 0.22 0.36 0.15 0.00 0.00 0.00
Anti-solidarity (high-level) 0.65 0.44 0.39 0.31 0.48 0.35 0.12 0.10
Mixed 0.30 0.40 0.37 0.16 0.32 0.30 0.15 0.26
None 0.62 0.54 0.47 0.41 0.66 0.41 0.21 0.40
Table 4: Agreement between GPT-4, Llama-3.3-70B, and human annotations on the Migrant dataset. Subtable LABEL:tab:model-human-agreement shows instance-level agreement; Subtable LABEL:tab:trend-agreement reports trend-level correlation across decades.
Comparison Cohen’s $\kappa$ / Macro F1 N
GPT-4 vs Llama-3.3-70B 0.47 (0.58) / — 17,696
GPT-4 vs Human 0.38 (0.57) / 0.34 (0.63) 2,084
Llama-3.3-70B vs Human 0.36 (0.50) / 0.35 (0.61) 2,084
Label Pearson (p)
Group-based Solidarity 0.560 ($p=2.41\times 10^{-2}$)
Exchange-based Solidarity 0.777 ($p=4.00\times 10^{-4}$)
Empathic Solidarity 0.108 ($p=0.691$)
Compassionate Solidarity 0.915 ($p=7.10\times 10^{-7}$)
Solidarity (high-level) 0.966 ($\boldsymbol{p=1.25\times 10^{-9}}$)
Group-based Anti-solidarity 0.945 ($p=3.65\times 10^{-8}$)
Exchange-based Anti-solidarity 0.898 ($p=2.29\times 10^{-6}$)
Empathic Anti-solidarity 0.160-0.160 ($p=0.553$)
Compassionate Anti-solidarity 0.841 ($p=4.44\times 10^{-5}$)
Anti-solidarity (high-level) 0.976 ($\boldsymbol{p=1.15\times 10^{-10}}$)
Mixed 0.685 ($p=0.0034$)
Figure 19: Distribution of all Migrant keywords over the years, normalized per keyword. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
Figure 20: Percentage of sentences showing solidarity/anti-solidarity per decade for all Migrant keywords from 1867 to 2025. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
Figure 21: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Migrant dataset.
Figure 22: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Frau dataset.
Figure 23: (Anti-)solidarity trends in 1867-2022, restricted to the keywords Flüchtlinge and Ausländer. Figure 23(a) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes predicted by Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes, so the curves in both panels together sum to 100% per decade. Figure 23(b) shows the aggregated proportions of solidarity, anti-solidarity, and mixed labels per decade; totals may not reach 100% because instances classified as ?None? are omitted (though included in the calculation). The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
(a) High-level prompt (Step 1).
Figure 24: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 25: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 26: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
(a) High-level prompt (Step 1).
Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
Keyword Solidarity Anti-Solidarity Mixed None Total
Asylsuchende 10 0 0 5 15
Ausländer 168 52 9 525 754
Aussiedler 99 1 5 61 166
Einwanderer 23 1 0 21 45
Emigranten 66 5 0 18 89
Flüchtlinge 1828 23 84 1179 3114
Heimatvertriebene 2471 20 68 1814 4373
Sowjetzonenflüchtlinge 264 5 10 178 457
Zuwanderer 32 0 2 46 80
Zuwanderung 7 0 1 29 37
Table 5: Distribution of Solidarity, Anti-Solidarity, Mixed, and None labels per keyword included in the post-war trends analysis (Section 4.2).
BP CDU/CSU DP DRP DZP FDP GB/BHE KPD SPD WAV Z
Group-based S. 95 707 56 13 23 221 118 50 667 93 5
Exchange-based S. 5 163 5 4 1 46 14 8 73 13 0
Compassionate S. 64 670 38 9 8 135 92 69 537 72 7
Empathic S. 6 91 8 1 0 39 29 4 89 8 1
Solidarity with no subtype 1 18 0 0 0 6 2 1 6 2 0
Group-based AS 28 100 9 3 1 22 9 30 68 23 0
Exchange-based AS 4 14 3 0 0 1 3 3 11 2 0
Compassionate AS 1 3 1 0 0 0 0 3 3 0 0
Empathic AS 0 1 0 0 0 0 0 0 0 0 0
Mixed 15 63 5 1 2 23 8 4 33 3 0
None 81 1564 78 10 11 389 94 54 754 47 4
Total per party 300 3394 203 41 46 882 369 226 2241 263 17
Table 6: Distribution of (anti-)solidarity subtypes by party included in the post-war trends analysis (Section 4.2). AS stands for Anti-Solidarity; S for Solidarity.
Original Text Translation
(1) Predicted label: group-based solidarity towards expellees
(May 14, 1952)
?[…] das Lastenausgleichsgesetz und die Eingliederung der Vertriebenen [steht] in keinem Widerspruch zu unserm Anspruch und unserm Recht auf die Heimat. Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden. […] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten; […] aber der Rechtsanspruch darf nicht so behandelt werden, als sei er nur etwas Zweitrangiges, angewiesen auf das, was übrigbleibt.? ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland. On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy. […] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations; […] but the legal entitlement should not be treated as secondary, dependent only on what is left over.?
(2) Predicted label: exchange-based solidarity towards expellees
(Feb 25, 1953)
?Ich darf […] auf einige Gesichtspunkte hinweisen […] zum Bundesvertriebenengesetz. Der Ihnen vorliegende Entwurf geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen. Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.? ?I would like to point out some aspects […] on the Federal Expellees Act. The draft before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy. The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
(3) Predicted label: compassionate solidarity towards Soviet-zone refugees
(June 15, 1955)
?[…] es ist notwendig, zu den noch immer vielseitigen Problemen der Sowjetzonenflüchtlinge Stellung zu nehmen […]. Wir erkennen selbstverständlich an, daß durch die Neuregelung des Härtefonds im Lastenausgleich eine Verbesserung für die Sowjetzonenflüchtlinge erfolgte, weil die Ermessensfragen, unter denen dieser Personenkreis bislang besonders zu leiden hatte, dadurch eingeschränkt werden. […] Wenn wir aber die Sicherheit haben wollen, daß sich diese Neuregelung auch wirklich positiv für den Flüchtling auswirkt, so ist es notwendig, daß die vorgesehene Erhöhung des Härtefonds […] wirklich realisiert wird.? ?[…] it is necessary to address the still manifold problems of the Soviet-zone refugees […]. We naturally recognize that the reform of the hardship fund in the Equalization of Burdens has brought improvements […] because the discretionary decisions that this group had particularly suffered from have been curtailed. But if we want to be sure that this reform really has a positive effect for the refugee, then […] the planned increase of the hardship fund […] must actually be implemented.?
(4) Predicted label: group-based anti-solidarity towards expellees
(Dec 13, 1950)
?In unseren Landkreisverwaltungen sind heute 42,2% Heimatvertriebene. […] Auch die einheimische Bevölkerung beklagt sich […] man höre auf den Landratsämtern überwiegend ostpreußisch, und das Plattdeutsche sei zur Seite gedrängt. Unter den Lehrern […] 51% Heimatvertriebene, unter den vollbeschäftigten Richtern […] mehr als 50%, unter der Polizei sogar 61%; zweifellos eine Überfremdung, die Besorgnis erregt! Aber all diese ehrlichen Bemühungen um Eingliederung der Heimatvertriebenen sind angesichts der ungeheuren Zahl vergeblich. Arbeitslosenzahl […] beträgt unter den Heimatvertriebenen fast 60%. Die Lasten sind unerschwinglich.? ?In our district administrations today, 42.2% are expellees. The native population […] complains […] that in the district offices one mostly hears East Prussian, and that Low German has been pushed aside. Among the teachers 51% are expellees, among the full-time judges also more than 50%, among the police even 61% – undoubtedly an over-foreignization that arouses concern! But all these honest efforts to integrate the expellees are in vain in view of their immense numbers. Unemployment […] almost 60% among the expellees […]. The burdens are unbearable.?
Table 7: Examples for postwar period discussed in Section 4.2. Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(5) Predicted label: group-based solidarity towards Soviet-zone refugees
(Sept 23, 1954)
?[…] den Ländern [wurden] für die wohnungsmäßige Unterbringung der […] Sowjetzonenflüchtlinge […] 350 Millionen DM […] zur Verfügung gestellt […]. Der Bundestag ist […] davon ausgegangen, daß die Sowjetzonenflüchtlinge im Lastenausgleich nicht berücksichtigt werden konnten und daß die Verweisung […] auf den Härtefonds nur als Interimslösung gedacht war. […] Die Bundesregierung prüft […] noch […] Maßnahmen, um den bei den Sowjetzonenflüchtlingen auftretenden Notständen abzuhelfen.? ?[…] the federal government provided 350 million DM for the housing of […] Soviet-zone refugees […]. Parliament assumed that Soviet-zone refugees could not be included in the Equalization of Burdens and that referral […] to the hardship fund was intended only as an interim solution. […] The federal government is still examining measures to address the emergencies arising among the Soviet-zone refugees.?
(6) Predicted label: group-based anti-solidarity towards foreigners
(June 21, 1956)
?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen. […] Für die Bundesrepublik kommt […] hinzu, daß sie […] Anlaß hat, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden. […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.? ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements. […] For the Federal Republic there is the added circumstance that it has reason to shield itself against tens of thousands of stateless foreigners who […] come with the intention of seeking economic support or political asylum here. […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
(7) Predicted label: group-based solidarity towards foreigners
(Oct 18, 1950)
?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige. Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen wie Deutschen gewährt werden. […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […]. Der Gesetzentwurf geht […] über den Konventionsentwurf, dessen Grundsätze Mindestforderungen für die Behandlung heimatloser Ausländer darstellen, hinaus. […]? ?[…] In public welfare too, they are to receive benefits equal to those of German citizens. They are to be granted access to schools of all kinds under the same conditions as Germans. […] the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state […]. The draft law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners. […]?
(8) Predicted label: group-based solidarity towards foreigners
(April 19, 1951)
?[…] Wir begrüßen die Verabschiedung des Gesetzes für heimatlose Ausländer, weil uns damit die Möglichkeit geboten war, dem Ausland zu beweisen, wie wir Angehörige anderer Nationen in unserem eigenen Staate zu behandeln gewillt sind, und weil wir uns […] erhoffen, daß sie […] dazu beitragen wird, die Frage der Internationalisierung unseres Flüchtlingsproblems und des Zuflusses von Mitteln aus dem Ausland zu fördern. […]? ?[…] We welcome the adoption of the law for stateless foreigners, because it gave us the opportunity to demonstrate to foreign countries how we intend to treat nationals of other nations within our own state, and because we […] hope that it will make a contribution to promoting the internationalization of our refugee problem and the inflow of funds from abroad. […]?
(9) Predicted label: empathic solidarity towards foreigners
(June 15, 1955)
?[…] Es ist kein Zweifel, daß die heimatlosen Ausländer natürlich vom Arbeitsministerium und vom Innenministerium […] eine rein materielle Unterstützung bekommen. Aber diese materielle Unterstützung gibt keine Möglichkeit, ihnen kulturell oder sonstwie unter die Arme zu greifen. […] [Es ist sehr schwer] diese Dinge heute den Ländern nahezulegen.? ?[…] There is no doubt that the stateless foreigners, of course, receive […] purely material support from the Ministry of Labor and the Ministry of the Interior. But this material support provides no opportunity to assist them culturally or in other ways. […] [It is very difficult] today to persuade the Länder on these matters. […]?
Table 8: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(10) Predicted label: exchange-based solidarity towards expellees,
Party: FDP
(Oct 20, 1949)
?Die werden […] von früh morgens bis spät abends arbeiten, wenn man ihnen das nötige Gelände, [Baumaterialien oder das Geld] zur Verfügung stellt. So werden sie in Gemeinschaftsarbeit wesentlich mehr leisten, als wir das sonst heute bei diesen Bauten sehen. In diesem Falle wäre das Geld, das wir dafür aufnehmen, praktisch nur als Kredit anzusehen. Wir hätten die Möglichkeit, daß diese Heimatvertriebenen sich ihr Heim selbst bauen, daß es allmählich ihr Eigentum wird. Wenn wir nur normale Mietsätze verrechnen und das Kapital ohne Zinsen oder zu geringen Zinssätzen verzinsen lassen, so wäre in wenigen, spätestens 15 Jahren, dieses Haus […] Eigentum des Betreffenden. ? ?They would […] work from early morning until late at night if they were provided with the necessary land, [building materials or the money]. In cooperative work they would achieve far more than we otherwise see with such construction today. In this case the money we raise would practically only be regarded as a loan. We would have the possibility that these expellees build their homes themselves, gradually becoming their property. If we only charged normal rents and let the capital bear no or low interest, then within at most fifteen years this house […] would become the property of the person concerned.?
(11) Predicted label: group-based solidarity towards expellees,
Party: GB/BHE
(Oct 29, 1953)
?[…] Wir dürfen die Augen auch nicht vor der Tatsache verschließen, daß Millionen von Vertriebenen und Kriegsgeschädigten noch nicht den Anschluß an unsere soziale Entwicklung gefunden haben. Auch hier mußten bisher Zehntausende arbeitswilliger tüchtiger Mitbürger abseitsstehen. Bei unserem Verlangen nach beschleunigter Eingliederung dieser Bevölkerungskreise geht es nicht nur um die Ablösung einer unproduktiven Fürsorge durch aufbauende Mitarbeit. Es geht auch nicht nur um die Erfüllung von ?Rechtsansprüchen?, und ich darf diesen Ausdruck Rechtsansprüche hier noch einmal betonen. ? ?[…] We must not close our eyes to the fact that millions of expellees and war victims have still not been integrated into our social development. Here too, tens of thousands of willing and capable fellow citizens have so far had to stand aside. In our demand for the accelerated integration of these groups, it is not only a matter of replacing unproductive welfare with constructive participation. It is also not only a matter of fulfilling legal claims — and I would like to emphasize this expression ?legal claims? once again.?
(12) Predicted label: compassionate solidarity towards refugees,
Party: WAV
(March 23, 1950)
?Die Aufnahme dieser Flüchtlinge wird in manchen Ländern bewußt erschwert, sabotiert und unmöglich gemacht. […] Seitdem die Spannung zwischen Osten und Westen zunimmt, hat ein neuer Strom von Flüchtlingen eingesetzt. […] Polen und die Tschechei [durchführen] die neue Ausweisungsaktion auf Befehl Moskaus, […] um das Flüchtlingsheer in Westdeutschland zu vermehren und noch mehr Arbeitslose zu schaffen, den Staat zu belasten und seine Wirtschaft zu zerstören. […]? ?The admission of these refugees is being deliberately obstructed, sabotaged, and made impossible in some states. […] Since the tension between East and West has increased, a new stream of refugees has begun. […] Poland and Czechoslovakia are carrying out the new expulsions on orders from Moscow, […] in order to enlarge the refugee army in West Germany, to create even more unemployed, to burden the state, and to destroy its economy. […]?
(13) Predicted label: group-based anti-solidarity towards expellees,
Party: BP
(Oct 20, 1949)
?Die Sudetendeutschen […] und die Südostdeutschen [stehen] dem süddeutschen, also dem bayerischen Kulturkreis am nächsten. […] Die aus dem Sudetenland bzw. aus Südosteuropa kommenden Deutschen [sollen] zunächst in Süddeutschland zusammen[gefasst werden], […] damit die Eingliederung […] schneller und reibungsloser vor sich gehen kann. [Wir] wollen […] nicht Schwierigkeiten schaffen, wo Schwierigkeiten nicht überwunden werden müssen, und nicht Gräben aufreißen, wo solche nicht vorhanden sind. Das hat mit Rassentheorie […] gar nichts zu tun. […] ? ?The Sudeten Germans […] and the Southeast Germans are closest to the South German, that is, the Bavarian cultural sphere. […] The Germans coming from the Sudetenland and from Southeastern Europe [should] initially be concentrated in southern Germany, […] so that integration […] can proceed more quickly and smoothly. […] we do not want to create difficulties where difficulties need not arise, and we do not want to open up rifts where none exist. This has nothing to do with racial theory […]. ?
Table 9: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(14) Predicted label: group-based solidarity towards refugees,
Party: KPD
(Dec 13, 1950)
?Wir glauben, daß das Problem der Beschaffung von Arbeit für [Umsiedler] unter dem kapitalistischen System nicht in der Form gelöst werden kann […]. Wir sind also der Auffassung, daß die Umsiedlung eine Notwendigkeit ist […]. Wenn hier der Sprecher der CDU das schöne, stolze und richtige Wort geprägt hat ?Unser Haus ist deutsch?, wenn er am Schluß gesagt hat ?Das ganze Deutschland soll es sein?, wie kann man dann […] gleichzeitig die Tatsache, daß sich die Vertriebenen, Flüchtlinge und Umsiedler, wie Sie sie bezeichnen, politisch zu Parteien zusammengeschlossen haben, als die Gefahr einer Überfremdung bezeichnen? ? ?We believe that the problem of providing work for [resettlers] cannot be solved under the capitalist system […]. We therefore take the view that resettlement is a necessity […]. If the CDU speaker here has coined the fine, proud, and correct phrase ?Our house is German?, and if he concluded by saying ?It shall be the whole of Germany?, how then can one at the same time describe the fact that the expellees, refugees, and resettlers, as you call them, have politically organized themselves into parties as a danger of being overrun by foreigners??
Table 10: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(15) Predicted label: compassionate solidarity towards refugees
Party: Die Grünen
(Nov 6, 2014)
?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen. […] Ich will ganz klar sagen: Das Boot ist nicht voll; das Boot ist niemals voll. Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […]. ? ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us. […] I want to state very clearly: the boat is not full; the boat is never full. Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
(16) Predicted label: group-based anti-solidarity towards refugees
(Oct 1, 2015)
?[…] Wenn die Zahlen weiterhin auf diesem hohen Niveau bleiben, dann wird Deutschland über kurz oder lang überfordert sein. Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […]. allein im September sind mehrere zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen. Sie wurden nicht registriert. Sie wurden nicht kontrolliert. […] Damit besteht auch ein großes Sicherheitsrisiko. ? ?[…] If the numbers remain at this high level, Germany will sooner or later be overburdened. Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […]. In September alone, several tens of thousands of refugees and asylum seekers came to Germany. They were not registered. They were not controlled. […] This poses a major security risk.?
(17) Predicted label: exchange-based anti-solidarity towards refugees
Party: AfD
(May 16, 2018)
?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug nach derzeitigem Stand um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind. In Deutschland, meine Damen und Herren, lebt bereits eine große soziale Unterschicht […]. Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant. Meine Damen und Herren, Deutschland ist kein reiches Land, sondern ein leistungsfähiges. ? ?[…] each year about 140,000 highly qualified people leave this country, while in return, according to the current situation, about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate. In Germany, ladies and gentlemen, there already lives a large social underclass […]. Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people. Ladies and gentlemen, Germany is not a rich country, but a capable one.?
Table 11: Examples for the 2009-2025 period discussed in Section 4.3. Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(18) Predicted label: group-based anti-solidarity towards refugees
Party: AfD
(Oct 19, 2022)
?[…] Echte Kriegsflüchtlinge aus der Ukraine brauchen Schutz – keine Frage. Wir sagen aber Nein zu Fluchttourismus durch halb Europa, und wir sagen Nein zu bevorzugter Behandlung auf Kosten der deutschen Steuerzahler. Wir fordern daher die sofortige Abschaltung dieses Sozialstaatsmagneten, […] Wir fordern außerdem die Umstellung von Geldleistungen auf Sachleistungen für Asylbewerber, damit Migrationsanreize gesenkt werden und der Missbrauch effektiv unterbunden wird. […] Und wer sich wie Frau Baerbock in erster Linie der ukrainischen Bevölkerung verpflichtet fühlt, der soll sich in der Ukraine um ein Amt bewerben, aber nicht in Deutschland. ? ?[…] Genuine war refugees from Ukraine need protection — no question. But we say no to flight tourism across half of Europe, and we say no to preferential treatment at the expense of German taxpayers. We therefore demand the immediate shutdown of this welfare magnet, […] We also demand the switch from cash benefits to benefits in kind for asylum seekers, so that incentives for migration are reduced and abuse is effectively curbed. […] And whoever, like Ms. Baerbock, feels primarily obligated to the Ukrainian population should run for office in Ukraine, but not in Germany.?
(19) Predicted label: group-based anti-solidarity towards refugees
Party: AfD
(Dec 5, 2023)
?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt. […] Dafür werden bereits Zimmer in angeblich christlichen Altenheimen gekündigt, weil angebliche Flüchtlinge lukrativer sind, Messerzuwanderung inklusive. […] Bei diesem Recht auf Wohnungstausch der Linken handelt es sich um eine Mischung aus Neid zulasten alter Leute und Zwang zulasten von Vermietern. Es ist die Fortsetzung einer inländerfeindlichen… ? ?[…] elderly people are supposed to move out of their large apartments because green-left politics invites millions of guests from Africa and the Middle East while simultaneously stifling housing construction. […] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration [^12] Messereinwanderung: a propagandistic term coined by the AfD implying migrants bring knife crime. included. […] This ?right to apartment swaps? proposed by the Left is a mixture of envy at the expense of elderly people and coercion at the expense of landlords. It is the continuation of an anti-native policy…?
(20) Predicted label: exchange-based solidarity towards immigrants
Party: FDP
(May 25, 2023)
?[…] Menschen, die bei uns arbeiten wollen, sollten wir eine echte Chance geben; denn viele, die an unserer Grenze ankommen, wollen ja gerade arbeiten. Wir arbeiten in der Ampel an der Reform des Fachkräfteeinwanderungsgesetzes. Da geht es um mehr Regulierung, gezielte Einwanderung in unseren Arbeitsmarkt. Menschen aus dem Ausland sollten eine realistische Chance bekommen, zu uns zu kommen und arbeiten zu können. Das Fachkräfteeinwanderungsgesetz wird das jetzt endlich besser machen und mehr Menschen die Chance geben, die unsere Wirtschaft auch dringend braucht. […] ? ?[…] People who want to work here should be given a real chance; because many who arrive at our border actually want to work. Within the coalition we are working on reforming the Skilled Immigration Act. This is about more regulation and targeted immigration into our labor market. People from abroad should have a realistic chance to come to us and be able to work. The Skilled Immigration Act will now finally improve this and give more people the opportunity that our economy urgently needs. […]?
(21) Predicted label: group-based anti-solidarity towards migrants
Party: CDU/CSU
(May 14, 2025)
?[…] Die gute Nachricht ist: Der Politikwechsel hat begonnen. […] Die illegale Migration der letzten zehn Jahre gefährdet die politische Stabilität Deutschlands und Europas. […] Unsere Städte und Gemeinden, die Landkreise: Alle sind über dem Limit. In den Kitas, in den Schulen, auf dem Wohnungsmarkt, bei der Sicherheit an den Bahnhöfen und auf den Marktplätzen wird diese Überforderung sehr konkret. Deswegen ist es gut, dass wir gemeinsam im Koalitionsvertrag die Migrationswende vereinbart haben: mehr Rückführungen, schnellere Verfahren und, ja, ebenso die Zurückweisung an den deutschen Grenzen auch bei Asylgesuchen. ? ?[…] The good news is: the policy shift has begun. […] The illegal migration of the last ten years endangers the political stability of Germany and Europe. […] Our towns and municipalities, the districts—everyone is over the limit. In day-care centers, in schools, in the housing market, in security at train stations and in marketplaces, this overburdening becomes very concrete. That is why it is good that we have agreed together in the coalition agreement on a migration turnaround: more deportations, faster procedures, and, yes, also rejections at the German borders, even in the case of asylum applications.?
Table 12: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(22) Predicted label: group-based anti-solidarity towards immigrants
Party: FDP
(Jan 30, 2025)
?Das Grundrecht auf Asyl darf nicht dazu führen, dass allein ein Antrag auf Asyl zu mehreren Jahren freiem und staatlich subventioniertem Aufenthalt in Deutschland führt, wenn er völlig ohne Grundlage erfolgt. Solche Missstände müssen wir beseitigen, damit wir uns auf jene konzentrieren können, die unsere Hilfe wirklich brauchen […]. SPD und Grüne […] wollen die illegale Einwanderung nach Deutschland kaum begrenzen und fordern lieber weiteren Familiennachzug. Diesen falschen Weg ist Deutschland leider viel zu lange gegangen. […] Eine immer linkere Politik wird […] dazu führen, dass die AfD ihr teuflisches Werk immer weiter fortsetzen kann. […] ? ?The fundamental right to asylum must not result in a situation where merely submitting an asylum application leads to several years of free and state-subsidized residence in Germany when it is entirely unfounded. Such abuses must be eliminated so that we can focus on those who truly need our help […]. The SPD and the Greens […] hardly want to limit illegal immigration to Germany and instead prefer to demand further family reunification. Unfortunately, Germany has followed this wrong path for far too long. […] Ever more left-leaning policies […] will only enable the AfD to continue its diabolical work […].?
(23) Predicted label: group-based anti-solidarity towards foreigners
Misclassified; true: group-based solidarity
Party: SPD
(Apr 10, 2024)
?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht. Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird. Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […]. Sie ?übersehen? in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben. Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen. Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist. […] ? ?Let us turn to the AfD’s draft and to what makes it simply atrocious. You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved. Not only does this place all foreigners under general suspicion […], but it also violates EU law […]. You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency. But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is. […]?
(24) Predicted label: group-based anti-solidarity towards migrants
Party: SPD
(Dec 9, 2024)
?Das Sicherheitspaket, das heute hier zur Debatte steht, ist der erste Schritt, um Konsequenzen aus der furchtbaren Katastrophe in Solingen zu ziehen. Waffenverbotszonen […] helfen [aber] Waffenverbotszonen nicht. Deshalb brauchen wir mehr Befugnisse für das Bundeskriminalamt und das Bundesamt für Migration. Sie müssen in die Lage versetzt werden, mit den öffentlich zugänglichen Informationen Straftaten zu verhindern. Das ist dringend erforderlich, und deshalb werden wir das jetzt umsetzen. Wir werden weitere Schritte veranlassen, damit ausreisepflichtige Personen unser Land verlassen. ? ?The security package being debated here today is the first step toward drawing consequences from the terrible catastrophe in Solingen. Weapons prohibition zones […] do not, however, help against terrorists. That is why we need more powers for the Federal Criminal Police Office and the Federal Office for Migration. They must be put in a position to use publicly available information to prevent crimes. That is urgently necessary, and that is why we will now implement it. We will take further steps to ensure that persons obliged to leave the country actually leave our country.?
(25) Predicted: group-based anti-solidarity towards migrants
Misclassified; true: empathic solidarity
Party: Die Linke
(June 5, 2025)
?[…] Migration wird pauschal als sicherheitspolitisches Problem dargestellt. Die AfD führt die sogenannte gescheiterte Migrationspolitik in einer Liste von Krisen auf – neben dem Krieg in der Ukraine, der Coronapandemie und der hybriden Bedrohung. Damit betreibt sie eine gefährliche und spaltende Gleichsetzung. Sie erklärt das bloße Vorhandensein von Migration zur Bedrohung der inneren und äußeren Sicherheit und nennt sie in einem Atemzug mit Terrorismus, Krieg und Katastrophen. ? ?Migration is presented across the board as a security policy problem. The AfD lists the so-called failed migration policy alongside the war in Ukraine, the coronavirus pandemic, and hybrid threats. In doing so, it engages in a dangerous and divisive equation. It declares the mere existence of migration to be a threat to internal and external security, mentioning it in the same breath as terrorism, war, and disasters.?
Table 13: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(26) Predicted label: exchange-based solidarity towards immigrants
Party: Die Grünen
(Jan 19, 2024)
?Wissenschaftliche Institute sprechen von 400 000 Menschen jedes Jahr. Darauf müssen wir die Gesellschaft und die Politik vorbereiten. Wir müssen also ohnehin die entsprechende Infrastruktur ausbauen – Wohnungen, Schulen, Kinderbetreuung. Wir brauchen gezielte Zuwanderung, und dafür haben wir das Fachkräfteeinwanderungsgesetz gemacht. Aber wir wären doch mit dem Klammerbeutel gepudert, wenn wir nicht auch das Potenzial der Menschen, die aus humanitären Gründen zu uns kommen, nutzten und dafür sorgten, dass sie aus- und weitergebildet werden und Teil unserer Gesellschaft werden. Das wäre die Hauptaufgabe. Und weil wir Zuwanderung brauchen, brauchen wir eine Willkommenskultur. ? ?Scientific institutes speak of 400,000 people every year. Society and politics must be prepared for that. We therefore have to expand the relevant infrastructure anyway — housing, schools, childcare. We need targeted immigration, and for that we created the Skilled Workers Immigration Act. But we would have to be out of our minds not to also make use of the potential of people who come to us for humanitarian reasons, and ensure that they receive education and training and become part of our society. That would be the main task. And because we need immigration, we need a culture of welcome.?
(27) Predicted label: empathic solidarity towards migrants
Party: Die Linke
(Jan 30, 2025)
?Sie sagen, Sie wollen Frauen und Kinder schützen. Sie wollen Kinder und Frauen in Kriegsgebiete abschieben. So ist es. Wir führen eine Debatte über Sicherheit und Migration. Und leider erklären alle Parteien außer der Linken, dass sie mit Abschottung dieses Land sicherer machen wollen, und reden von ?Abschiebungen im großen Stil?, als seien Migranten grundsätzlich eine Gefahr. So werden Menschen unter Generalverdacht gestellt. Weniger Migration mache das Land sicherer, das ist doch Unsinn. ? ?You say you want to protect women and children. You want to deport children and women to war zones. That is the reality. We are having a debate about security and migration. And unfortunately, all parties except the Left declare that they want to make this country safer through isolation, and talk about ?deportations on a large scale,? as if migrants were inherently a danger. In this way, people are placed under general suspicion. Less migration makes the country safer — that is nonsense.?
Table 14: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.