### v2

# LLM Analysis of 150+ Years of German Parliamentary Debates on Migration Reveals a Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

Abstract
Migration has been a core topic in German political debate, from the postwar displacement of millions of expellees to labor migration and recent refugee movements. Studying political speech across such wide-ranging phenomena in depth has traditionally required extensive manual annotation, limiting analysis to small subsets of the data. Large language models (LLMs) offer a potential way to overcome this constraint. Using a theory-driven annotation scheme, we examine how well LLMs annotate subtypes of solidarity and anti-solidarity in German parliamentary debates and whether the resulting labels support valid downstream inference. We first provide a comprehensive evaluation of multiple LLMs, analyzing the effects of model size, prompting strategies, fine-tuning, historical versus contemporary data, and systematic error patterns. We find that the strongest models, especially GPT-5 and gpt-oss-120B, achieve human-level agreement on this task, although their errors remain systematic and bias downstream results. To address this issue, we combine soft-label model outputs with Design-based Supervised Learning (DSL) to reduce bias in long-term trend estimates. Beyond the methodological evaluation, we interpret the resulting annotations from a social-scientific perspective to trace trends in solidarity and anti-solidarity toward migrants in postwar and contemporary Germany. Our approach shows relatively high levels of solidarity in the postwar period, especially in group-based and compassionate forms, and a marked rise in anti-solidarity since 2015, framed through exclusion, undeservingness, and resource burden. We argue that LLMs can support large-scale social-scientific text analysis, but only when their outputs are rigorously validated and statistically corrected.
1 Introduction
Migration is a defining feature of the 21st century(Wilkinson et al., 2022). Economic imbalances, demographic divergence, and climate change make migration are increasing pressures for migration, especially from the Global South to the Global North (World Bank, 2023; Almulhim et al., 2024; Marois et al., 2020). If governed responsibly, it can contribute to prosperity and support sustainable development (Gavonel et al., 2021; Adger et al., 2024).
Within Europe, Germany is a major migration destination (European Commission, Eurostat, 2024), and migration has profoundly shaped its development, from post-Second World War expellee resettlement to labor migration in the 1950s–1970s and refugee movements in the 2010s-2020s (Kaya, 2017; Fröhlich, 2025). As the EU’s largest economy, Germany remains an attractive destination, while demographic decline and skilled-labor shortages make continued immigration economically beneficial (Bertoli et al., 2016; Donal Smith and Huang, 2025; Eissel et al., 2021; Angenendt et al., 2023). At the same time, anti-migration stances have become more visible in political debate, often in relation to perceived risks such as pressure on public services, labor-market and resource competition, and the rise of right-wing populism (Frey, 2020; Gessler and Hunger, 2023; Messing and Ságvári, 2021; Esguerra et al., 2023). This raises a central question: how does German politics discuss migration, and to what extent does it express solidarity or anti-solidarity toward migrants?
Answering such questions requires systematic analysis of political discourse across long time spans to assess continuity and change in migration discourse. Such analysis has traditionally relied on manual textual analysis and qualitative interpretation (Randour et al., 2020). However, such methods are labor-intensive (Bhattacherjee, 2012) and, hence, often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized (Babalola and Nwanzu, 2021). Computational social science (CSS) emerged partly in response to the scale and complexity of modern data, which exceed the analytic reach of traditional approaches (Lazer et al., 2009). More recently, large language models (LLMs) have been introduced into CSS as a means of scaling annotation and automating parts of the research workflow (Pavlovic and Poesio, 2024). At the same time, their use raises several methodological challenges. (i) Although LLMs show strong performance on tasks such as ideological stance detection (Ziems et al., 2024), it remains unclear whether they can perform more complex, theory-driven sociopolitical annotation tasks. (ii) Even when LLMs appear to perform well, recent work questions whether their outputs can support valid downstream inference or serve as ground-truth labels, given their sensitivity to model choice, prompt design, and configuration, as well as their potential to introduce systematic bias (Baumann et al., 2025; Gligoric et al., 2025; Krsteski et al., 2025; Egami et al., 2023). (iii) These concerns connect to ongoing debate in CSS emphasizing that computational methods and digital data must be tightly integrated with social scientific theory, concepts, and theoretically meaningful research questions (Lazer and Ognyanova, 2024; Jungherr and Theocharis, 2017; Theocharis and Jungherr, 2021; Soemer et al., 2025).
In a prior conference paper, we addressed challenge (i) by showing that LLMs can approximate human performance on a complex annotation task focused on (anti-)solidarity in parliamentary text and grounded in a theoretically informed, fine-grained taxonomy (Kostikova et al., 2024). In this paper, we extend that analysis in three ways: the first two are methodological, and the third is substantive. First, to address (i) more rigorously, we evaluate a broader range of LLMs, compare their predictions against a larger set of human reference annotations, and systematically assess the effects of prompts, model size, and historical versus contemporary data. Second, to address (ii), we examine whether strong annotation performance is sufficient for valid downstream inference by analyzing systematic model errors, testing ensemble predictions, and applying statistical bias correction (Egami et al., 2023). Third, to address (iii), we apply this validated pipeline to the analysis of solidarity and anti-solidarity toward migrants in German parliamentary speeches and interpret the resulting trends between 1949 and 1957 and between 2009 and 2025 in light of German political history.
Figure 1: Overview of the annotation scheme adapted from Thijssen (2012). Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none. Solidarity and anti-solidarity instances are then further categorized into subtypes based on the underlying rationale: group-based, exchange-based, compassionate, and empathic. The example sentence in the figure illustrates exchange-based anti-solidarity.
We focus on solidarity because it is a particularly demanding test case for operationalization and annotation, both for human annotators and for LLMs. Solidarity is an essentially contested concept: it is difficult to define, value-laden, ambiguous, multi-faceted, context-dependent, and open to change over time (Kozłowska, 2024). This makes annotation demanding, as coders must map concrete and historically variable political statements onto abstract categories of (anti-)solidarity. To address this challenge, we approach solidarity as a theoretically informed category and apply solidarity typology of Thijssen (2012), which offers a structured, annotatable framework. It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure˜1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification). Germany is a particularly suitable context for such a study since solidarity is deeply rooted in its political vocabulary (Fiegle, 2003; Stjernø_2005; Reisz, 2006), and its multi-party system enables to trace how meanings of (anti-)solidarity vary across parties and over time.
Overall, we find that in our task, the strongest models approach human agreement, and in some settings even surpass it, with GPT-5 performing best overall and gpt-oss-120B being the strongest open model. However, we conclude that raw model predictions are not sufficient for reliable downstream inference: although strong models achieve near-human agreement at the instance level, their errors remain structured and correlated, limiting the benefits of simple ensemble methods. At the same time, human agreement itself remains limited, especially at the fine-grained level, indicating substantial label variation. In response, combining soft-label model outputs with statistical correction, such as Design-based Supervised Learning (DSL) (Egami et al., 2023), achieves more accurate recovery of human-derived temporal trends. Substantively, this validated pipeline reveals high levels of migrant-directed solidarity in the postwar period, especially forms centered on shared belonging and support for vulnerable groups (group-based and compassionate), thereby corroborating existing qualitative research on early postwar inclusion politics. For the contemporary period, it reveals a marked shift since 2015 toward anti-solidarity in German parliamentary discourse, framed through exclusion, undeservingness, and burdens on collective resources (group-based, compassionate and exchange-based). To our knowledge, this is the first large-scale quantitative evidence of that shift.
1.1 Related Work
Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; (ii) the use of LLMs as annotation tools in political and social science contexts; and (iii) valid downstream inference from imperfect machine-generated labels.
Solidarity in Political Discourse and NLP
The theoretical foundation for our analysis is the solidarity framework of Thijssen (2012), which extends Durkheim’s distinction between mechanical and organic solidarity by incorporating recognition-based theories of Honneth (1996, 2014). Thijssen distinguishes four subtypes of solidarity—group-based, compassionate, exchange-based, and empathic—applicable to both historical and contemporary discourse. Thijssen and Verheyen (2022) apply this typology to Belgian party manifestos, showing how parties emphasize different frames in line with ideological divides. Survey and interview studies also examine solidarity longitudinally (Schönweitz et al., 2024; Sengupta et al., 2023), but these approaches are not suited for large-scale text analysis or fine-grained classification. We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs.
Solidarity has received relatively little attention in computational text analysis. Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements (Weber et al., 2024b; Eger et al., 2022), the COVID-19 pandemic (Kalaloi et al., 2021), and natural disasters (Santhanam et al., 2019). These studies focus on specific, time-bound crises and typically treat (anti-)solidarity as a binary classification task, whereas our work applies a differentiated, theory-based typology to parliamentary discourse.
LLMs as Annotation Tools in Political and Social Science
LLMs are increasingly used in political and social science research. Prior work has considered annotation of political ideology (e.g. Heseltine and Clemm von Hohenberg, 2024), sentiment (Lashitew and Mu, 2025; Fu et al., 2024), stance (e.g. Cao et al., 2024), and misinformation (e.g. Huang et al., 2025). For broader surveys of LLM applications in social and political science, see Thapa et al. (2025); Li et al. (2024); Ziems et al. (2024). Conclusions about LLM performance vary: some find near-human (Rytting et al., 2023) or superior results on short texts (Le Mens and Gallego, 2025; Heseltine and Clemm von Hohenberg, 2024), while others show smaller fine-tuned models perform better with labeled data or expert taxonomies (Wang et al., 2024; Burnham et al., 2024; Ziems et al., 2024). Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories. Erhard et al. (2025) develop PopBERT, a fine-tuned transformer for detecting populist rhetoric in German Bundestag speeches, while Card et al. (2022) analyze 140 years of U.S. political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, frames, and dehumanizing metaphors. Other work fine-tune models for pragmatic and rhetorical categories in German debates, such as speaker attribution, speech acts, and opinion roles (Reinig et al., 2024; Rehbein et al., 2024; Rehbein and Ponzetto, 2024). In contrast to fine-tuned approaches, Ranjit et al. (2024) use general-purpose LLMs (e.g., GPT-4, Flan-T5) with a custom typology of attitudes toward homelessness to annotate large-scale Twitter data and study aggregate trends.
Valid Inference from Imperfect LLM Labels
Recent work moves beyond raw annotation accuracy to ask whether machine-generated labels actually support valid downstream inference. As Egami et al. (2023) stress, downstream uses of surrogate labels often ignore measurement error, that is, the unknown, heterogeneous mismatch between gold-standard and surrogate labels, which can distort estimates and invalidate inference even when surrogate labels appear highly accurate (Egami et al., 2023; Knox et al., 2022). In CSS, such error is especially difficult to avoid because annotation tasks are often intrinsically challenging (Ziems et al., 2024), and prior work shows that LLM annotations often diverge from expert coding decisions, especially on more complex tasks (Lin and Zhang, 2025; de Pieuchon et al., 2024). In response, bias-corrected estimators combine machine-generated labels with smaller sets of gold-standard annotations to support valid downstream inference. These include Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023, 2024), Design-Based Supervised Learning (DSL) (Egami et al., 2023), and Confidence-Driven Inference (CDI) (Gligorić et al., 2025). We build most directly on DSL because it was developed with CSS applications in mind and fits our design, which combines a fixed decade-stratified human-coded sample with multiple historical trend analyses. Recent benchmarking also suggests that DSL often yields lower bias and more precise estimates than PPI in text classification tasks (de Pieuchon et al., 2025).
DSL has been evaluated in a limited number of benchmark and applied settings. Egami et al. (2023) assess DSL in a custom regression task and in class prevalence estimation across CSS text classification datasets from Ziems et al. (2024). Rister Portinari Maranca et al. (2025) apply DSL to AI-assisted image labeling, and Krsteski et al. (2025) to LLM-generated survey responses. In settings closer to ours, Stuhler et al. (2026) study U.S. media representations of the timing and urgency of climate change using a hybrid pipeline of rule-based extraction and RoBERTa-based classifiers, and apply DSL to debias trend and regression estimates. Liu (2025) examines Chinese-language social media discourse around two major 2024 U.S. political events using LLM-based coding of misinformation, partisan stance, affective polarization, and attack language. There, LLM annotation is the main measurement framework, while DSL is presented as a robustness check and is mainly implemented in the user-engagement analysis rather than across all analyses.
Across work using both raw and bias-corrected model predictions, our study differs in several respects. Unlike Erhard et al. (2025); Card et al. (2022); Stuhler et al. (2026), we use state-of-the-art prompted LLMs and avoid the need for task-specific fine-tuning, and apply DSL directly to their outputs. Unlike Ranjit et al. (2024) and Liu (2025), which also use prompted LLMs, we apply a theory-driven typology to more than 150+ years of parliamentary speech rather than social media text, and we make LLM validation and DSL-based bias correction central to inference across all analyses rather than treating them as robustness checks, thus enabling fine-grained, longitudinal insight into political discourse.
2 Methodology
2.1 Data
We build on the DeuParl dataset introduced in Walter et al. (2021), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022111Protocols from the Volkskammer (the GDR parliament) are not included due to lack of availability. As a result, GDR-specific narratives and solidarity frames are absent from the corpus.. We extend this corpus to June 2025, following the preprocessing steps from Walter et al. (2021) and including metadata such as date, session number, and legislative period.
We reuse the keyword sets (see Appendix˜A in the Appendix) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in Kostikova et al. (2024), comprising 32 and 18 terms, respectively. Based on these, we extract 63k migrant-related and 138k women-related sentences (hereafter, instances) from DeuParl, including the three preceding and three following sentences as context. We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) (Porada, 2019); see Appendix˜B in the Appendix for details. We include women alongside migrants in the evaluation to test model generalizability across distinct marginalized groups, though our substantive analysis focuses on migrants.
Figure˜2 shows the number of instances over time and their share relative to all DeuParl sentences. Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015 during refugee crisis debates. Notably, mentions of migrants also surpass women-related discourse in the immediate post-war years.
Figure 2: Absolute number of instances (left) and their share relative to all DeuParl sentences (right) for the Woman and Migrant categories.
2.2 Data Annotation
To validate our models and enable qualitative analysis, we use a human-annotated dataset created in two batches. The first, introduced in Kostikova et al. (2024), comprises 2,864 instances (1,428 migrant-related, 1,437 woman-related) from DeuParl, annotated by five trained student assistants. The annotation guidelines were refined over multiple rounds of discussion among the annotators and authors; details of the annotation process and the final guidelines are provided in Appendix˜F and Appendix˜G in the Appendix. The second adds 756 migrant-related instances with broader historical coverage (pre-1950), annotated by four of the same assistants following the same procedure.
Each instance is labeled according to a fine-grained solidarity framing scheme distinguishing four high-level stance categories – solidarity, defined as willingness to share resources Lahusen and Grasso (2018); Ils et al. (2021), anti-solidarity (restriction or exclusion), mixed (both supportive and opposing elements), and none (neutral or unrelated content). Solidarity and anti-solidarity are further assigned one of four subtypes adapted from Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance), and empathic (respect for or rejection of diversity). We refer to the subtypes as fine-grained annotation.
The final gold-standard label for each instance is determined by majority vote.For the first batch, inter-annotator agreement reached a Cohen’s $\kappa$ of 0.62 for the high-level labels and 0.42 for the fine-grained labels. For the second batch, agreement reached 0.47 for the high-level labels and 0.37 for the fine-grained labels. We attribute the lower agreement in the second batch to the inclusion of older historical data, which we discuss further in Section˜3.1. Statistics on the human annotated dataset are shown in Figure˜3. Label distributions differ by target group: woman-related instances are predominantly solidarity (74.1%), while migrant-related instances are more diverse, with less solidarity (51.6%) and more anti-solidarity (21.3%).
Figure 3: Distribution of human annotated instances by time and target group. Total indicates the number of instances per group (including None); see Table˜6 in the Appendix for exact counts.
2.3 Models and Experiments
We evaluate a diverse set of models to cover variation in model size, provider, and access setting (open vs. closed), while also including supervised classifier baselines. Our selection comprises:
  1. (i)
    smaller pre-trained LMs such as BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details are provided in Appendix˜C in the Appendix;
  2. (ii)
    closed-source LLMs such as gpt-4-1106-preview and gpt-5-chat;
  3. (iii)
    open-weight LLMs, including:
    • gpt-oss-20B and gpt-oss-120B;
    • Mistral-2-Large (123b parameters);
    • Llama-4-Scout-17B-16E-Instruct;
    • Llama-3.3-70B-It, Llama-3-8B-It;
    • Qwen2.5-72B-Instruct;
    • Gemma-2-9B-It and Gemma-2-27B-It.
Data and Metrics
We adopt a 70/15/15 train/dev/test split of the first annotated batch (Kostikova et al., 2024), generating three random splits of about 2,000 training, 400 development, and 430 test (Test 1) instances in total, covering both target groups. Supervised models are trained and evaluated on these splits, while inference-only experiments use two test sets: Test 1 (original) and Test 2 (756 newly added migrant-related instances, 1860s–1950s). We report Macro F1 to account for class imbalance, along with per-class F1 for high-level (four stance categories) and fine-grained (ten labels) classification. [^2] High-level: solidarity, anti-solidarity, mixed, none. Fine-grained: four solidarity subtypes, four anti-solidarity subtypes, plus mixed and none.
Prompting Setup
For the inference-only experiments with LLMs, We use manually designed prompts for each target group. Each prompt consists of a classification instruction concatenated with the parliamentary instance (the keyword sentence plus three sentences of left/right context as defined in Section˜2.1). Speaker and party metadata are available in the corpus but removed from the model input. Instructions include (i) chain-of-thought cues in zero-/few-shot settings (think step by step); (ii) label definitions; and (iii) a two-step format: high-level classification followed by subtype prediction. [^3] For GPT-4, both steps are combined in a single prompt (Figure 25 and Figure 26 in the Appendix); for open-weight models, we use two prompts: one for high-level classification and, conditional on that, a second for subtype selection (Figure 27 and Figure 30 in the Appendix). In few-shot mode, one representative example per label is included.
Model Ensemble
We additionally evaluate a model ensemble to test whether combining strong open-weight LLMs improves performance relative to individual models. The ensemble setup also resembles the human annotation procedure in that final labels are derived from multiple judgments rather than a single annotator. The ensemble combines the best-performing open models, each in its strongest configuration (Table˜1), using majority vote. The label with the most votes is selected as the final prediction; in tie cases, votes are weighted by each model’s macro-F1 score as measured on the validation set. We report ensemble results in Section˜3.1.3.
Design-based Supervised Learning (DSL)
When employing LLMs as annotators, it may be that single labels are systematically preferred compared to human annotators, yielding biased estimates and incorrect conclusions. We use Design-based Supervised Learning (DSL; Egami et al. 2023) to correct such systematic biases. DSL combines large-scale surrogate outputs with a subsample of human-coded data to construct bias-corrected pseudo-outcomes for inference. In our application, the surrogate outputs are soft label distributions derived from a committee of LLMs.
For each parliamentary speech $i$, we observe surrogate labels $Q_{i}$ and covariates $X_{i}$. For a (small) subset, we also have human annotations $Y_{i}$. The surrogate labels $Q_{i}$ are soft-label distributions over annotation categories, computed as the proportion of committee models assigning each label, while the human labels $Y_{i}$ are distributions over categories constructed from multiple human annotators. Human annotations are observed only for a subset of instances under a decade-stratified annotation design. We assign each instance a decade-specific inclusion probability $\pi_{i}$, computed as the proportion of instances in that decade that received human annotation, and treat these as fixed design probabilities in the DSL correction.
DSL corrects downstream estimands by constructing a bias-adjusted pseudo-outcome
$\tilde{Y}_{i}=\hat{g}(Q_{i},X_{i})+\frac{R_{i}}{\pi_{i}}\big(Y_{i}-\hat{g}(Q_{i},X_{i})\big),$ (1)
where $R_{i}$ indicates whether instance $i$ is human-annotated, and $\hat{g}(Q_{i},X_{i})$ estimates the conditional expectation of human labels given surrogate outputs and covariates. We model $\hat{g}$ using a multinomial regression trained on soft targets and implement $K$-fold cross-fitting.
Because outcomes are multi-class, DSL is applied in a vector-valued form, giving a vector of DSL-adjusted class scores $\tilde{Y}_{i}$ for each instance. We use them directly in downstream analyses.
We estimate separate DSL models for the high-level and fine-grained tasks to avoid carrying subtype-level error from fine-grained predictions into aggregated high-level estimates.
3 Results
In this section, we evaluate the models in two steps. We first assess instance-level annotation performance against human labels, including overall performance, task-specific challenges, and systematic error patterns in Section˜3.1. We then examine whether these model outputs support reliable downstream inference by comparing hard-label, soft-label, and DSL-corrected trend estimates against human-derived reference trends in Section˜3.2.
3.1 Annotation Performance at the Instance Level
Table˜1 reports macro F1 scores for each model under inference and fine-tuned settings. We benchmark against the leave-one-out F1 score of human annotators, [^4] For each annotator, we exclude their labels, compute a consensus from the others, and calculate the macro F1 score between the excluded annotator and the consensus; scores are then averaged across annotators. and show results for both high-level and fine-grained tasks (the latter in parentheses). We evaluate instance-level annotation performance along four dimensions: (i) overall performance of individual models; (ii) task- and label-related challenges; (iii) model ensemble performance; and (iv) error analysis.
Model Method Woman Migrant (T1) Migrant (T2)
Closed-weight models
GPT-4 Zero-shot 0.37 (0.60) 0.42 (0.73) 0.29 (0.62)
Few-shot 0.37 (0.54) 0.43 (0.63) 0.26 (0.55)
GPT-5 Zero-shot 0.36 (0.53) 0.42 (0.67) 0.30 (0.60)
[…]
Large open-weight models
gpt-oss-20B Zero-shot 0.25 (0.43) 0.27 (0.55) 0.15 (0.49)
Few-shot 0.28 (0.47) 0.40 (0.56) 0.22 (0.49)
gpt-oss-120B Zero-shot 0.28 (0.55) 0.38 (0.66) 0.25 (0.55)
[…]
Mistral-2-Large Zero-shot 0.20 (0.50) 0.27 (0.56) 0.24 (0.49)
Few-shot 0.26 (0.51) 0.38 (0.57) 0.26 (0.46)
Llama-4-Scout Zero-shot 0.25 (0.46) 0.28 (0.52) 0.20 (0.50)
Few-shot 0.25 (0.46) 0.41 (0.63) 0.25 (0.50)
Qwen-2.5 Zero-shot 0.21 (0.40) 0.37 (0.62) 0.23 (0.52)
Few-shot 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
Llama-3.3-70B Zero-shot 0.23 (0.50) 0.32 (0.54) 0.25 (0.47)
[…]
Mid- and small-scale open-weight models
Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.15 (0.40)
Few-shot 0.14 (0.35) 0.19 (0.34) 0.12 (0.42)
Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.21 (0.46)
Few-shot 0.25 (0.55) 0.33 (0.48) 0.21 (0.47)
Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.21 (0.47)
Few-shot 0.23 (0.44) 0.31 (0.54) 0.23 (0.47)
Classical baselines
BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.22 (0.48)
SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.17 (0.37)
Human reference 0.32 (0.67) 0.48 (0.78) 0.30 (0.58)
Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks. ?Migrant (T1)? is the original benchmark; ?Migrant (T2)? is newly introduced in this paper. Green cells mark the best score per target group (model’s best setting); yellow cells mark the best score among open models (second best overall). Gray text indicates suboptimal settings for a given model.
3.1.1 Overall model performance
  • GPT-5 leads, but some open models offer a strong alternative. Across Woman and both Migrant test sets, GPT-5 (few-shot) matches human annotators on the high-level task and surpasses them on the fine-grained task. Among open models, gpt-oss-120B matches the human reference value in fine-grained performance and comes close to GPT-4 and GPT-5 at the high level.
    Among the remaining open-weight models, Llama-4-Scout, Qwen-2.5, and Llama-3.3-70B show the strongest performance. These models achieve comparable results across targets and evaluation levels (e.g., on Migrant T1, fine-grained F1 scores range from 0.38 to 0.43 and high-level scores from 0.59 to 0.65). While their relative strengths vary by setting, where Qwen-2.5 and Llama-4-Scout achieve the strongest fine-grained scores, and Llama-3.3-70B the highest high-level scores, their overall performance remains closely clustered.
  • Model size correlates with performance, but not perfectly. Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure˜4). However, this relationship is non-monotonic: moving from small to mid-sized models brings clear benefits, with notable exceptions such as Llama3-8B, which underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B. Among large models ($\approx$70B–120B+ parameters), performance gains diminish. Moreover, mid-sized models can match or outperform larger systems; for example, gpt-oss-20B clusters with Qwen-72B and Mistral-Large-2, while gpt-oss-120B approaches GPT-4/5 on the fine-grained benchmark. These results suggest that factors beyond parameter count influence model performance.
    Figure 4: Model size vs. performance on high-level and fine-grained classification in Migrant (T1). Each model is plotted with its best-performing configuration (zero- or few-shot). The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
  • Examples help mainly mid-sized models. Most models benefit from few-shot prompting, with the most consistent improvements observed for mid-sized models such as Gemma models, Mistral-2-Large, Llama-3.3-70B and Llama-4-Scout. Smaller models like Llama-3-8B show only marginal gains, while GPT-4 demonstrates minimal improvement, likely due to its strong zero-shot performance. Overall, the effect of few-shot prompting is widespread but not uniform across model scales.
3.1.2 Task and label-related challenges
  • All models (and human annotators) perform worse on the fine-grained task, which indicates its complexity.
  • Solidarity is easier to detect than anti-solidarity or mixed stance (Figure˜5). Even smaller models perform well on solidarity, especially in the Woman category (e.g., BERT 0.77, Gemma-9B 0.70). In contrast, anti-solidarity and mixed stances remain challenging, with anti-solidarity being most difficult for Woman. This pattern is likely due in part to class imbalance (Table˜6 in the Appendix), especially because anti-solidarity is very rare for Woman and mixed is less common in both target groups. That said, GPT models (both closed-source models and gpt-oss-120B) achieve the strongest anti-solidarity performance (0.66–0.70 for Migrant; 0.53–0.65 for Woman), while mid-sized models reach 0.56–0.59. Only GPT-4/5, gpt-oss-120B and Llama-3.3-70B consistently capture mixed stances, with F1 scores between 0.30 and 0.50. Llama-4-Scout shows some asymmetry for Migrant: strong performance on solidarity (0.78) and anti-solidarity (0.63) but near failure on mixed stance (0.08).
  • On the fine-grained level, GPT-5, gpt-oss-120B, Llama-3.3-70B and Llama-4-Scout offer the most balanced performance on the fine-grained level (see Table˜7 in the Appendix). Best fine-grained macro F1 is achieved by GPT-5. Among non-GPT-4/5 models, Llama-3.3-70B and Llama-4-Scout follow. Nevertheless, all models perform poorly on specific subtypes, including empathic solidarity (e.g., 0.11 for Migrant with Llama-3.3-70B) and compassionate anti-solidarity (e.g., 0.06 for Migrant with gpt-oss-120B). Empathic anti-solidarity fails across all models. For Woman, all models consistently struggle with fine-grained anti-solidarity.
    Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For Migrant, the two test sets (Test 1 and Test 2) are combined.
  • Performance drops on historical data. All models, including the strongest, perform worse on the historical Migrant (T2) set. This reflects a broader trend: performance declines on older data (see Figure˜18 in the Appendix), as does human agreement (Cohen’s Kappa). We examine errors on the historical data in more detail in Section˜3.1.4.
3.1.3 Model Ensemble Performance
We construct a majority-vote ensemble from the best-performing open models in Table˜1: Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B. We exclude GPT-5 due to its higher cost and because the subsequent social-scientific analysis focuses on open models.
We find that the ensemble does not outperform the strongest individual model (gpt-oss-120B), indicating that simple majority voting does not consistently improve performance in this setting (see Table˜2). This is in line with prior work showing that ensemble methods are most effective when individual models achieve comparable performance while making at least partly uncorrelated errors (Trad and Chehab, 2024).
Model Woman Migrant (T1) Migrant (T2)
gpt-oss-120B (Few-shot) 0.38 (0.55) 0.48 (0.69) 0.28 (0.57)
Llama-3.3-70B (Few-shot) 0.30 (0.50) 0.38 (0.65) 0.29 (0.58)
Qwen-2.5 (Few-shot) 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
Ensemble 0.34 (0.60) 0.47 (0.70) 0.29 (0.60)
Table 2: Performance (macro F1) of models included in the ensemble and the resulting majority-weighted ensemble on fine-grained and high-level tasks (in parentheses). In tie cases, votes are weighted by validation-set macro-F1. Tie-breaking weights are based on validation-set performance, ranking gpt-oss-120B >> Llama-3.3-70B >> Qwen-2.5 for the migrant setting and Llama-3.3-70B >> gpt-oss-120B >> Qwen-2.5 for the woman setting.
In summary, GPT-5 achieves the highest overall performance, surpassing human annotators on fine-grained level and approaching them on the high-level task. However, its closed-weight nature and high cost make it less suitable for our large-scale social scientific analysis. Among open models, gpt-oss-120B performs comparably to GPT-5 and approaches human-level performance (even matching it on the fine-grained level), making it the strongest open-weight alternative.
That said, the fact that these models reach or exceed human reference values does not imply that they can be used as direct substitutes of human annotators. First, the observed macro-F1 scores remain low enough that substantial and systematic model errors can still occur, including errors that human annotators are less likely to make; moreover, recent work shows that even high agreement with human consensus can coexist with systematic misclassification patterns (Bavaresco et al., 2025). We therefore examine these patterns through error analysis (Section˜3.1.4).
Second, in our setup, the human reference is derived from leave-one-out agreement and therefore inherits the uncertainty present in the human annotations. As a result, perfect F1 may itself be unattainable because some disagreement among human annotators likely reflects genuine label variation. In a subjective task like ours, this motivates moving beyond hard labels to consider label distributions instead (Section˜3.2.1). We address both issues next.
3.1.4 Error Analysis
We analyze errors of the strongest models through their confusion matrices on the full Migrant test set, comparing them with a human reference aggregated from leave-one-out (LOO) annotator comparisons. [^5] Because the human panel compares each annotator to a LOO consensus, whereas the model panels compare predictions to gold labels, we use row-normalized confusion matrices so that error patterns for each reference label are more directly comparable. Figure˜6 shows the top models; full matrices appear in Figure˜19 and Figure˜20 in the Appendix. We consider four aspects of model error: (i) high-level confusion; (ii) confusion within (anti-)solidarity subtypes; (iii) errors in historical instances (1867-1900); and (iv) ensemble failure analysis, with particular attention to error correlation across models. Since our substantive analysis focuses on migrant-related discourse, we restrict the error analysis to the Migrant category.
(a) High-level confusion matrices.
(b) Fine-grained confusion matrices.
Figure 6: Row-normalized confusion matrices (%) for the best-performing configuration of each of the top models by macro-F1, as reported in Table˜1, on the two Migrant test sets combined (Test 1 and Test 2). Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
(i) High-level confusion
Most models recover the broad high-level structure of the task. GPT-5 and gpt-oss-120B show the clearest separation between solidarity, anti-solidarity, and none, but even these stronger models do not fully match human error patterns. Relative to the human LOO panel, models systematically overassign cases to none, suggesting that they often fail to detect low-intensity or indirect stance expressions, despite the prompt instruction to consider even slight expressions of solidarity regardless of the main topic of the text Example 11 in Table˜14).
The models also preserve mixed much less often: for reference-mixed cases, the share predicted as mixed is 42% for humans, compared with 27% for GPT-5, 22% for gpt-oss-120B and Qwen, and 33% for the ensemble; Llama-3.3-70B is the only exception, reaching 47%. Instead, models resolve these cases into one-sided solidarity or anti-solidarity labels, suggesting that models overlook weaker cues of the opposing stance, as in Example 11.
A smaller but qualitatively important error type arises when speakers from party A quote or paraphrase anti-migrant rhetoric from party B to criticize party B and its stance. In such cases, models sometimes misclassify the speaker’s stance as anti-solidarity (e.g. Llama and Qwen in Examples 11 and 12).
(ii) Confusion within (anti-)solidarity subtypes
Within solidarity, and focusing on the better-represented labels (Table˜6), human disagreement is greatest between group-based and compassionate solidarity: in the human LOO matrix, 18% of group-based solidarity cases are labeled as compassionate solidarity by the held-out annotator. Models show less confusion in this direction (5-7%), but more often predict group-based solidarity for human-labeled compassionate solidarity instead (e.g. 32% for gpt-oss-120B). Humans also identify compassionate solidarity more consistently overall (66% vs. roughly 40% for models).
The two subtypes can co-occur in migration debates. Human annotators reflect this ambiguity, whereas models tend to shift such cases toward group-based solidarity, apparently relying more heavily on cues such as integration, rights, and legal inclusion (Examples 12, 12, and 13). However, rights language alone does not imply group-based solidarity. When rights claims are framed in terms of vulnerability, deprivation, or need, compassionate solidarity is the stronger reading, as it aims to place the other ?in an equal position as I am?. When such claims are framed in terms of shared membership or common rights and duties, group-based solidarity is more appropriate (Thijssen, 2012; Thijssen and Verheyen, 2022).
A related pattern is observed for other subtypes: confusion occurs when broader integration language associated with group-based solidarity overlaps with characteristics of other frames, as in Example 13. Since Thijssen and Verheyen (2022) describe the categories as conceptually close and the coding as interpretive, we treat such variation as meaningful uncertainty rather than noise. Within anti-solidarity, the sharpest subtype confusions appear in the rarer categories, especially empathic and compassionate anti-solidarity. These cases are often collapsed into group-based anti-solidarity, suggesting that models over-rely on broad cues of protecting the interests of the native in-group (see Example 14). Nevertheless, some models also show confusion among the better-represented anti-solidarity categories such as exchange-based anti-solidarity and group-based anti-solidarity (22% cases for Llama-3.3-70B, 21% for humans, see Example 13).
(a) High-level confusion matrices.
(b) Fine-grained confusion matrices.
Figure 7: Row-normalized high-level and fine-grained confusion matrices (%) for the best-performing configurations of the top models on the combined Migrant test sets, restricted to 1867-1900. As in Figure˜6, rows sum to 100, Human Annotators (LOO) denotes aggregated leave-one-out annotator comparisons, and the Ensemble aggregates Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
(iii) Confusion for historical instances
The historical sample (restricted to 1867-1900, Figure˜7) shows an even stronger overprediction of none. Across nearly all models, errors for many true classes collapse into none, more than in the human LOO panel; this is particularly visible for Qwen-2.5-72B, Llama-3.3-70B, and the ensemble. Human disagreement also increases on the historical subset, butannotators more often distribute errors across neighboring stance categories. Examples 14 and 14 indicate why this matters for historical inference. Here, solidarity is expressed through legal and constitutional reasoning rather than overt emotional language, and models fail to recognize such low-intensity stance at all.
Overall, strong models reproduce the broad label structure of the task, but their errors remain systematic rather than random. As a result, trends in (anti-)solidarity and mixed stance are likely to be underestimated, especially in older instances. Within subtypes, group-based solidarity is likely to absorb other solidarity claims, while rarer categories are often collapsed or disappear entirely. The ensemble mitigates these biases only modestly, if at all (e.g. it still overpredicts none and underpredicts mixed relative to humans). This highlights the need for correction techniques such as DSL, which we analyze in Section˜3.2.2.
(iv) Ensemble failure analysis
We begin by asking how often ensemble errors contain any useful signal at all. On the combined Test 1 and Test 2 data, the ensemble makes 1,112 errors. Of these, 747 (67.2%) occur on instances where all models are wrong, whereas only 365 (32.8%) occur when at least one model predicts the correct label. Thus, most ensemble failures occur on examples where none of the component models predicts the correct label.
The remaining 365 cases suggest that complementary signal exists in principle. To quantify how much performance could be recovered from this signal, we construct an oracle that selects a correct model prediction whenever at least one model is correct. The oracle reaches markedly higher macro-F1 than the ensemble–0.66 (0.87) on Test 1 and 0.45 (0.73) on Test 2, compared with 0.47 (0.70) and 0.27 (0.56), respectively. This indicates that although useful complementary signal is present in the model pool, simple ensembling fails to recover much of it in practice, which explains the ensemble’s substantially lower empirical F1.
A closer look at model overlap helps explain why this signal is difficult to exploit. Across the pooled Test 1 and Test 2 data, pairwise error Jaccard scores are consistently high (0.57-0.64; see Appendix˜D in the Appendix and Table˜4), indicating that models often fail on the same instances. Moreover, conditional on two models both being wrong, they predict the same wrong label in roughly half of cases (45-49%), which means that shared errors are often reinforced rather than corrected.
Taken together, these results suggest that the main limitation is limited error diversity across models on difficult instances. Section˜3.2 examines how the biases discussed in this section affect downstream trend estimates.
3.2 Improving Downstream Reliability
The preceding analyses identify two challenges for using LLM annotations in downstream inference. First, disagreement between human annotators does not necessarily reflect error, but may indicate intrinsic ambiguity and multiple defensible interpretations. Following work that discusses such cases under the umbrella term of label variation (Plank, 2022), we therefore move beyond hard labels and compare label distributions. Second, the error analysis shows that model predictions are systematically biased, with some labels over- or under-predicted relative to human annotations. Such bias can distort downstream analyses when predictions are aggregated into temporal or group-level trends. We address both issues in this section: we first assess the similarity of model and human label distributions (Section˜3.2.1), and then test whether debiasing with DSL improves recovery of human-derived trends (Section˜3.2.2).
3.2.1 Label distribution similarity
For each instance with at least three human annotations, we derive a human label distribution from the annotators’ votes. For the model ensemble, we analogously construct a label distribution from the votes of the three models (Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B). For individual models, we use sharp distributions that assign 100% probability mass on the predicted label. We compare these distributions using Jensen–Shannon divergence (JSD) and total variation distance (TV), where lower values indicate higher similarity (see Appendix˜E in the Appendix for definitions). The analysis covers both the high- and fine-grained levels. We compare model–human distances against three baselines: (i) human-human distance, averaged over ten random annotator splits; (ii) a randomized model baseline, with predictions shuffled across instances; and (iii) single-model baselines comparing each model to the human distributions.
Table˜5 in the Appendix shows that the ensemble (TV 0.49; JSD 0.39) is closer to human label distributions (TV 0.45; JSD 0.35) than any individual model (TV 0.57–0.59; JSD 0.48–0.50) and far closer than the randomized baseline (TV 0.75; JSD 0.69), with the same pattern observed at the high level. These results indicate that strong LLMs reproduce human-like patterns of disagreement. At the same time, the ensemble–human distance does not surpass the human–human baseline.
These results have the following implications: First, since model disagreement parallels human uncertainty, this variation should be preserved through soft labels rather than reduced to single hard predictions. Second, because models do not surpass human–human similarity and their errors remain correlated, these soft labels still contain structured bias. As a result, model predictions cannot be used directly without correcting for bias. We address this issue in the next section.
3.2.2 Bias Correction with DSL
We validate all trend estimators considered below using the complete set of 2,184 Migrant-related instances that were annotated by humans (hereafter the evaluation subset; see Table˜6 in the Appendix). Trends computed from these annotations serve as the reference. Because the evaluation subset is fully human-annotated, we simulate a partial-label setting for DSL evaluation. Specifically, we randomly retain human labels for only a subset of instances within each decade and treat the remaining labels as unobserved. We then apply DSL using the corresponding decade-specific sampling probabilities and compare the resulting trends with those computed from the full human annotations.
We compare the human-derived reference trends against several alternative estimators. We first evaluate uncorrected estimators based on raw model outputs: (i) single-model hard-label predictions from the strongest models (Llama-3.3-70B, Llama-4-Scout, Qwen-2.5, gpt-oss-120B, GPT-5), (ii) a majority-vote ensemble of the best open models (Llama-3.3-70B, Qwen-2.5, gpt-oss-120B), and (iii) soft-label distributions obtained from these three open models’ outputs. We then evaluate debiased estimators obtained by applying DSL to (iv) a single strong model (Llama-3.3-70B) and (v) the soft-label ensemble. For each estimator, we compute temporal trends by grouping instances by decade and computing the share of instances assigned to each label in each decade. For DSL-based approaches, we use the bias-corrected pseudo-outcomes $\tilde{Y}_{i}$ directly in these aggregations. We then quantify agreement with the reference model by computing Pearson and Spearman correlations, as well as root mean squared error (RMSE), between model-based and human-based time series. We compute correlations separately for each label and report the average across labels.
Table˜3 shows that the soft-label ensemble corrected with DSL recovers human-derived temporal trends more accurately than other approaches, including a single-model DSL estimator, at both the high-level and fine-grained levels.
Model / Estimator Pearson Spearman RMSE
High-level trends
Llama-3.3-70B 0.87 0.86 0.08
Llama-4-Scout 0.74 0.76 0.08
Qwen-2.5 0.72 0.71 0.11
gpt-oss-120B 0.84 0.80 0.07
GPT-5 0.82 0.82 0.07
Majority vote ensemble 0.88 0.87 0.08
DSL (Llama-3.3-70B) 0.85 0.83 0.06
DSL (soft-label ensemble) 0.89 0.91 0.06
Fine-grained trends
Llama-3.3-70B 0.64 0.58 0.09
Llama-4-Scout 0.58 0.58 0.07
Qwen-2.5 0.36 0.38 0.11
gpt-oss-120B 0.44 0.52 0.10
GPT-5 0.56 0.50 0.08
Majority vote ensemble 0.64 0.67 0.10
DSL (Llama-3.3-70B) 0.50 0.49 0.07
DSL (soft-label ensemble) 0.68 0.67 0.04
Table 3: Comparison with human-based decade-level trends across estimators. The table reports average Pearson and Spearman correlations and RMSE (reference: human majority vote).
At the high level, the strongest single models and ensemble variants already achieve high agreement with human trends, with Pearson and Spearman correlations above 0.80 and low RMSE. In this setting, DSL shows moderate but consistent improvements: DSL applied to the soft-label ensemble achieves the highest correlations (Pearson 0.89; Spearman 0.91) and the lowest RMSE (0.06).
(a) Human majority vote
(b) gpt-oss-120B
(c) DSL (soft labels)
Figure 8: High-level (anti-)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2). The grey shaded area marks the period 1930–1949, excluded due to limited data availability during and immediately after the NS dictatorship. Figure˜23 in the Appendix shows the same trends for the remaining configurations.
In contrast, fine-grained trends are substantially harder to approximate from raw model predictions alone. Single-model and ensemble estimators show lower correlations and higher error: notably, even models that achieve the highest macro F1 scores in Section˜3.1.1 show only moderate agreement with human-derived fine-grained trends (for example, gpt-oss-120B reaches a Pearson correlation of 0.44 with an RMSE of 0.10, and the majority-vote ensemble reaches 0.51 with an RMSE of 0.08). This indicates that strong instance-level performance does not guarantee accurate recovery of fine-grained temporal trends. Here, DSL has a clear effect, particularly in reducing error: DSL applied to soft-label ensembles achieves the lowest RMSE (0.04) and higher correlations (Pearson 0.68) than all uncorrected approaches.
(a) Human majority vote
(b) gpt-oss-120B
(c) DSL (soft labels)
Figure 9: Fine-grained (anti)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship. Figure˜24 in the Appendix shows the same trends for the remaining configurations.
Figure˜8 and Figure˜9 illustrate these results. First, DSL reduces overuse of the None category and correspondingly increases estimated trends for solidarity and anti-solidarity, for example for Llama-3.3-70B. This pattern is also reflected in Table˜8 in the Appendix, which compares per-label correlations with human-annotated decade-level trends for Llama-3.3-70B and soft-label DSL. There, DSL reduces the RMSE for None from 0.13 to 0.07.
Box 3.1: Label Shares Over Time Each speech belongs to a time period $t$ (e.g. a decade or a year) with $N_{t}$ speeches and receives a high-level or fine-grained label $\ell$. The estimated label share $\hat{p}_{t,\ell}$ in period $t$ is defined as follows: (a) Hard-label trends (e.g. human majority vote, gpt-oss-120B): $\hat{p}_{t,\ell}=\frac{N_{t,\ell}}{N_{t}}$, where $N_{t,\ell}$ denotes the number of speeches in period $t$ assigned label $\ell$. (b) Soft-label trends [^6] Reported in Figure˜24 in the Appendix.: $\hat{p}^{\text{soft}}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}Y_{i,\ell}$, where $Y_{i,\ell}$ denotes the probability of label $\ell$ for speech $i$, obtained from the distribution of annotator or model labels. (c) DSL (soft-label) trends: $\hat{p}^{DSL}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
At the fine-grained level, raw model outputs, including gpt-oss-120B, Llama-3.3-70B, Qwen-2.5, and their ensemble variants, show unstable trends, with dominance shifting between group-based, exchange-based, and compassionate categories across decades. DSL achieves more stable trends that follows with the human reference more closely, particularly within solidarity. Most notably, all raw model trends overemphasize group-based solidarity at the expense of compassionate solidarity. DSL corrects this imbalance and restores the dominance of compassionate solidarity observed in human annotations. Table˜8 in the Appendix shows that these corrections correspond to substantial error reductions, with RMSE for group-based solidarity decreasing from 0.20 to 0.06 and for compassionate solidarity from 0.20 to 0.08. A similar pattern is observed for fine-grained anti-solidarity. In raw model outputs, anti-solidarity is often limited to a single dominant subtype or disappears entirely; for example, empathic and compassionate anti-solidarity are absent from trends derived from Llama-3.3-70B predictions. DSL restores these subtypes and recovers differentiation between group-based and exchange-based anti-solidarity.
In sum, these results suggest that combining soft-label model outputs with DSL is a viable strategy for studying long-run trends when full human annotation is not feasible.
4 Temporal Trends in Solidarity Towards Migrants
In this section, we analyze temporal trends in the full Migrant dataset (63k instances) using DSL-corrected soft labels. We relate the resulting patterns to findings from political science, sociology, and history on migration to assess whether the temporal trends recovered by the DSL-corrected estimates are consistent with established interpretations.
For these analyses, DSL uses the decade-stratified human annotation sample of 2,184 Migrant instances (3.5% of the full corpus) to estimate bias-corrected label distributions for the full corpus of 63k speeches. Since human annotations were obtained under a decade-stratified sampling design (see Section˜2.2), we estimate DSL with decade covariates and decade-specific labeling probabilities $\pi_{i}$, and reuse the resulting pseudo-outcomes $\tilde{Y}_{i}$ for all analyses that condition only on time, including the sub-period analyses introduced in Section˜4.1.2. For analyses that condition on additional variables, such as party or keyword, we re-estimate DSL on the relevant analysis sample, since the prediction function $\hat{g}(Q_{i},X_{i})$ changes when additional covariates are included. In these analyses, party or keyword indicators enter the DSL prediction step as additional covariates alongside decade indicators, while labeling probabilities $\pi_{i}$ remain defined at the decade level, consistent with the annotation design.
We first present the DSL-corrected trends for the full dataset, summarizing the broad long-term pattern in parliamentary discourse about migrants and briefly assessing whether the aggregate trends are robust to variation in keyword composition. We then identify two periods with particularly notable shifts in solidarity–the post-war period and the recent past–which we analyze in more detail in Section˜4.2 and Section˜4.3.
4.1 Overall Solidarity Trends with DSL, 1867-2025
Figure 10: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the full 63K dataset. Figure˜10(a) shows the high-level (anti-)solidarity trends, while Figure˜10(b) shows distribution of fine-grained (anti-)solidarity subtypes over time. The period 1930–1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
The overall trends from 1867 to 2025 are shown in Figure˜10. Across most of the observable period, and especially from the beginning of the 20th century through the 2010s, solidarity exceeds anti-solidarity in parliamentary discourse about migrants. This pattern is driven mainly by compassionate solidarity, with group-based solidarity playing a less prominent but persistent role. Solidarity peaks in the postwar decades, while anti-solidarity remains comparatively low until the late twentieth century. From the 1980s onward, anti-solidarity rises (especially its group-based, compassionate, and exchange-based subtypes), with the sharpest jump occurring between the 2010s and 2020s.
Before turning to specific historical periods, we assess whether these aggregate trends are robust to changes in keyword composition over time.
4.1.1 Robustness to Keyword Composition
As Figure˜21 and Figure˜22 in the Appendix show, individual keywords vary widely in temporal distribution in the balance of solidarity vs. anti-solidarity associated with them. To ensure our long-term trends are not driven by such effects, we conduct two robustness tests.
First, we conduct a stability test by randomly sampling 200 subsets of at least five keyword groups (grouping inflectional variants). For each subset, we re-estimate the high-level DSL trends (as defined in Section˜3.2.2) and compare the resulting time series to the full-sample trend using Pearson’s correlation coefficient. The trends remain broadly stable, with mean correlations of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; for more detail on keywords and the stability test, see Appendix˜A in the Appendix. Second, we restrict the corpus to the two most frequent and temporally widespread terms, Ausländer (en.: foreigners) (19.7k mentions) and Flüchtlinge (en.: refugees) (12k), and re-estimate DSL on this reduced sample. Both checks confirm that overall solidarity and anti-solidarity trends are preserved both under this restricted vocabulary (see Figure˜33) and across different keyword selections.
4.1.2 Notable Trends in 1867-2025
Inspecting Figure˜10, we identify three periods with notable shifts in solidarity: 1890-1900s, the post-war period, and the recent period after 2010. We note that the 1930s and 1940s are outliers, as the parliament was swiftly sidelined by the Nazi regime and the corpus contains only 109 available instances from 1933 to 1949. Hence, we exclude these data points from further analysis.
The first of these, 1890-1900, is difficult to interpret with confidence. In this early part of the series, even the strongest model in our ensemble performs worse than in later decades in terms of macro-F1, and agreement between human annotators is also lower (see Figure˜18). Although DSL corrects the resulting estimates, we do not consider this early trend sufficiently reliable for detailed substantive analysis. This period also contains fewer instances than the post-war and recent periods that we analyze in detail below (see Figure˜2).
For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look.
4.2 Germany after the Second World War, 1949-1957
The Second World War triggered the largest wave of forced migration in European history, with tens of millions of people displaced or resettled across Central and Eastern Europe (Wessel, 2017). The affected populations were heterogeneous—refugees, displaced persons (DPs), forced laborers—and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans. For those who settled in postwar Germany, integration posed political and societal challenges shaped by regional disparities, economic hardship, and tensions with locals (Chevalier et al., 2024).
This subsection examines the early postwar period (1949-1957), when the Federal Republic institutionalized solidarity as part of nation-building. This period is analytically important because solidarity became a tool of governance, embedded in laws like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy (Kittel, 2020). These measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups. Since solidarity was neither static nor uniform, shaped by diverse migrant groups, competing ideologies, and contested notions of deservingness (Ther, 1996), we employ Thijssen’s typology of solidarity (Thijssen, 2012) to guide our analysis. By distinguishing identity-, need-, and contribution-based forms, we analyze how solidarity was constructed and sustained amid institutional strain.
4.2.1 Overall solidarity distribution
During 1949–1957, solidarity dominated parliamentary discourse on postwar population movements. As Figure˜11 shows, [^7] We report (DSL-adjusted) counts rather than proportions (see definition (d) in Section 4.2.1) because counts make spikes visible whenever migration-related laws are debated and enable comparison of debate volume across months. mentions of solidarity consistently exceeded anti-solidarity and mixed statements, at times by a factor of ten.
Figure 11: Estimated counts of (anti-)solidarity mentions in Bundestag debates, 1949–1957 (DSL-adjusted, see definition (d) in Section˜4.2.1). LAG == Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG == Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG == Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
Mentions of solidarity closely track key legislative moments. A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced. A second cluster in early 1952 reflects debates on the LAG, preceding its adoption in May (see Example 15 in Table˜19 in the Appendix). Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified integration measures (Example 15 ). Later peaks in September 1954 and June 1955 reflect debate over LAG implementation and budgets. Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
Box 4.1: Estimated Monthly Counts of (Anti-)Solidarity Speeches are grouped by month $t$, with $N_{t}$ speeches in month $t$. We define the estimated number of mentions of label $\ell$ in month $t$ as: (d) DSL-adjusted monthly counts: $\hat{C}^{DSL}_{t,\ell}=\sum_{i=1}^{N_{t}}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
Figure 12: Distribution of (anti-)solidarity subtypes, 1949–1957 (DSL-adjusted expected counts, see definition (d) in Section˜4.2.1). Inset shows the evolving composition of solidarity subtypes, normalized within each six-month period as shares of all solidarity framing (see definition (c) in Section˜3.2.2).
The prominence of solidarity in this period can be explained by the fact that debates used it not only to justify compensation but also to manage postwar social and economic fragmentation and competing claims to belonging (Hughes, 1999). The distribution of all four subtypes across key legislative moments (Figure˜12) reflects this mix of solidarity rationales. Compassionate and group-based frames dominate throughout the period (see inset of Figure˜12). By emphasizing shared national belonging (group-based) and the suffering of affected populations (compassionate), these frames construct a victim-centered Schicksalsgemeinschaft (en.: community of fate), uniting Germans through shared loss (Gengler, 2020) (see Example 15in Table˜19 in the Appendix [^8] Example labels are based on the dominant DSL subtype, that is, the subtype with the highest DSL-adjusted score. This is useful for illustrating which subtype is most strongly reflected in a given speech within the DSL-based analysis; however, since DSL is designed for aggregate inference, these labels do not guarantee corrected true labels for individual instances.). This coexistence of the two subtypes is consistent with Thijssen (2012), who argues that compassionate and group-based solidarity form a dialectical pair: identification with a collective (?we?) generates obligations toward its members, while emphasizing their suffering reinforces that collective bond.
Additionally, a small but persistent share of exchange-based solidarity appears throughout the period. It reflects debates over efficient allocation and the reintegration of displaced populations into labor and housing markets, with measures such as the LAG framed as tools to rebuild productive capacity (Examples 15 and 18).
These patterns show that solidarity in parliamentary debates mediated the challenges of postwar reconstruction through key legislative reforms. However, since these reforms differentiated between migrant groups, we examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in this period.
4.2.2 Solidarity Frames Across Keywords in 1949-1957
The analysis includes migrant-related keywords for which the DSL-adjusted expected number of stanced instances (i.e., solidarity, anti-solidarity, or mixed rather than none) is at least 150. Keywords dominated by none or with low DSL-adjusted expected counts (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants)) are excluded; see Table˜9 in the Appendix for full distributions. Figure˜13 shows the distribution of (anti-)solidarity frames across these terms.
Vertriebene (en.: expellees), a category reserved primarily for Volkszugehörige (en.: ethnic Germans), are most often associated with compassionate solidarity (about half of all expressions with a stance), and also with group-based solidarity (nearly 25%). This reflects their dual position: as key representatives of German victimhood during and after the World War II (Moeller, 2001), they were both central to West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice.
Box 4.2: Distribution of (Anti-)Solidarity Subtypes Across Groups Speeches are grouped by a categorical variable $g$ (a keyword or a political party). $I_{g}$ is the set of speeches belonging to group $g$. The share of subtype $\ell$ within group $g$ is defined as: (e) Subtype shares within groups: $\hat{s}_{g,\ell}=\frac{\hat{M}_{g,\ell}}{\hat{M}_{g}}$, where $\hat{M}_{g,\ell}=\sum_{i\in I_{g}}\tilde{Y}_{i,\ell}$ is the DSL-adjusted mass of subtype $\ell$ in group $g$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2. $\hat{M}_{g}=\sum_{\ell\in L}\hat{M}_{g,\ell}$ is the total mass of group $g$ across all labels $L$ (four solidarity subtypes, four anti-solidarity subtypes, and mixed stance). Shares sum to one within each group $g$.
Still, around 10% of stance expressions toward Heimatvertriebene were anti-solidarity, particularly group-based (solidarity limited to ?us?, the in-group), reflecting concerns that expellees strained local infrastructures or were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment (Kossert, 2008). Perceptions of injustice grew with the uneven distribution of expellees across West German states (Braun and Mahmoud, 2014), especially where their prominence in institutions and politics sparked fears of Überfremdung (en.: overforeignization) and unequal burden-sharing (Levy and Sznaider, 2005) (see Example 15 in Table˜19).
Figure 13: Relative distribution of (anti-)solidarity subtypes across selected migrant-related keywords, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2). Bars show each subtype’s share within a keyword group; shares sum to 100%. Full label distributions appear in Table˜9 in the Appendix.
Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are also primarily framed through compassionate solidarity (50–72%), followed by group-based solidarity (15%). The label Flüchtlinge began as a broad catch-all in the late 1940s but, even after the 1953 BVFG differentiated it into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler, it retained connotations of marginality and foreignness (Nachum and Schaefer, 2018). This helps explain a more notable share of anti-solidarity (around 10%). Sowjetzonenflüchtlinge, framed as Cold War refugees from communist oppression, were less integrated into the nation-building narrative than Vertriebene, prompting more compassionate than group-based solidarity (Nachum and Schaefer, 2018; Böke, 1996) (see Example 16).
Among the migrant categories, Ausländer (en.: foreigners) shows one of the highest relative shares of anti-solidarity. A discursively constructed ?other?, including DPs [^9] With the 1951 HAuslG, authority over the remaining DPs was transferred from Allied administration to the Federal Republic, and they were reclassified as heimatlose Ausländer (en.: stateless foreigners), a noncitizen status that granted them most civil rights without granting citizenship (Holian, 2011)., Eastern European border crossers, wartime laborers, and early foreign workers (Smith, 1994), they were contrasted with groups like Vertriebene or Einheimische and framed as non-belonging (evident in group-based anti-solidarity share), economically burdensome (exchange-based anti-solidarity), or even criminalized (Holian and Wilhelm, 2017) (see Example 16). Yet solidarity, especially group-based, still exceeded anti-solidarity, partly due to the category’s fluidity: it sometimes included ethnic Germans not yet formally recognized as such in the early postwar years (Ruhkopf, 2023). Speakers reinforced this dynamic by linking DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community (Ruhkopf, 2023) (Example 16).
The second-highest subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy (Ruhkopf, 2023), for example through the 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure (Holian and Wilhelm, 2017) (Example 17). However, solidarity also extended to non-ethnic German foreigners, particularly in CDU and SPD statements, showing that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 16 and 17).
In sum, compassionate framing dominated across most categories, reflecting a broader discourse of shared suffering in which displaced populations were positioned as victims of the war and its consequences. Nevertheless, solidarity was still shaped by perceived ?deservingness? and ethnic identity. While Vertriebene were more firmly incorporated into the national community as ethnic Germans, others, especially Ausländer, received comparatively less support and conditional inclusion (Nachum, 2013). These differences reflected both group characteristics and the ideological strategies of political parties. The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.
4.2.3 Distribution across parties
Figure˜14 shows party distributions of solidarity and anti-solidarity in 1949–1957. Solidarity dominated overall, but parties framed it differently, shaped by ideology and strategy.
The main parties – governing CDU/CSU and FDP, as well as opposition SPD – supported expellee and refugee integration through legal and institutional measures, primarily expressing compassionate or group-based solidarity (see Figure˜14). The CDU/CSU led in proposing the LAG, framing solidarity as a national duty toward fellow ethnic Germans, and even included expellee representatives in the party ranks (Hughes, 1999). The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers (Connor, 2018), resulting in occasional anti-solidarity statements. The FDP, while formally aligned with CDU/CSU and SPD on refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation (Connor, 2018) (see Example 17 in Table˜19).
Figure 14: Distribution of (anti-)solidarity subtypes across Bundestag parties, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2). Bars show each subtype’s share within a party, summing to 100%. The analysis includes parties with at least 150 DSL-adjusted stanced statements. Full label distributions appear in Table˜10 in the Appendix.
The expellee advocacy party GB/BHE [^10] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights). The DP (Deutsche Partei (en.: German Party)) also addressed expellee demands, especially in Lower Saxony because of their electoral weight there, but it began as a Heimatpartei (en.: ?homeland party?) rather than a refugee party (Aschoff, 2014). This is visible in the party’s share of group-based anti-solidarity, reflected in its defense of local interests (Example 19). positioned itself as a political representative of displaced Germans. Following multiple successes in state elections, GB/BHE, founded by and for expellees, entered the federal coalition in 1953 and pressured other parties to address expellee demands (Connor, 2018). It expressed compassionate and group-based solidarity, portraying expellees as full members of the national community entitled to compensation and integration (see Example 17). Additionally, the GB/BHE shows the highest share of empathic solidarity (15%), consistent with its role as an expellee party defining expellees as a distinct community. Over time, CDU and SPD managed to absorb major expellee demands, which made GB/BHE less successful (Connor, 2018; Kittel, 2020).
Regionalist parties such as WAV and BP [^11] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized compassionate solidarity, but restricted it to ethnic Germans. The short-lived WAV appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories (Kranenpohl, 1998) (Example 18). The BP, responding to Bavaria’s disproportionate refugee burden, called for invoked collective responsibility, fairer redistribution and favored groups like the Sudeten Germans (Braun and Dwenger, 2020), which corresponds to its high share of group-based anti-solidarity (see Example 18). Both parties also expressed the highest levels of mixed stance, either by excluding non-German migrants (WAV) (Connor, 2018) or opposing specific groups such as Silesians (BP).
The KPD [^12] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) also addressed displaced populations, framing them through socialist integration and emphasizing compassionate and group-based solidarity, though in an assimilationist rather than inclusive sense. Avoiding the label Vertriebene in line with Soviet policy, the party instead spoke of ?refugees?, portraying them as victims of fascism and capitalism (Heidemeyer, 2007; Connor, 2018).
Parties like KPD, WAV, and BP did not stay in parliament for long, but nonetheless shaped early postwar debatesand helped define solidarity and its boundaries of deservingness. As suggested by Thijssen and Verheyen (2022), such framings were conditioned by the period’s political and social challenges and reflected broader efforts to maintain or renegotiate societal cohesion.
4.3 Recent German history 2009-2025
Between 2009 and 2015, annual migration to Germany rose from just under 800,000 to more than 2 million people, peaking in 2015 in the context of the Syrian war (Gessler and Hunger, 2023). This period has since been associated with the label ?migration crisis? (Gessler and Hunger, 2023; Fröhlich, 2025). Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017. Another wave of migration followed the 2022 Russian invasion of Ukraine (Federal Office for Migration and Refugees, 2025). Alongside these events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
4.3.1 Overall solidarity distribution
(a) Solidarity and anti-solidarity distribution
(b) Solidarity and anti-solidarity frames distribution
Figure 15: (Anti-)solidarity trends for the Migrant category, 2009–2025. Figure˜15(a): Yearly DSL-adjusted shares of solidarity, anti-solidarity, and mixed stance (see definition (c) in Section˜3.2.2). The none category is included in the normalization but omitted from the visualization. Figure˜15(b): Yearly DSL-adjusted shares of solidarity (left) and anti-solidarity (right) subtypes (see definition (c) in Section˜3.2.2). Within each year, subtype shares are renormalized across the eight solidarity and anti-solidarity categories (excluding mixed and none), summing to 100%.
Figure˜15 shows (anti-)solidarity trends from 2009–2025 (DSL-adjusted). Solidarity declined from 2014 to 2019, increased modestly between 2020 and 2022, and then saw a slight decline after 2022. This decrease in solidarity is associated with an almost corresponding increase in anti-solidarity. In terms of (anti-)solidarity subtypes, compassionate solidarity peaks around 2014–2016, drops sharply, briefly rebounds in 2021, [^13] This rebound coincides with the COVID-19 context, rising asylum applications, humanitarian responses to the Taliban takeover in Afghanistan, and the formation of the SPD–Greens–FDP coalition government (Brandt et al., 2024). and declines again. The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity. In terms of anti-solidarity, group-based anti-solidarity rose strongly after 2017, accounting for over one third of all stanced migrant-related statements. Compassionate anti-solidarity spikes around 2020 and then declines after 2022.
Overall, the period since 2019 marks the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements in the German parliament. This key finding is visible both in the DSL-adjusted estimates but also in human annotations (Figure˜2). We next contextualize it with recent Germany history and a deeper look into party distribution.
The 2014 spike in compassionate solidarity coincided with heightened public support for Syrian refugees: 15.4% of German survey respondents named solidarity their most important value in 2014, up from 12.3% in 2013 (De Leur et al., 2023, p. 38). Refugees were commonly framed as needing help, motivating calls for a ?German welcome culture? (Gebauer, 2023) (see Example 20 in Table˜23 in the Appendix). By early 2016, a counter-narrative emerged portraying refugees as overburdening German capacities and as contributing to a ?refugee crisis? (Gebauer, 2023) (see Example 20). During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests, shifted further towards the right of the political spectrum, and entered parliament in 2017 after major electoral gains (Frey, 2020; Gebauer, 2023). Regarding the broader European context, De Leur et al. (2023) diagnose a post-2016 ?solidarity fatigue?, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level. These counter-narratives match several notions of anti-solidarity according to Thijssen and Verheyen (2022): refugees as an outgroup competing with German population (group-based, see Example 22), as undeserving of compassion (compassionate, see Example 23), and as straining resources without contributing (exchange-based, see Example 21). We also note a general shift towards right-wing sentiments in the general German population (Zick et al., 2023), which may be associated with heightened anti-solidarity.
More recently, the Russian invasion of Ukraine and the resulting arrival of Ukrainian refugees in Germany, marking the largest net migration since 1950 (Atanisev et al., 2024), triggered another, though less prominent, spike in public solidarity (Weber et al., 2024a). In our data, solidarity increases slightly in 2022, followed by a continued decline. Bojarskich et al. (2025) identified several factors promoting anti-solidarity towards Ukraine among the German public, such as conspiracy mentality (i.e.  receptivity to contrarian narratives and disinformation emphasizing the risks of supporting Ukraine) and support for populist parties, but also fears about economic and societal consequences, indicating group-based (competition between groups) or exchange-based (insufficient contribution) anti-solidarity (Example 21). Anti-Ukrainian framing may also have been reinforced by the AfD’s pro-Russian stance, including opposition to sanctions and military aid (Arzheimer, 2023). In contrast to the postwar period, (anti-)solidarity shifts between 2009 and 2025 appear less tied to specific legislation, likely because much of German migration policy is negotiated at the EU level (Geddes et al., 2020). Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics. As such, we focus less on legislation and more on party-specific trends.
Figure 16: Raw yearly counts of parliamentary statements relating to migrants by selected political parties, 2009–2025.
4.3.2 Distribution across parties
(a) AfD
(b) CDU/CSU
(c) SPD
(d) FDP
(e) Die Linke
(f) Die Grünen
Figure 17: DSL-adjusted yearly shares of (anti-)solidarity subtypes across Bundestag parties, 2009–2025 (see definition (c) in Section˜3.2.2). Within each year and party, subtype shares are renormalized across the eight solidarity and anti-solidarity categories so that they sum to 100%. Note: the AfD entered the Bundestag only in 2017, while the FDP was absent from 2013–2017; accordingly, no observations appear for these periods.
Figure˜16 shows the raw numbers of migration-related statements per party and year, Figure˜17 shows the fine-grained solidarity trends.
The right-wing AfD
expresses almost no solidarity, with anti-solidarity stable over time and dominated by group-based frames, fluctuating around 50% throughout 2017-2025, although a notable number of instances with compassionate and exchange-based anti-solidarity (ca. 25%) occurs, with the two frames alternating across years. Though one of the smallest factions in parliament, the AfD accounts for most anti-solidarity instances overall (53.3% over the whole period). The sharp rise of anti-solidarity after 2017 (Figure˜15(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure˜16). Many AfD speeches, across topics, link migrants to illegal border crossings, criminality, or additional burdens on the German society, without necessarily addressing the actual topic of legislative debate (see Example 21 in Table˜23 in the Appendix ).
The center-right CDU/CSU and liberal FDP
changed the most in our data. In 2009-2015, most statements expressed solidarity, with both parties focusing mostly on compassionate solidarity. However, the main difference between the two parties lies in their framing: the CDU/CSU emphasizes group-based solidarity, whereas the FDP foregrounds exchange-based solidarity, for instance by stressing the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: Skilled Labor Migration Law) in 2023 (see Example 21). From the late 2010s onward, both parties shift toward anti-solidarity, most visibly in the CDU/CSU, where group-based anti-solidarity and compassionate anti-solidarity frames rise after the 2015 refugee crisis, decline during the Ukrainian refugee inflow in 2022, and resurge by 2025, reaching 33% and 24% of all frames, respectively. For the FDP, a comparable shift appears by 2025, where exchange-based, group-based, and compassionate anti-solidarity account for 16%, 12%, and 12%, respectively. When inspecting the instances, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, for which group-based anti-solidarity has the highest DSL estimate among the anti-solidarity subtypes (see Example 22) and portraying migration as misuse of welfare benefits, corresponding to exchange-based anti-solidarity (see Example 22).
The shift toward compassionate and group-based anti-solidarity over the last three years is insufficiently explained by recent literature. One possibility is that CDU/CSU and FDP adopted populist rhetoric from AfD after 2017 (Esguerra et al., 2023). However, Lewandowsky et al. (2022) and Kesting et al. (2018) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality. Notably, the first shift occurred even before the AfD entered parliament. Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population (Zick et al., 2023), appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD. Overall, our findings encourage further research into this striking and rapid trend.
The center-left SPD
shows predominantly compassionate solidarity, with a spike in 2020, and is closely followed by group-based solidarity, which gains prominence post-2019, and almost taking over in 2025. This is consistent with the party’s humanitarian commitments and stronger emphasis on integration (Orhan, 2025). Additionally, exchange-based solidarity is present, at times reaching around 20%. In line with this perspective, SPD has repeatedly framed Germany as an Einwanderungsland (en.: country of immigration) and supported immigration policies aimed at attracting qualified workers to meet labor market demands (Sozialdemokratische Partei Deutschlands, 2021). However, there are also instances where SPD speakers express anti-solidarity by referring positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 22).
The left Green and Linke parties
only display very few instances classified as anti-solidarity. Overall, since 2021, Die Grünen moved from compassionate solidarity to group-based appeals, stressing integration and rights. In the 2024/25 campaign, and in the 2024/25 campaign strongly emphasized exchange-based arguments about migrants’ economic and social contributions (Bündnis 90/Die Grünen, 2025) (see Example 23). Die Linke likewise emphasized compassionate solidarity but since 2021 started increasingly framing migration as integral to society, advancing a detailed pro-immigration agenda and strengthening group-based solidarity. In 2024/25, the party continued to oppose stereotyping and position themselves against exclusionary discourses (Die Linke, 2021, 2025) (see Example 23). A special case is the BSW, which split from Linke party in 2024, partially due to explicit anti-migration positions. However, our data set contains only very few instances from this party (38 statements), so we cannot provide a reliable quantitative analysis.
5 Discussion and Conclusion
In this paper, we have evaluated LLMs on the complex, theory-driven task of annotating (anti-)solidarity in German parliamentary speeches and assessed whether their outputs support valid downstream social-scientific inference. We then use a validated and statistically corrected annotation pipeline to derive new insights into solidarity and anti-solidarity toward migrants in the postwar period as well as the recent past. In more detail, we provide both methodological and social scientific insights:
Methodological perspective
Our results show that strong instruction-tuned LLMs can reach human agreement in annotating solidarity frames in parliamentary debates, where GPT-5 performs best overall, while open models such as gpt-oss-120B and Llama-3.3-70B provide competitive and more cost-effective alternatives. At the same time, performance varies systematically by target group, historical period, prompting strategy, and model family, with older data and fine-grained labels proving substantially harder. Generally, existing research presents mixed evidence on the reliability of LLMs for annotation tasks: some studies mention limitations in handling subjective or expert-defined taxonomies (Ziems et al., 2024) or find that smaller fine-tuned models are more reliable in political science applications (Wang et al., 2024), while others show that LLMs can substitute for or outperform human coders (Heseltine and Clemm von Hohenberg, 2024; Törnberg, 2024). Our findings are in line with a more cautious interpretation: LLMs are useful annotation tools, but they should not be viewed as neutral coders, nor should their outputs be treated as ?ground truth? data; rather, they are ?measurement tools that require rigorous validation, calibration, and documentation? (Baumann et al., 2025). This idea is reinforced by our error analysis, which shows that even the strongest models in our study make systematic mistakes. For example, Llama-3.3-70B occasionally misclassifies anti-solidarity, mistaking criticism of anti-migrant rhetoric for anti-solidarity itself, and all models show a general tendency to overpredict the None label. More broadly, they underestimate Mixed cases and compress fine-grained subtype distinctions into broader categories, especially group-based subtypes of (anti-)solidarity. These errors become more severe on historical instances, where less explicit forms of solidarity are often missed. More importantly, these errors are not independent across models: they are often correlated, which limits the benefit of ensemble methods. This matters for downstream analysis because small but systematic misclassifications can accumulate into biased temporal trends.
Recent work highlights this problem: high agreement with human annotators can coexist with systematic misclassifications, meaning that benchmark performance alone does not guarantee valid downstream inference (Bavaresco et al., 2025; Gligoric et al., 2025). Even highly accurate LLM-based surrogates can therefore result in misleading substantive conclusions when used directly in downstream analyses, because doing so ignores measurement error between gold-standard and surrogate labels (Egami et al., 2023). Consistent with this, our study finds that DSL-corrected soft-label outputs best recover human-derived trends, especially in fine-grained categories. This suggests that large-scale LLM annotation, combined with a limited set of expert annotations (around 2k in our case), is a viable strategy for downstream inference. The benefits are greatest when soft labels are aggregated across multiple strong models and corrected with DSL, rather than treated as hard predictions from a single model. Recent benchmarking further suggests that DSL performs especially well in low-sample settings, often outperforming alternative correction methods (de Pieuchon et al., 2025). From a resource-allocation perspective, this pipeline combines expensive expert labels with low-cost model labels in a way that improves inference. In our project, expert annotation cost over €29,000 for 3,500 instances, whereas model annotation scaled much more cheaply, whether via open-weight models such as gpt-oss-120B and Llama-3.3-70B on in-house NVIDIA A40 GPUs or via GPT-4, which annotated one third of the dataset for about €500. A smaller set of costly human labels provides the reference basis for validation and correction, while a much larger set of low-cost model labels extends coverage to corpus scale.
Social science perspective
Beyond methodological contributions, our study shows how an LLM-based pipeline can support theoretically informed quantitative analysis through two case studies: one historical (1949–1957) and one contemporary (2009–2025).
For the early postwar period (1949–1957), we showed how solidarity was shaped and institutionalized in national reconstruction. While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insights. First, it moves beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity and shifting rhetorical strategies across migrant groups. Second, our approach integrates migration histories that are often treated separately (expellees, displaced persons, and refugees) into a shared analytical framework, enabling direct comparison of how solidarity was articulated. This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis. This may reflect solidarity fatigue (De Leur et al., 2023) or a larger societal right-wing shift (Zick et al., 2023) on migration. While stricter migration stances across Europe have been studied, our study is the first to quantify this shift in German parliamentary debates. Our data also revealed that the shift is not only explained by the right-wing AfD party entering parliament, but it also visible in the conservative CDU/CSU party and the liberal FDP party, motivating further research. This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives (Aktas, 2024), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists (Bayerlein, 2021). Comparative surveys in Europe likewise show anti-immigrant attitudes remain a central driver of right-wing populist support (Messing and Ságvári, 2021).
Taken together, our findings address a broader question in CSS: whether LLMs can support valid, theory-driven analysis of complex political concepts at scale. Our answer is yes, but only under clear conditions. LLMs are not neutral replacements for expert coders, and their outputs cannot be used uncritically as if they were ground-truth labels, particularly for theory-grounded annotation tasks. Yet when combined with human reference annotation, systematic validation, error analysis, and DSL-based correction, they can support serious social-scientific research by revealing patterns at a scale that would otherwise be accessible only through qualitative analysis. In this sense, our pipeline bridges ?distant reading? approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded and systematic interpretation of parliamentary speech while retaining full corpus coverage.
6 Limitations and Future Work
While our analysis was performed carefully to avoid methodological pitfalls, we nonetheless acknowledge remaining limitations of our study.
Human Coded Reference
Our evaluation and DSL correction rely on human annotations as the best available expert-coded reference. At the same time, some degree of disagreement remains, especially at the fine-grained level, even after extensive training and refinement of the annotation manual. This reflects the intrinsic difficulty of interpreting theoretically informed categories in parliamentary speech and the possibility of reasonable disagreement (e.g.  over whether competition over resources is group- or exchange-based anti-solidarity, or whether a text is better read as compassionate or group-based solidarity.). DSL improves correlation with human annotations, but it cannot remove uncertainty where the reference itself remains interpretive.
Limits of Correction with DSL
DSL improves recovery of human-derived trends, especially relative to raw hard-label predictions, but it does not remove all sources of error. Its performance depends on the quality of the surrogate model outputs, the coverage of the human-coded subset, and the sampling design. In settings with rare labels or sparse historical coverage, correction may remain unstable.
Temporal Reliability
Both model performance and human agreement fluctuate over time (e.g.  the 1960s for several models; Figure˜18), reflecting the challenge of historical language. Although our correction framework mitigates part of this problem, historical variation remains a challenge for both annotation and downstream inference. Still, because our focus is on broad temporal trends rather than individual years, these fluctuations are less consequential for our main substantive conclusions.
(Anti-)Solidarity towards Women
Similar to migrants, one could analyze (anti-)solidarity trends toward women. Diabaté et al. (2023) notes a leftward shift in Germany on gender issues, potentially contrasting with our migration findings. Due to space and time constraints, we do not pursue this here, but consider it a promising avenue for future work.
Political Bias
LLMs compared in this paper may be subject to political bias as investigated by several prior works (Motoki et al., 2024; Rettenberger et al., 2025), with left-wing political bias potentially leading to more anti-solidarity being annotated. Yet our error analysis (Figure˜6) shows no systematic skew toward anti-solidarity; models more often mislabel instances as ?None?. Annotator bias towards certain labels is also possible but would need to be uniform across annotators, which we do not observe. Even if present, bias would affect only overall levels of (anti-)solidarity, not the differences and trends central to our analysis.
Appendix
Appendix A List of Keywords
For Frau (en.: woman), we use Frauen (and Frau), Mütter (and Mutter, Müttern), Großmutter, Hausfrauen (and Hausfrau), Ehefrauen (and Ehefrau), Mädchen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
Since Frau is the German word for woman but also for Mrs./Ms., we only include occurrences of the word Frau that are not followed by a capitalized word (which are probably surnames).
For Migrant (en.: migrant) we use Flüchtlinge (and Flüchtling, Flüchtlingen), Ausländer (and Ausländern, Ausländerinnen), Asylsuchende (and Asylsuchende, Asylsuchenden), Asylbewerber (and Asylbewerbern), Heimatvertriebene (and Heimatvertriebenen), Vertriebene (and Vertriebenen, Vertriebener), Aussiedler (and Aussiedlern), Einwanderer (and Einwanderung), Ansiedler (and Ansiedlern), Zuwanderer (and Zuwanderern, Zuwanderung), Migranten (and Migrantinnen, Migration), Sowjetzonenflüchtlinge, Bürgerkriegsflüchtlinge, Kriegsflüchtlinge, Emigranten, and Immigranten.
For robustness, we conducted a stability test by randomly sampling 200 subsets of at least 5 keyword groups (grouping inflectional variants) each, ensuring that each instance covered at least 10% of the dataset and spanned at least 75% of the timeline (such that enough data is present to create the plots). For each subset, we re-estimated the high-level DSL trends (as defined in Section˜3.2.2) and compared the resulting time series to the full-sample trend using Pearson’s correlation coefficient. The correlations remain high on average, with mean values of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; the 25% quantiles are 0.89, 0.85, and 0.75, respectively. This indicates that the main aggregate trends are broadly robust to keyword composition, although robustness is weaker for mixed.
For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred. The trends are shown in the Figure˜21 in the Appendix.
Appendix B Parties
For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag. For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by Porada (2019).
Each new intervention is marked in the XML by a <redner> element containing both the speaker’s name and their party. During preprocessing (XML parsing, transcript cleaning, and sentence segmentation), we convert these into standardized <SPEAKER> tags and extract the associated party information. During extraction of migrant- or women-related sentences, the most recent <SPEAKER> marker (with party) is assigned to each sentence until a new one appears.
List of the parties included in the dataset, along with the variations of their names or abbreviations as they have been recorded: AfD (Alternative for Germany); Die Linke (The Left) with variations such as PDS, Gruppe der PDS; Bündnis 90/Die Grünen (Alliance 90/The Greens); CDU/CSU (Christian Democratic Union/Christian Social Union); SPD (Social Democratic Party of Germany); FDP (Free Democratic Party); DP (German Party) with variations such as DP/DPB, DP/FVP, FVP; GB/BHE (All-German Bloc/League of Expellees and Deprived of Rights); KPD (Communist Party of Germany); BP (Bavarian Party); WAV (Economic Reconstruction Union).
Appendix C Models Inference and Training Details
Hardware and Inference Settings.
Inference was conducted on mixed CPU and GPU environments. Closed-source models (gpt-4-1106-preview, gpt-5-chat) were accessed via API.
Open-weight models were run locally:
  • gpt-oss-20B on two NVIDIA A40 (80 GB) GPUs in bfloat16 precision;
  • Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it in full precision on a single A40 GPU;
  • Mistral-Large-Instruct-2407 on a rented NVIDIA A100 in 4-bit quantization to accommodate memory requirements;
  • Llama-3.3-70B-Instruct in 4-bit quantization and Qwen2.5-72B-Instruct on two A40 GPUs in 8-bit quantization;
  • and gpt-oss-120B and Llama-4-Scout-17B-16E-Instruct in 4-bit quantized form (Q4_K_M and Q4_K_XL, respectively) on CPU nodes (32 threads, 128 GB RAM) via llama.cpp [^14] https://github.com/ggml-org/llama.cpp.
All generations used temperature =0.6=0.6 and top-p =0.9=0.9.
Prompt Format.
Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure˜27 for Migrant and in Figure˜30 in the Appendix for Woman).
In few-shot settings, one representative example is included for each class label. GPT-4 receives all instructions in a single prompt (see the prompt text in Figure˜25 for Migrant and Figure˜26 in the Appendix for Woman). Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
Baseline fine-tuning details.
For our baseline, we use a BERT-based pipeline with 110M parameters, comprising a high-level category classifier and two subtype models for solidarity and anti-solidarity. All models share a similar architecture, processing inputs with a target (Frau or Migrant) and full text comprising the focus sentence and left and right context. We add a fully connected layer with softmax activation atop the pooled output of the BERT-based models, with 4 output units for each model. To counter class imbalance, minority classes are oversampled to parity with the majority class. We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer (Loshchilov and Hutter, 2017) and categorical crossentropy loss.
Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 (Reimers and Gurevych, 2019). We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions. The model incorporates ReLU activation and dropout (0.1), and maps the embeddings to four and ten target classes for high-level and fine-grained tasks, respectively. To manage class imbalance, we use oversampling. The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
Appendix D Error Correlation
For two models $m_{i}$ and $m_{j}$, let $\mathcal{E}_{i}$ and $\mathcal{E}_{j}$ denote the sets of instances misclassified by each model, and let $\mathcal{E}^{\text{same}}_{ij}$ be the subset where both models predict the same wrong label. We compute two measures of pairwise error overlap, reported in Table˜4.
Error Jaccard
$J_{ij}=\frac{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}{|\mathcal{E}_{i}\cup\mathcal{E}_{j}|}$
This measures how often two models make errors on the same instances relative to their total error sets.
Same-Wrong Rate
$P_{\text{same-wrong}\mid\text{both-wrong}}=\frac{|\mathcal{E}^{\text{same}}_{ij}|}{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}$
This measures the proportion of jointly misclassified instances on which two models predict the same incorrect label. Table˜4 reports the resulting values for all model pairs.
Model 1 Model 2 Error Jaccard P(Same Wrong $\mid$ Both Wrong)
gpt-oss-120B Qwen-2.5 0.637 0.494
Llama-3.3-70B Qwen-2.5 0.610 0.448
Llama-3.3-70B gpt-oss-120B 0.574 0.456
Table 4: Pairwise overlap in model prediction errors. Error Jaccard measures the proportion of shared errors relative to the union of all errors between two models. P(Same Wrong $\mid$ Both Wrong) measures how often two models predict the same wrong label when both are incorrect.
Appendix E Label Distribution Distance Metrics
We compare model and human label distributions (Section˜3.2.1) using Jensen–Shannon Divergence and total variation distance.
Jensen–Shannon Divergence.
Let $p$ and $q$ denote discrete probability distributions over the same label set, and let $m=\frac{1}{2}(p+q)$. The Jensen–Shannon divergence is defined as
$\mathrm{JSD}(p\,\|\,q)=\frac{1}{2}\mathrm{KL}(p\,\|\,m)+\frac{1}{2}\mathrm{KL}(q\,\|\,m),$
where $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence. Lower values indicate more similar label distributions.
Total Variation Distance.
Total variation distance measures absolute differences in probability mass:
$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{i}|p_{i}-q_{i}|.$
It ranges from 0 (identical distributions) to 1 (disjoint support).
Comparison Total Variation Jensen–Shannon Divergence
Model ensemble vs humans 0.49 (0.32) 0.39 (0.22)
Human vs human (avg 10 splits) 0.45 (0.19) 0.35 (0.12)
Randomized model baseline 0.75 (0.59) 0.69 (0.47)
gpt-oss-120B vs humans 0.58 (0.36) 0.48 (0.29)
Qwen-2.5 vs humans 0.59 (0.42) 0.50 (0.34)
Llama-3.3-70B vs humans 0.57 (0.40) 0.48 (0.32)
GPT-5 vs humans 0.57 (0.33) 0.47 (0.25)
Table 5: Distribution distances between human and model label distributions (lower = closer alignment). Values are shown as fine-grained (high-level).
Label Woman Migrant Total per label
Group-based Solidarity 112 (7.8%) 270 (12.4%) 382 (10.6%)
Exchange-based Solidarity 54 (3.8%) 79 (3.6%) 133 (3.7%)
Empathic Solidarity 125 (8.7%) 32 (1.5%) 157 (4.3%)
Compassionate Solidarity 732 (51.0%) 692 (31.7%) 1424 (39.3%)
Solidarity (no subtype) 41 (2.9%) 55 (2.5%) 96 (2.7%)
Total Solidarity 1064 (74.1%) 1128 (51.6%) 2192 (60.5%)
Group-based Anti-solidarity 10 (0.7%) 288 (13.2%) 298 (8.2%)
Exchange-based Anti-solidarity 0 (0.0%) 55 (2.5%) 55 (1.5%)
Empathic Anti-solidarity 17 (1.2%) 6 (0.3%) 23 (0.6%)
Compassionate Anti-solidarity 7 (0.5%) 97 (4.4%) 104 (2.9%)
Anti-solidarity (no subtype) 5 (0.3%) 19 (0.9%) 24 (0.7%)
Total Anti-solidarity 39 (2.7%) 465 (21.3%) 504 (13.9%)
Mixed 60 (4.2%) 149 (6.8%) 209 (5.8%)
None 273 (19.0%) 442 (20.2%) 715 (19.8%)
Total 1436 (100.0%) 2184 (100.0%) 3620 (100.0%)
Table 6: Distribution of labels in the human-annotated dataset for Woman and Migrant instances. Percentages are relative to each column total.
Figure 18: Macro F1 scores over time for each model (best configuration), and average pairwise Cohen’s Kappa for human annotators. Evaluated on the full test set, including Frau and both test sets for Migrant, grouped by decade.
Figure 19: High-level row-normalized confusion matrices (%) for the best-performing configuration of each model on the two migrant test sets (Test 1 and Test 2, combined). Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
Figure 20: Fine-grained row-normalized confusion matrices (%) showing the best-performing configuration for each of the selected models on the two migrant test sets (Test 1 and Test 2), combined. Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
Table 7: F1 scores for selected models on fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman. The top section contains fine-grained categories, and the bold rows show performance on high-level classification. Scores are chosen for the best performing configuration for the model in terms of Macro F1.
Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
Group-based S 0.42 0.33 0.37 0.39 0.29 0.35 0.31 0.25 0.31
Exchange-based S 0.43 0.60 0.55 0.42 0.39 0.44 0.38 0.43 0.22
Empathic S 0.48 0.05 0.26 0.26 0.11 0.17 0.30 0.11 0.00
Compassionate S 0.60 0.50 0.48 0.50 0.57 0.59 0.42 0.46 0.53
Solidarity (high-level) 0.81 0.80 0.78 0.69 0.71 0.78 0.54 0.58 0.67
Group-based AS 0.47 0.51 0.46 0.39 0.47 0.42 0.31 0.44 0.32
Exchange-based AS 0.43 0.48 0.47 0.40 0.42 0.48 0.30 0.30 0.10
Empathic AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Compassionate AS 0.20 0.00 0.06 0.16 0.08 0.19 0.16 0.21 0.18
Anti-solidarity (high-level) 0.70 0.69 0.66 0.59 0.56 0.63 0.57 0.58 0.39
Mixed 0.32 0.46 0.32 0.24 0.48 0.08 0.26 0.17 0.26
None 0.66 0.71 0.68 0.65 0.66 0.68 0.64 0.59 0.62
Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
Group-based S 0.45 0.28 0.40 0.41 0.25 0.25 0.24 0.20 0.24
Exchange-based S 0.50 0.31 0.40 0.00 0.00 0.00 0.00 0.22 0.35
Empathic S 0.27 0.27 0.34 0.46 0.41 0.31 0.36 0.33 0.22
Compassionate S 0.62 0.43 0.59 0.42 0.51 0.34 0.32 0.37 0.62
Solidarity (high-level) 0.86 0.85 0.82 0.76 0.79 0.76 0.67 0.75 0.77
Group-based AS 0.22 0.31 0.25 0.27 0.22 0.22 0.36 0.27 0.00
Exchange-based AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Empathic AS 0.50 0.67 0.47 0.00 0.25 0.36 0.00 0.00 0.00
Compassionate AS 0.40 0.50 0.50 0.24 0.22 0.30 0.33 0.15 0.00
Anti-solidarity (high-level) 0.59 0.65 0.53 0.30 0.38 0.39 0.44 0.48 0.12
Mixed 0.50 0.30 0.38 0.33 0.37 0.43 0.40 0.32 0.15
None 0.65 0.62 0.48 0.51 0.47 0.26 0.54 0.66 0.21
Figure 21: Distribution of all Migrant keywords over the years, normalized per keyword. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
Figure 22: (Anti-)solidarity trends for the Migrant category by keyword and decade, 1867-2025. Each panel shows DSL-adjusted decade-level shares of solidarity and anti-solidarity for one keyword (see definition (c) in Section˜3.2.2). The mixed and none categories are included in the four-class distribution but omitted from the visualization. Keywords are sorted by frequency, so estimates become less reliable toward the bottom right.
(a) Human majority vote
(b) Human soft labels
(c) gpt-oss-120B
(d) Model majority vote
(e) Llama-3.3-70B
(f) Qwen-2.5-72B
(g) Model soft labels
(h) DSL (soft labels)
Figure 23: High-level solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
(a) Human majority vote
(b) Human soft labels
(c) gpt-oss-120B
(d) Model majority vote
(e) Llama-3.3-70B
(f) Qwen-2.5-72B
(g) Model soft labels
(h) DSL (soft labels)
Figure 24: Fine-grained solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
Label Model Pearson Spearman RMSE
High-level trends (Human majority vote)
Solidarity Llama-3.3-70B 0.907 0.894 0.101
DSL (soft labels) 0.910 0.938 0.075
Anti-solidarity Llama-3.3-70B 0.886 0.847 0.065
DSL (soft labels) 0.928 0.921 0.041
Mixed Llama-3.3-70B 0.810 0.830 0.029
DSL (soft labels) 0.860 0.873 0.027
None Llama-3.3-70B 0.910 0.906 0.134
DSL (soft labels) 0.868 0.899 0.074
Average Llama-3.3-70B 0.878 0.869 0.082
DSL (soft labels) 0.892 0.908 0.054
Fine-grained trends (Human majority vote)
Group-based S Llama-3.3-70B 0.412 0.358 0.202
DSL (soft labels) 0.679 0.678 0.062
Exchange-based S Llama-3.3-70B 0.588 0.458 0.073
DSL (soft labels) 0.461 0.508 0.055
Empathic S Llama-3.3-70B 0.763 0.827 0.044
DSL (soft labels) 0.559 0.519 0.022
Compassionate S Llama-3.3-70B 0.823 0.772 0.204
DSL (soft labels) 0.872 0.912 0.080
Group-based AS Llama-3.3-70B 0.855 0.824 0.087
DSL (soft labels) 0.930 0.908 0.058
Exchange-based AS Llama-3.3-70B 0.706 0.642 0.024
DSL (soft labels) 0.728 0.729 0.021
Empathic AS Llama-3.3-70B 0.497 0.435 0.017
DSL (soft labels) 0.365 0.442 0.018
Compassionate AS Llama-3.3-70B 0.447 0.492 0.064
DSL (soft labels) 0.842 0.660 0.036
Average Llama-3.3-70B 0.636 0.601 0.089
DSL (soft labels) 0.680 0.670 0.044
Table 8: Per-label correlation between model predictions and human-annotated decade-level trends. Higher is better for Pearson and Spearman; lower is better for RMSE.
Figure 25: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Migrant dataset.
Figure 26: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Frau dataset.
(a) High-level prompt (Step 1).
Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B. The model is first given the high-level classification prompt in Figure˜30(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
(a) High-level prompt (Step 1).
Figure 30: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset. The model is first given the high-level classification prompt in Figure˜30(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 31: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 32: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
Figure 33: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the restricted corpus containing only the keyword groups Flüchtlinge and Ausländer. Figure˜33(b) shows the high-level (anti-)solidarity trends, while Figure˜33(a) shows the distribution of fine-grained (anti-)solidarity subtypes over time. In the fine-grained panel, percentages are normalized within each decade across the eight solidarity and anti-solidarity subtypes. The period 1930-1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
Keyword Solidarity Anti-Solidarity Mixed None N (statements)
Asylsuchende 4.11 2.06 0.48 3.35 10
Ausländer 291.43 118.86 21.43 326.28 758
Aussiedler 118.43 12.48 15.22 25.87 172
Einwanderer 2.84 3.21 0.20 19.75 26
Emigranten 21.14 11.08 2.61 10.18 45
Flüchtlinge 1917.08 321.69 174.29 859.94 3273
Heimatvertriebene 3158.67 422.75 167.72 1006.85 4756
Sowjetzonenflüchtlinge 308.99 27.13 2.63 138.25 477
Zuwanderer 58.24 6.54 0.86 60.36 126
Table 9: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per keyword included in the post-war trends analysis Section˜4.2. Values reflect summed DSL-adjusted expected counts. Totals correspond to the number of keyword-coded statements in the analyzed subset.
Party Solidarity Anti-Solidarity Mixed None N (statements)
BP 134.13 54.77 67.51 46.60 303
CDU/CSU 2040.48 229.98 94.93 1050.62 3416
DP 138.21 40.81 -0.54 26.52 205
DRP 34.53 1.47 1.74 3.26 41
FDP 539.79 38.02 56.78 251.41 886
GB/BHE 318.75 27.61 9.03 17.62 373
KPD 151.75 57.43 -1.14 18.96 227
SPD 1595.06 153.56 37.13 478.25 2264
WAV 191.81 19.72 37.12 15.35 264
Zentrum 53.61 3.02 1.35 5.01 63
Table 10: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per party included in the post-war trends analysis Section˜4.2. Values reflect summed DSL-corrected expected counts. Totals correspond to the number of party-coded statements in the analyzed subset.
Original Text Translation
[…]
? […] das Interessante […] ist, daß nicht nur die östliche Landwirtschaft sondern auch die Industrie Ausländer braucht, daß die Zahlen der industrietätigen Ausländer im Lause der Jahre viel mehr gewachsen sind als die Zahlen der von der Landwirtschaft gebrauchten Arbeiter. […] Das Maß des Bedürfnisses der Beschäftigung von Ausländern ist natürlich auch hier örtlich und nach Betrieben verschieden. Es gibt gewisse Betriebe, für die inländische Arbeiter kaum noch zu bekommen sind. […] Es gibt aber auch industrielle Gebiete, die ohne diese Ausländer gar nicht weiter existieren können. ? ? […] what is interesting […] is that not only eastern agriculture but also industry needs foreigners, and that the number of foreigners employed in industry has grown over the years much more than the number of workers needed in agriculture. […] The extent of the need for the employment of foreigners naturally also varies here by locality and by sector. There are certain sectors in which domestic workers can scarcely be found anymore. […] There are also industrial regions that simply could not continue to exist without these foreigners. ?
[…]
?[…] Die Beschleunigung der Asylverfahren ist überfällig, und eine konsequente Rückführung […], und […] eine konsequente Abschiebung. Übrigens sind zügige Verfahren auch und gerade im Interesse der Menschen mit Bleibeperspektive, die wir dann umso besser integrieren können, weil wir ihnen eine Perspektive geben können. Zu guter Letzt müssen wir die Einwanderung nach Deutschland insgesamt neu ordnen, mit einem Einwanderungsgesetz, das diesen Namen verdient. Herr Kollege Castellucci, möglicherweise wird umgekehrt ein Schuh daraus: Wer Akzeptanz für legale Migration erhöhen will, der muss auch entschlossen illegale Migration bekämpfen. […]? ?Speeding up asylum procedures is overdue, as are consistent returns […] and […] consistent deportations. Swift procedures are also, and especially, in the interest of people who have a prospect of staying, because we can then integrate them all the better by giving them a perspective. Finally, we must reorganize immigration to Germany as a whole, with an immigration law worthy of the name. Mr Castellucci, perhaps the reverse is true: anyone who wants to increase acceptance of legal migration must also resolutely combat illegal migration. […]?
[…]
?[…] Sie [die europäischen Staaten] sind aber nicht bereit, Flüchtlinge vor Menschenrechtsverletzungen zu schützen, sondern diffamieren sie als Schmarotzer und Bedrohung unserer Sicherheit, wie dies leider auch gerade durch den Kollegen der SPD wieder geschehen ist. Der Vertrag von Maastricht sieht die gemeinsame Einführung eines Visumzwangs vor, wenn auf Grund einer Notlage in einem Land eine größere Fluchtbewegung zu erwarten ist. So wird ganz bewußt verfolgten oder unter sonstigen Notlagen leidenden Menschen die Flucht erheblich erschwert, oft sogar unmöglich gemacht. Flüchtlinge werden damit geradezu in die Arme von Schlepperorganisationen getrieben, über die man sich hier so gerne beschwert.? ?[…] But they [the European states] are not willing to protect refugees from human rights violations; instead, they defame them as parasites and a threat to our security, as unfortunately just happened again in the remarks of the SPD colleague. The Maastricht Treaty provides for the joint introduction of a visa requirement if, because of an emergency in a country, a larger refugee movement is to be expected. In this way, people who are persecuted or otherwise suffering acute hardship are deliberately making it much harder to flee, often even impossible. Refugees are thereby virtually driven into the arms of smuggling organizations, about which people here so readily complain.?
Table 11: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
Example (4): Predictions: group-based solidarity (gpt-oss, Qwen), group-based anti-solidarity (Llama). Human reference: group-based solidarity (10 Apr 2024)
?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht. Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird. Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […]. Sie ?übersehen? in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben. Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen. Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist. […] ? ?[…] Let us turn to the AfD’s draft and to what makes it simply atrocious. You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved. Not only does this place all foreigners under general suspicion […], but it also violates EU law […]. You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency. But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is. […]?
[…]
? […] Diese Rückführung [der Evakuierten] muß beschleunigt werden, und wir müssen gesetzgeberische Maßnahmen ergreifen, damit die Evakuierten nicht immer wieder auf lange Zeit vertröstet werden, sondern in absehbarer Zeit zu ihren Rechten, zu einer menschenwürdigen Behandlung kommen. […] Gerade das Vertrauen dieser Millionen von Menschen verpflichtet uns, sie auch in den nächsten Jahren nicht zu enttäuschen, […] diesen Eingliederungsprozeß auf allen Gebieten mit dem größten Nachdruck zu fördern […]. In der politischen Diskussion hat zwar das Problem der sachlichen, wirtschaftlichen und sozialen Interessen der Vertriebenen und aller Kriegsgeschädigten kaum noch eine Rolle gespielt […]. ? ? […] This return [of evacuees] must be accelerated, and we must take legislative measures so that evacuees are not kept waiting indefinitely time and again, but instead obtain their rights and humane treatment within a foreseeable period of time. […] It is precisely the trust of these millions of people that obliges us not to disappoint them in the years ahead either, […] and to promote this process of integration in every sphere with the greatest possible determination […]. In political debate, the problem of the material, economic, and social interests of the expellees and all other victims of war has in fact hardly played any role any longer […]. ?
[…]
?Die Bundesrepublik hat […] einen Vorbehalt zur UN-Kinderrechtskonvention zurückgenommen, der ihr zusichern sollte, dass sie ausländische Kinder schlechter behandeln kann als inländische. […] In den vorliegenden Anträgen […] ist dargelegt, was alles zu tun wäre: Heraufsetzung der Verfahrensmündigkeit von minderjährigen Flüchtlingen von 16 auf 18 Jahre, Ausbau und Verbesserung der Betreuung unbegleiteter minderjähriger Flüchtlinge durch die Jugendfürsorge, ihre Herausnahme aus dem Flughafenverfahren, keine Abschiebehaft für Minderjährige, keine Unterbringung von Kindern und Jugendlichen in Sammelunterkünften für Asylsuchende und Flüchtlinge. Unbegleitete Kinder und Jugendliche, deren Eltern nicht ausfindig gemacht werden können, sollen genauso behandelt werden wie elternlose deutsche Kinder. […] ? ?The Federal Republic has […] withdrawn a reservation to the UN Convention on the Rights of the Child that had ensured it could treat foreign children worse than German children. […] The motions submitted […] outline what still needs to be done: raising the age of legal capacity in proceedings for minor refugees from 16 to 18; expanding and improving youth welfare support for unaccompanied minor refugees; removing them from the airport procedure; no detention pending deportation for minors; no accommodation of children and adolescents in collective facilities for asylum seekers and refugees. Unaccompanied children and adolescents whose parents cannot be located should be treated in the same way as German children without parents. […] ?
Table 12: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
[…]
?Das Hamburger Modell ist geradezu vorbildlich; denn hier werden nicht Menschen mit geringem Einkommen und Geflüchtete gegeneinandergestellt, sondern alle haben den gleichen Anspruch auf Mobilität, mit dem Unterschied, dass die Geflüchteten das von ihrem Taschengeld zwangsweise abgezogen bekommen. […] ?Wir sind auf Migration angewiesen, aber die Flüchtlinge…?, könnte man sagen. […] Ich bin so erzogen: Wenn jemand Hilfe braucht und du kannst helfen, dann hilf. Ich bin stolz und froh, dass unser Land helfen kann […].? ?The Hamburg model is virtually exemplary, because here people with low incomes and refugees are not being pitted against one another; rather, everyone has the same entitlement to mobility, with the difference that refugees have it forcibly deducted from their allowance. […] One might say, ?We depend on migration, but the refugees…?. […] I was raised to believe this: if someone needs help and you can help, then help. I am proud and glad that our country is able to help […].?
[…]
?[…] Die große Aufgabe, die uns hier gesetzt war, war der Versuch, einen Kompromiß […] zwischen den Interessen der Ausländer in der Bundesrepublik Deutschland und den berechtigten Ansprüchen der deutschen Bevölkerung auf Aufrechterhaltung der öffentlichen Ordnung und Sicherheit zu finden. Das Gesetz ist ein Ausdruck unserer freiheitlichen demokratischen Grundordnung, indem es den Ausländern in der Bundesrepublik ein sehr hohes Maß von Entfaltungsmöglichkeiten gewährt. […] Das Gesetz nimmt auch besonders Bedacht auf eine Verbesserung der Rechtslage der ausländischen Flüchtlinge […], die ihre Heimat unter dem Druck totalitärer Systeme verlassen mußten. […]? ?[…] The major task before us was to try to find a compromise […] between the interests of foreigners in the Federal Republic of Germany and the legitimate claims of the German population to the maintenance of public order and security. This law is an expression of our liberal democratic constitutional order in that it grants foreigners in the Federal Republic a very high degree of scope for self-development. […] The law also gives special consideration to improving the legal position of foreign refugees […] who were forced to leave their homeland under pressure from totalitarian systems. […]?
[…]
?[…] Mit diesem auf Angebotsorientierung im Migrationsrecht zielenden Gesetzentwurf wollen die Grünen im Kern nichts anderes, als die Einwanderung massiv vereinfachen und die Tore weit aufstoßen. Wir als Unionsfraktion stellen uns gegen die Beliebigkeit im Einwanderungsrecht und stehen daher auch klar für eine gezielte Auswahl bei der Migration ein. Der Staat und die Politik haben die Pflicht, bewusst darauf zu achten, dass die Migration für unser Land vor allem einen Mehrwert und keine Belastung darstellt, und genau darin liegt der Unterschied zum Asyl- und Flüchtlingsrecht. […] ? ?[…] With this bill, which is aimed at a supply-oriented approach to migration law, the Greens essentially want nothing other than to make immigration much easier and throw the gates wide open. As the CDU/CSU parliamentary group, we oppose arbitrariness in immigration law and therefore clearly support a targeted selection in migration. The state and policymakers have a duty to ensure consciously that migration represents above all an added value for our country and not a burden, and that is precisely the difference from asylum and refugee law. […]?
Table 13: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
[…]
?52 Prozent der britischen Muslime sagen nicht nur: ?Homosexualität ist moralisch falsch?, sondern: Homosexualität sollte verboten werden. […] bei fortgesetzter Einwanderung [wird sich] der Anteil der Muslime in Deutschland in den kommenden Jahren auf 20 Prozent erhöhen. In vielen Städten und Gemeinden sind sie dann die Mehrheit. Der Islam kennt keine Minderheitenrechte. Das sind die dunklen Wolken, die über den Homosexuellen in Westeuropa aufziehen.? ?52 percent of British Muslims say not only, ?Homosexuality is morally wrong?, but also: homosexuality should be banned. […] if immigration continues, the share of Muslims in Germany will rise to 20 percent in the coming years. In many cities and municipalities, they will then be the majority. Islam recognizes no minority rights. These are the dark clouds gathering over homosexuals in Western Europe.?
[…]
?Ausländer bleibt unerachtet solchen Aufenthalts innerhalb der Mecklenburgischen Lande immer Ausländer, und einem solchen die Trauung innerhalb der diesseitigen Lande zu gestatten, ist allein das Großherzogliche Ministerium des Innern competent. […] Vielmehr muß in solchen Fällen der Wiederverheirathung der Mecklenburger sein Recht zur Verheiratung durch einen gesetzlichen Domizilschein über das ihm für einen Mecklenburgischen Ort ertheilte Recht der Niederlassung und Verheiratung nachweisen und der Ausländer den Konsens des Großherzoglichen Ministeriums des Innern beibringen, bevor die Trauung geschehen darf. Ja, meine Herren, da sehen Sie nun, was in Mecklenburg das Indigenat aus dem Artikel 2 der Bundesverfassung für eine Bedeutung hat, und wie nothwendig es ist, daß irgend eine Ergänzung des Freizügigkeitsgesetzes […] erfolge, damit das Gesetz dort nicht blos eine Schale ohne Kern bleibe.? ?A foreigner remains a foreigner within the lands of Mecklenburg despite such residence, and only the Grand Ducal Ministry of the Interior is competent to permit such a person to marry within these territories. […] Rather, in such cases of remarriage, the Mecklenburger must prove his right to marry by means of a lawful certificate of domicile showing that he has been granted the right of residence and marriage in a Mecklenburg locality, and the foreigner must produce the consent of the Grand Ducal Ministry of the Interior before the marriage may take place. Gentlemen, there you can see what significance indigeneity under Article 2 of the federal constitution has in Mecklenburg, and how necessary it is that some supplementation of the Freedom of Movement Act […] should be made, so that the law there may not remain merely a shell without a kernel.?
[…]
?Wann und warum ein Ausländer ohne ein vorangegangenes gerichtliches Erkenntniß aus dem Reich oder einem Partikularstaat ausgewiesen werden darf, darüber existiren zwar keine gesetzlichen Normen und Bestimmungen […]. Rechtfertigende Gründe der Ausweisung eines Ausländers nach dem allgemein anerkannten Völkerrechte sind folgende: erstens Gefährdung der öffentlichen Sicherheit durch ein feindliches Auftreten des Ausgewiesenen […]; zweitens […]; drittens[…], und viertens […]. Damit sind alle Fälle erschöpft, in welchen nach dem modernen Völkerrecht Ausweisungen von einzelnen Personen stattfinden dürfen, und alle Lehrer des Völkerrechts sind um so vorsichtiger in der Empfehlung auch nur eines dieser Mittel, als sich das moderne Völkerrecht immer mehr der Theorie zuwendet, daß ein domizilirter Ausländer zu den wirklichen Staatsangehörigen gehört und als solcher zu behandeln ist. […]? ?Although there are no legal norms or provisions governing when and why a foreigner may be expelled from the Empire or from an individual state without prior judicial judgment […], the reasons justifying the expulsion of a foreigner under generally recognized international law are the following: first, endangerment of public security through hostile conduct on the part of the person expelled […]; second […]; third […]; and fourth […]. These exhaust all the cases in which, under modern international law, expulsions of individual persons may take place, and all teachers of international law are all the more cautious in recommending even one of these measures because modern international law is increasingly turning toward the theory that a domiciled foreigner belongs among the true members of the state and is to be treated as such. […]?
Table 14: Error analysis examples. Bold text is the main sentence, the other sentences are for context. The reference label is the final human consensus label. Where relevant, we also report the distribution of individual human annotations.
Original Text Translation
(1) Dominant DSL subtype score: group-based solidarity towards expellees
(May 14, 1952)
?[…] das Lastenausgleichsgesetz und die Eingliederung der Vertriebenen [steht] in keinem Widerspruch zu unserm Anspruch und unserm Recht auf die Heimat. Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden. […] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten […]. ? ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland. On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy. […] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations […]. secondary, dependent only on what is left over. ?
(2) Dominant DSL subtype score: group-based solidarity towards expellees
(Feb 25, 1953)
?[…] Der Ihnen vorliegende Entwurf [zum Bundesvertriebenengesetz] geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen. Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.? ?[…] The draft [on the Federal Expellees Act] before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy. The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
(3) Dominant DSL subtype score: compassionate solidarity towards Soviet-zone refugees
(June 15, 1955)
?[…] es ist notwendig, zu den noch immer vielseitigen Problemen der Sowjetzonenflüchtlinge Stellung zu nehmen […]. Wir erkennen selbstverständlich an, daß durch die Neuregelung des Härtefonds im Lastenausgleich eine Verbesserung für die Sowjetzonenflüchtlinge erfolgte, weil die Ermessensfragen, unter denen dieser Personenkreis bislang besonders zu leiden hatte, dadurch eingeschränkt werden. […] Wenn wir aber die Sicherheit haben wollen, daß sich diese Neuregelung auch wirklich positiv für den Flüchtling auswirkt, so ist es notwendig, daß die vorgesehene Erhöhung des Härtefonds […] wirklich realisiert wird.? ?[…] it is necessary to address the still manifold problems of the Soviet-zone refugees […]. We naturally recognize that the reform of the hardship fund in the Equalization of Burdens has brought improvements […] because the discretionary decisions that this group had particularly suffered from have been curtailed. But if we want to be sure that this reform really has a positive effect for the refugee, then […] the planned increase of the hardship fund […] must actually be implemented.?
(4) Dominant DSL subtype score: group-based anti-solidarity towards expellees
(Dec 13, 1950)
?In unseren Landkreisverwaltungen sind heute 42,2% Heimatvertriebene. […] Auch die einheimische Bevölkerung beklagt sich […] man höre auf den Landratsämtern überwiegend ostpreußisch, und das Plattdeutsche sei zur Seite gedrängt. Unter den Lehrern […] 51% Heimatvertriebene, unter den vollbeschäftigten Richtern […] mehr als 50%, unter der Polizei sogar 61%; zweifellos eine Überfremdung, die Besorgnis erregt! Aber all diese ehrlichen Bemühungen um Eingliederung der Heimatvertriebenen sind angesichts der ungeheuren Zahl vergeblich. Arbeitslosenzahl […] beträgt unter den Heimatvertriebenen fast 60%. Die Lasten sind unerschwinglich.? ?In our district administrations today, 42.2% are expellees. The native population […] complains […] that in the district offices one mostly hears East Prussian, and that Low German has been pushed aside. Among the teachers 51% are expellees, among the full-time judges also more than 50%, among the police even 61% – undoubtedly an over-foreignization that arouses concern! But all these honest efforts to integrate the expellees are in vain in view of their immense numbers. Unemployment […] almost 60% among the expellees […]. The burdens are unbearable.?
Table 15: Examples for postwar period discussed in Section˜4.2. Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(5) Dominant DSL subtype score: group-based/compass. solidarity towards Soviet-zone refugees
(Sept 23, 1954)
?[…] [Sowjetzonenflüchtlinge] meinen, es könnte ihnen sonst bei mangelnder Bereitwilligkeit der Bevölkerung zu neuen Opfern vielleicht nicht einmal mehr das zur Verfügung stehen, was ihnen heute aus dem Härtefonds zur Verfügung steht. […] Wir halten es für unvertretbar, daß […] diese nach wie vor auf den genannten, im Rahmen des Gesamtlastenausgleichs verhältnismäßig geringen Betrag ohne Rechtsanspruch angewiesen sein sollen. [Die], deren Urstatus dem der Heimatvertriebenen in Westdeutschland doch gleich ist, im Lastenausgleich diesen Heimatvertriebenen nicht gleichgestellt sind.? ?[…] [Soviet-zone refugees] fear that, if public willingness for further sacrifices declines, they may no longer receive even the support currently provided through the hardship fund […]. [It is] indefensible that […] they should still be dependent on the aforementioned, comparatively small amount within the overall equalization scheme without any legal entitlement. […] those whose original legal status is equivalent to that of expellees in West Germany are not treated equally under the Equalization of Burdens system.?
(6) Dominant DSL subtype score: group-based anti-solidarity towards foreigners
(June 21, 1956)
?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen. […] die Bundesrepublik […] [hat] Anlaß, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden. […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.? ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements. […] the Federal Republic […] has reason to shield itself against […] stateless foreigners who […] come with the intention of seeking economic support or political asylum here. […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
(7) Dominant DSL subtype score: group-based solidarity towards foreigners
(Oct 18, 1950)
?[…] [HAuslG] […] [wird] den Besatzungsmächten den Entschluß erleichtern, der Bundesrepublik die volle Verantwortung für die heimatlosen Ausländer zu übertragen. Wenn […] das deutsche Volk seine Bereitschaft bekundet, die Eingliederung der heimatlosen Ausländer in die deutsche Betreuung […] zu erleichtern, so gehen wir dabei von der Erwartung aus, daß auch die heimatlosen Ausländer bemüht sein werden, sich in die nun einmal gegebene Lage einzufügen. Sie werden sich dabei bewußt sein müssen, daß auch die Lage der deutschen Bevölkerung schwierig ist und daß viele Zehntausende insbesondere von Heimatvertriebenen noch immer arbeitslos sind und über keinen angemessenen Wohnraum verfügen. […] der heimatlose Ausländer [wird] viele der Härten dieser Nachkriegszeit mit den Deutschen im Gebiete der Bundesrepublik teilen müssen. ? ? […] [HAuslG] […] will make it easier for the occupying powers to decide to transfer full responsibility for stateless foreigners to the Federal Republic. If […] the German people declare their willingness to facilitate […] the integration of stateless foreigners into German administration and care, we proceed on the assumption that the stateless foreigners will also strive to adapt to the situation as it now stands. They must remain aware that the situation of the German population is also difficult and that many tens of thousands, especially expellees, remain unemployed and lack adequate housing. […] stateless foreigners will have to share many of the hardships of this postwar period with the Germans in the territory of the Federal Republic. ?
(8) Dominant DSL subtype score: group-based solidarity towards foreigners
(Oct 18, 1950)
?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige. Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen… gewährt werden. […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […]. Der Gesetzentwurf geht […] über den Konventionsentwurf, dessen Grundsätze Mindestforderungen für die Behandlung heimatloser Ausländer darstellen, hinaus. […]? ?[…] In public welfare too, they are to receive benefits equal to those of German citizens. They are to be granted access to schools of all kinds under the same conditions […]. the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state. The law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners. […]?
Table 16: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(9) Dominant DSL subtype score: exchange-based solidarity towards foreigners
(April 19, 1951)
?[…] Wir begrüßen die Verabschiedung des Gesetzes für heimatlose Ausländer, weil uns damit die Möglichkeit geboten war, dem Ausland zu beweisen, wie wir Angehörige anderer Nationen in unserem eigenen Staate zu behandeln gewillt sind, und weil wir uns […] erhoffen, daß sie […] dazu beitragen wird, die Frage der Internationalisierung unseres Flüchtlingsproblems und des Zuflusses von Mitteln aus dem Ausland zu fördern. […]? ?[…] We welcome the adoption of the law for stateless foreigners, because it gave us the opportunity to demonstrate to foreign countries how we intend to treat nationals of other nations within our own state, and because we […] hope that it will make a contribution to promoting the internationalization of our refugee problem and the inflow of funds from abroad. […]?
(10) Dominant DSL subtype score: group-based solidarity towards foreigners
(June 15, 1955)
?[…] Es ist kein Zweifel, daß die heimatlosen Ausländer natürlich vom Arbeitsministerium und vom Innenministerium […] eine rein materielle Unterstützung bekommen. Aber diese materielle Unterstützung gibt keine Möglichkeit, ihnen kulturell oder sonstwie unter die Arme zu greifen. […] [Es ist sehr schwer] diese Dinge heute den Ländern nahezulegen.? ?[…] There is no doubt that the stateless foreigners, of course, receive […] purely material support from the Ministry of Labor and the Ministry of the Interior. But this material support provides no opportunity to assist them culturally or in other ways. […] [It is very difficult] today to persuade the Länder on these matters. […]?
(11) Dominant DSL subtype score: exchange-based solidarity towards expellees,
Party: FDP
(Oct 20, 1949)
?Die werden […] von früh morgens bis spät abends arbeiten, wenn man ihnen das nötige Gelände, [Baumaterialien oder das Geld] zur Verfügung stellt. So werden sie in Gemeinschaftsarbeit wesentlich mehr leisten, als wir das sonst heute bei diesen Bauten sehen. In diesem Falle wäre das Geld, das wir dafür aufnehmen, praktisch nur als Kredit anzusehen. Wir hätten die Möglichkeit, daß diese Heimatvertriebenen sich ihr Heim selbst bauen, daß es allmählich ihr Eigentum wird. Wenn wir nur normale Mietsätze verrechnen und das Kapital ohne Zinsen oder zu geringen Zinssätzen verzinsen lassen, so wäre in wenigen, spätestens 15 Jahren, dieses Haus […] Eigentum des Betreffenden. ? ?They would […] work from early morning until late at night if they were provided with the necessary land, [building materials or the money]. In cooperative work they would achieve far more than we otherwise see with such construction today. In this case the money we raise would practically only be regarded as a loan. We would have the possibility that these expellees build their homes themselves, gradually becoming their property. If we only charged normal rents and let the capital bear no or low interest, then within at most fifteen years this house […] would become the property of the person concerned.?
(12) Dominant DSL subtype score: group-based/
compassionate solidarity towards expellees,
Party: GB/BHE
(Oct 29, 1953)
?[…] Wir dürfen die Augen auch nicht vor der Tatsache verschließen, daß Millionen von Vertriebenen und Kriegsgeschädigten noch nicht den Anschluß an unsere soziale Entwicklung gefunden haben. Auch hier mußten bisher Zehntausende arbeitswilliger tüchtiger Mitbürger abseitsstehen. Bei unserem Verlangen nach beschleunigter Eingliederung dieser Bevölkerungskreise geht es nicht nur um die Ablösung einer unproduktiven Fürsorge durch aufbauende Mitarbeit. Es geht auch nicht nur um die Erfüllung von ?Rechtsansprüchen?, und ich darf diesen Ausdruck Rechtsansprüche hier noch einmal betonen. ? ?[…] We must not close our eyes to the fact that millions of expellees and war victims have still not been integrated into our social development. Here too, tens of thousands of willing and capable fellow citizens have so far had to stand aside. In our demand for the accelerated integration of these groups, it is not only a matter of replacing unproductive welfare with constructive participation. It is also not only a matter of fulfilling legal claims — and I would like to emphasize this expression ?legal claims? once again.?
Table 17: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(13) Dominant DSL subtype score: compassionate solidarity towards refugees,
Party: WAV
(March 23, 1950)
?Die Aufnahme dieser Flüchtlinge wird in manchen Ländern bewußt erschwert, sabotiert und unmöglich gemacht. […] Seitdem die Spannung zwischen Osten und Westen zunimmt, hat ein neuer Strom von Flüchtlingen eingesetzt. […] Polen und die Tschechei [durchführen] die neue Ausweisungsaktion auf Befehl Moskaus, […] um das Flüchtlingsheer in Westdeutschland zu vermehren und noch mehr Arbeitslose zu schaffen, den Staat zu belasten und seine Wirtschaft zu zerstören. […]? ?The admission of these refugees is being deliberately obstructed, sabotaged, and made impossible in some states. […] Since the tension between East and West has increased, a new stream of refugees has begun. […] Poland and Czechoslovakia are carrying out the new expulsions on orders from Moscow, […] in order to enlarge the refugee army in West Germany, to create even more unemployed, to burden the state, and to destroy its economy. […]?
(14) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
Party: BP
(Oct 20, 1949)
?Die Sudetendeutschen […] und die Südostdeutschen [stehen] dem süddeutschen, also dem bayerischen Kulturkreis am nächsten. […] Die aus dem Sudetenland bzw. aus Südosteuropa kommenden Deutschen [sollen] zunächst in Süddeutschland zusammen[gefasst werden], […] damit die Eingliederung […] schneller und reibungsloser vor sich gehen kann. [Wir] wollen […] nicht Schwierigkeiten schaffen, wo Schwierigkeiten nicht überwunden werden müssen, und nicht Gräben aufreißen, wo solche nicht vorhanden sind. Das hat mit Rassentheorie […] gar nichts zu tun. […] ? ?The Sudeten Germans […] and the Southeast Germans are closest to the South German, that is, the Bavarian cultural sphere. […] The Germans coming from the Sudetenland and from Southeastern Europe [should] initially be concentrated in southern Germany, […] so that integration […] can proceed more quickly and smoothly. […] we do not want to create difficulties where difficulties need not arise, and we do not want to open up rifts where none exist. This has nothing to do with racial theory […]. ?
(15) Dominant DSL subtype score: group-based solidarity towards refugees,
Party: KPD
(Dec 13, 1950)
?Wir glauben, daß das Problem der Beschaffung von Arbeit für [Umsiedler] unter dem kapitalistischen System nicht in der Form gelöst werden kann […]. Wir sind also der Auffassung, daß die Umsiedlung eine Notwendigkeit ist […]. Wenn hier der Sprecher der CDU das schöne, stolze und richtige Wort geprägt hat ?Unser Haus ist deutsch?, wenn er am Schluß gesagt hat ?Das ganze Deutschland soll es sein?, wie kann man dann […] gleichzeitig die Tatsache, daß sich die Vertriebenen, Flüchtlinge und Umsiedler, wie Sie sie bezeichnen, politisch zu Parteien zusammengeschlossen haben, als die Gefahr einer Überfremdung bezeichnen? ? ?We believe that the problem of providing work for [resettlers] cannot be solved under the capitalist system […]. We therefore take the view that resettlement is a necessity […]. If the CDU speaker here has coined the fine, proud, and correct phrase ?Our house is German?, and if he concluded by saying ?It shall be the whole of Germany?, how then can one at the same time describe the fact that the expellees, refugees, and resettlers, as you call them, have politically organized themselves into parties as a danger of being overrun by foreigners??
(16) Dominant DSL subtype score: exchange-based solidarity towards refugees
Party: CDU/CSU
(Jan 31, 1951)
?[…] Es ist Voraussetzung, dass das deutsche Volk für das System des Lastenausgleichs einen Weg findet, damit das, was aus der deutschen Produktion und Wirtschaftskraft herausgenommen werden muss, […] dafür verwendet wird, die Kräfte der Kriegsgeschädigten, der Flüchtlinge und der Heimatvertriebenen wieder in die deutsche Produktion hineinzustellen und diese damit zu stärken. […] Was wir an Eingliederungshilfe […], an Wohnraumhilfe […] und an Förderungshilfe aller Art […] verwenden, ist […] produktive Ausgabe. […] das ist rund die Hälfte des gesamten Aufkommens. ? ?[…] It is a prerequisite that the German people find a way, within the system of burden sharing, to ensure that what must be taken from German production and economic strength is used […] to restore the productive capacity of war victims, refugees, and expellees and to reintegrate them into German production, thereby strengthening it. […] What we allocate for integration assistance […], for housing assistance […], for all kinds of development assistance is […] productive expenditure. […] that is roughly half of the total revenue.?
Table 18: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(17) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
Party: DP
(March 18, 1953)
?Also auch der dritte, vierte und fünfte Sohn eines ostdeutschen Bauern […] kann alle […] Vergünstigungen erhalten. […] Welcher einheimische Bauer kann […] seinen drei oder vier Jungen eine solche Mitgift erarbeiten? Wir schaffen, wenn wir die alte Vorlage wiederherstellen, zweierlei Recht. Das dürfte so viel böses Blut zur Folge haben, dass das auch nicht im Interesse der Vertriebenen liegen kann. Herr Kollege Dr. Kather hat vorhin gesagt, dass 300 000 bäuerliche Familien aus dem Osten geflüchtet sind. Infolge der Formulierung des § 35 ist die Zahl der Bewerber um einen Betrieb weitaus größer geworden. Ich meine, wir sind unehrlich, wenn wir den Leuten Hoffnungen machen, die doch nicht zu erfüllen sind. ? ?Thus even the third, fourth, and fifth son of an East German farmer […] can receive all the benefits. […] Which local farmer […] can earn such a dowry for three or four of his sons? If we restore the old version of the bill, we create two kinds of law. That would cause so much bad blood that it cannot possibly be in the interest of the expellees themselves. […] Dr. Kather said that 300,000 farming families fled from the East. Because of the wording of § 35, the number of applicants for a farm has become far greater. I believe we are being dishonest if we give people hopes that cannot in fact be fulfilled.?
Table 19: Examples for postwar period discussed in Section˜4.2 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(18) Dominant DSL subtype score: compassionate solidarity towards refugees
Party: Die Grünen
(Nov 6, 2014)
?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen. […] Ich will ganz klar sagen: Das Boot ist nicht voll […]. Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […]. ? ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us. […] I want to state very clearly: the boat is not full […]. Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
(19) Dominant DSL subtype score: group-based anti-solidarity towards refugees
(Oct 1, 2015)
?[…] Deutschland [wird] über kurz oder lang überfordert sein. Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […]. allein im September sind zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen. Sie wurden nicht registriert. Sie wurden nicht kontrolliert. […] Damit besteht auch ein Sicherheitsrisiko. ? ?[…] Germany will sooner or later be overburdened. Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […]. In September alone, several tens of thousands of refugees and asylum seekers came to Germany. They were not registered. They were not controlled. […] This poses a major security risk.?
Table 20: Examples for the 2009-2025 period discussed in Section˜4.3. Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(20) Dominant DSL subtype score: exchange-based anti-solidarity towards refugees
Party: AfD
(May 16, 2018)
?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug […] um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind. In Deutschland lebt bereits eine große soziale Unterschicht […]. Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant. […] Deutschland ist kein reiches Land, sondern ein leistungsfähiges. ? ?[…] each year about 140,000 highly qualified people leave this country, while in return, […] about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate. In Germany, there already lives a large social underclass […]. Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people. […] Germany is not a rich country, but a capable one.?
(21) Dominant DSL subtype score: group-based anti-solidarity towards refugees
Party: AfD
(Oct 19, 2022)
?[…] Echte Kriegsflüchtlinge aus der Ukraine brauchen Schutz – keine Frage. Wir sagen aber Nein zu Fluchttourismus durch halb Europa, und wir sagen Nein zu bevorzugter Behandlung auf Kosten der deutschen Steuerzahler. Wir fordern daher die sofortige Abschaltung dieses Sozialstaatsmagneten, […] Wir fordern außerdem die Umstellung von Geldleistungen auf Sachleistungen für Asylbewerber, damit Migrationsanreize gesenkt werden und der Missbrauch effektiv unterbunden wird. […] Und wer sich wie Frau Baerbock in erster Linie der ukrainischen Bevölkerung verpflichtet fühlt, der soll sich in der Ukraine um ein Amt bewerben, aber nicht in Deutschland. ? ?[…] Genuine war refugees from Ukraine need protection — no question. But we say no to flight tourism across half of Europe, and we say no to preferential treatment at the expense of German taxpayers. We therefore demand the immediate shutdown of this welfare magnet, […] We also demand the switch from cash benefits to benefits in kind for asylum seekers, so that incentives for migration are reduced and abuse is effectively curbed. […] And whoever, like Ms. Baerbock, feels primarily obligated to the Ukrainian population should run for office in Ukraine, but not in Germany.?
(22) Dominant DSL subtype score: group-based anti-solidarity towards refugees
Party: AfD
(12 May, 2023)
?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt. […] Dafür werden bereits Zimmer in angeblich christlichen Altenheimen gekündigt, weil angebliche Flüchtlinge lukrativer sind, Messerzuwanderung inklusive. […] Bei diesem Recht auf Wohnungstausch der Linken handelt es sich um eine Mischung aus Neid zulasten alter Leute und Zwang zulasten von Vermietern. Es ist die Fortsetzung einer inländerfeindlichen… ? ?[…] elderly people are expected to leave their large apartments because green-left politics invites millions of guests from Africa and the Middle East while choking off housing construction. […] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration included. […] This ?right to apartment swaps? proposed by the Left is a mixture of envy at the expense of the elderly and coercion at the expense of landlords. It is the continuation of an anti-native [policy]…?
(23) Dominant DSL subtype score: exchange-based solidarity towards immigrants
Party: FDP
(May 25, 2023)
?[…] Menschen, die bei uns arbeiten wollen, sollten wir eine echte Chance geben; denn viele, die an unserer Grenze ankommen, wollen ja gerade arbeiten. Wir arbeiten in der Ampel an der Reform des Fachkräfteeinwanderungsgesetzes. Da geht es um mehr Regulierung, gezielte Einwanderung in unseren Arbeitsmarkt. Menschen aus dem Ausland sollten eine realistische Chance bekommen, zu uns zu kommen und arbeiten zu können. Das Fachkräfteeinwanderungsgesetz wird das jetzt endlich besser machen und mehr Menschen die Chance geben, die unsere Wirtschaft auch dringend braucht. […] ? ?[…] People who want to work here should be given a real chance; because many who arrive at our border actually want to work. Within the coalition we are working on reforming the Skilled Immigration Act. This is about more regulation and targeted immigration into our labor market. People from abroad should have a realistic chance to come to us and be able to work. The Skilled Immigration Act will now finally improve this and give more people the opportunity that our economy urgently needs. […]?
Table 21: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(24) Dominant DSL subtype score: group-based anti-solidarity towards migrants
Party: CDU/CSU
(May 14, 2025)
?[…] Die gute Nachricht ist: Der Politikwechsel hat begonnen. […] Die illegale Migration der letzten zehn Jahre gefährdet die politische Stabilität Deutschlands und Europas. […] Unsere Städte und Gemeinden, die Landkreise: Alle sind über dem Limit. In den Kitas, in den Schulen, auf dem Wohnungsmarkt, bei der Sicherheit an den Bahnhöfen und auf den Marktplätzen wird diese Überforderung sehr konkret. Deswegen ist es gut, dass wir gemeinsam im Koalitionsvertrag die Migrationswende vereinbart haben: mehr Rückführungen, schnellere Verfahren und, ja, ebenso die Zurückweisung an den deutschen Grenzen auch bei Asylgesuchen. ? ?[…] The good news is: the policy shift has begun. […] The illegal migration of the last ten years endangers the political stability of Germany and Europe. […] Our towns and municipalities, the districts—everyone is over the limit. In day-care centers, in schools, in the housing market, in security at train stations and in marketplaces, this overburdening becomes very concrete. That is why it is good that we have agreed together in the coalition agreement on a migration turnaround: more deportations, faster procedures, and, yes, also rejections at the German borders, even in the case of asylum applications.?
(25) Dominant DSL subtype score: exchange-based anti-solidarity towards immigrants
Party: FDP
(Jan 30, 2025)
?Das Grundrecht auf Asyl darf nicht dazu führen, dass allein ein Antrag auf Asyl zu mehreren Jahren freiem und staatlich subventioniertem Aufenthalt in Deutschland führt, wenn er völlig ohne Grundlage erfolgt. Solche Missstände müssen wir beseitigen, damit wir uns auf jene konzentrieren können, die unsere Hilfe wirklich brauchen […]. SPD und Grüne […] wollen die illegale Einwanderung nach Deutschland kaum begrenzen und fordern lieber weiteren Familiennachzug. Diesen falschen Weg ist Deutschland leider viel zu lange gegangen. […] Eine immer linkere Politik wird […] dazu führen, dass die AfD ihr teuflisches Werk immer weiter fortsetzen kann. […] ? ?The fundamental right to asylum must not result in a situation where merely submitting an asylum application leads to several years of free and state-subsidized residence in Germany when it is entirely unfounded. Such abuses must be eliminated so that we can focus on those who truly need our help […]. The SPD and the Greens […] hardly want to limit illegal immigration to Germany and instead prefer to demand further family reunification. Unfortunately, Germany has followed this wrong path for far too long. […] Ever more left-leaning policies […] will only enable the AfD to continue its diabolical work […].?
(26) Dominant DSL subtype score: group-based anti-solidarity towards migrants
Party: SPD
(12 Sept, 2024)
?Das Sicherheitspaket, das heute hier zur Debatte steht, ist der erste Schritt, um Konsequenzen aus der furchtbaren Katastrophe in Solingen zu ziehen. Waffenverbotszonen […] helfen [aber] Waffenverbotszonen nicht. Deshalb brauchen wir mehr Befugnisse für das Bundeskriminalamt und das Bundesamt für Migration. Sie müssen in die Lage versetzt werden, mit den öffentlich zugänglichen Informationen Straftaten zu verhindern. Das ist dringend erforderlich, und deshalb werden wir das jetzt umsetzen. Wir werden weitere Schritte veranlassen, damit ausreisepflichtige Personen unser Land verlassen. ? ?The security package being debated here today is the first step toward drawing consequences from the terrible catastrophe in Solingen. Weapons prohibition zones […] do not, however, help against terrorists. That is why we need more powers for the Federal Criminal Police Office and the Federal Office for Migration. They must be put in a position to use publicly available information to prevent crimes. That is urgently necessary, and that is why we will now implement it. We will take further steps to ensure that persons obliged to leave the country actually leave our country.?
Table 22: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Original Text Translation
(27) Dominant DSL subtype score: exchange-based solidarity towards immigrants
Party: Die Grünen
(Jan 19, 2024)
?Wissenschaftliche Institute sprechen von 400 000 Menschen jedes Jahr. Darauf müssen wir die Gesellschaft und die Politik vorbereiten. Wir müssen also ohnehin die entsprechende Infrastruktur ausbauen – Wohnungen, Schulen, Kinderbetreuung. Wir brauchen gezielte Zuwanderung, und dafür haben wir das Fachkräfteeinwanderungsgesetz gemacht. Aber wir wären doch mit dem Klammerbeutel gepudert, wenn wir nicht auch das Potenzial der Menschen, die aus humanitären Gründen zu uns kommen, nutzten und dafür sorgten, dass sie aus- und weitergebildet werden und Teil unserer Gesellschaft werden. Das wäre die Hauptaufgabe. Und weil wir Zuwanderung brauchen, brauchen wir eine Willkommenskultur. ? ?Scientific institutes speak of 400,000 people every year. Society and politics must be prepared for that. We therefore have to expand the relevant infrastructure anyway — housing, schools, childcare. We need targeted immigration, and for that we created the Skilled Workers Immigration Act. But we would have to be out of our minds not to also make use of the potential of people who come to us for humanitarian reasons, and ensure that they receive education and training and become part of our society. That would be the main task. And because we need immigration, we need a culture of welcome.?
(28) Dominant DSL subtype score: compassionate solidarity towards migrants
Party: Die Linke
(Jan 30, 2025)
?Sie sagen, Sie wollen Frauen und Kinder schützen. Sie wollen Kinder und Frauen in Kriegsgebiete abschieben. So ist es. Wir führen eine Debatte über Sicherheit und Migration. Und leider erklären alle Parteien außer der Linken, dass sie mit Abschottung dieses Land sicherer machen wollen, und reden von ?Abschiebungen im großen Stil?, als seien Migranten grundsätzlich eine Gefahr. So werden Menschen unter Generalverdacht gestellt. Weniger Migration mache das Land sicherer, das ist doch Unsinn. ? ?You say you want to protect women and children. You want to deport children and women to war zones. That is the reality. We are having a debate about security and migration. And unfortunately, all parties except the Left declare that they want to make this country safer through isolation, and talk about ?deportations on a large scale,? as if migrants were inherently a danger. In this way, people are placed under general suspicion. Less migration makes the country safer — that is nonsense.?
(29) Dominant DSL subtype score: compassionate anti-solidarity towards asylum seekers
Party: CDU/CSU
(19 Feb 2016)
?Das sind aber nicht nur Flüchtlinge, die Anspruch auf Schutz haben. […] bei mehr als der Hälfte [konnte] keine Schutzbedürftigkeit festgestellt werden. Wenn wir bei Asylbewerbern, bei denen schon bei der Ankunft absehbar ist, dass sie keine Bleibeaussicht haben, nicht in kürzester Zeit entscheiden und sie in ihre Herkunftsländer zurückführen, werden wir noch mehr Anreize schaffen, ohne berechtigten Grund nach Deutschland zu kommen. […] Das ist nicht nur ein klares Signal im Sinne von ?Wir nehmen nur wirklich Schutzbedürftige auf?, sondern entlastet auch deutlich unsere Kommunen bei der Unterbringung. Man kann dabei auch nicht von einer Einschränkung oder Aushöhlung des Asylrechts sprechen. ? ?These are not only refugees who are entitled to protection. […] in more than half of the cases no need for protection could be established. If, in the case of asylum seekers for whom it is already apparent upon arrival that they have no prospect of remaining, we do not make a decision as quickly as possible and return them to their countries of origin, we will create even more incentives to come to Germany without justified reason. […] This is not only a clear signal in the sense of ?We only take in those who genuinely need protection,? but it also significantly relieves our municipalities with regard to accommodation. Nor can one speak of a restriction or hollowing-out of the right to asylum in this context.?
Table 23: Examples for the 2009-2025 period discussed in Section˜4.3 (continued). Bold text is the main sentence, the other sentences are for context. The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
Appendix F Annotation Process Details
Annotators first assigned a high-level stance label and, where applicable, an (anti-)solidarity subtype. To document their decisions, they also identified the specific resource at issue (e.g.  time, money, rights, or access to education), selected an indicator associated with the chosen category, highlighted supporting text spans, and provided a brief free-text explanation (1-2 sentences). Where relevant, annotators also marked expressions of the speaker’s own position and references to others’ positions. These additional annotations supported interpretation and documentation during the annotation process, but were not used in the analyses reported in this paper. The full annotation guidelines are provided in Appendix˜G in the Appendix, and Figure˜34 shows an example from the annotation file.
(a) Columns for high-level and (anti-)solidarity categorizations.
(b) Columns for providing explanations.
Figure 34: Example of the annotation process from the annotation file. Figure˜34(a) illustrates the step where annotators choose a high-level label and an (anti-)solidarity subcategory, if applicable. Figure˜34(b) shows columns for detailed explanations, including the choice of a resource, an indicator, and providing free-text commentary
Appendix G Annotation Guidelines Excerpt
This appendix reproduces the annotation guidelines used by the annotators, with minor formatting edits.
General Information
Goal of the study and annotation
The purpose of this study is to examine expressions of solidarity and anti-solidarity towards two specific vulnerable groups—women and migrants—within German parliamentary proceedings. The definitions of (anti-)solidarity are derived from the work Ils et al. (2021), and Eger et al. (2022). We distinguish four subtypes of solidarity and anti-solidarity based on Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of vulnerable groups), exchange-based (contribution, reciprocity, or burden), and empathic (respect for or rejection of diversity).
Details of the Annotation Process
Annotation instructions
In the next section (Annotation Categories), we provide general theoretical information about the categories used in the annotation. This material is intended to support a better understanding of the concepts underlying the task. It is critical that annotators provide a clear explanation for the chosen category or categories in the Explanation column, while keeping comments short. The Other comments column can be used for any additional observations.
Contextual evaluation
Pay close attention to the middle text, which is the sentence whose stance is to be assessed. The surrounding sentences (previous and next) are provided to support contextual interpretation.
Highlighting sentences
When labeling the instances, annotators highlight parts of or full sentences that guided the label choice. Green is used for solidarity categories and red for anti-solidarity categories. In cases of mixed stance, solidarity-expressing parts are highlighted in green and anti-solidarity-expressing parts in red.
Pay attention to the target group
If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None. For example, if the target group is women, then a sentence such as ?all refugees must now leave this country? should also be marked as None.
Structure of the Annotation File
The annotation file contains the following fields: ID: ID of the sentence. The first four numbers after the word debates indicate the year. Category: Indicates the target group (woman or migrant). Pay close attention to the target group. If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None. Sentences: The text to annotate, consisting of Previous, Middle, and Next. The surrounding sentences are provided for contextual interpretation. However, the central focus of the decision should be the middle text. Main category: One of Solidarity, Anti-solidarity, Mixed stance, or None. (Anti-)solidarity subcategory: One of the subcategories, selected only for instances labeled as Solidarity or Anti-solidarity. Explanation: Resource: The resource on the basis of which solidarity or anti-solidarity is expressed (e.g., access to Germany, rights and well-being). Indicators: An indicator motivating the selected anti-solidarity or solidarity type. Indicators are given in drop-down boxes for the respective (anti-)solidarity type and are also listed in the Annotation Categories section. Free text: A short explanation for the chosen label (1–2 sentences). Annotators should provide a clear explanation for the selected category. Other comments: Any additional observations.
Annotation Categories
Definition of solidarity (based on Ils et al. (2021)): solidarity is the readiness to share one’s own resources with others (be it time, financial means, rights, access to education etc.). 1. Solidarity Group-based solidarity Definition. Group-based solidarity stems from a perceived similarity among members of the same group, leading to a shared identity. The group-based solidarity frames often refer to a certain (desired) commonality and a sense of togetherness due to common interests and goals, shared values and norms, or common rights and duties. Characteristics. To identify group-based solidarity, look for: 1. References to a shared identity, common interests, goals, values, or norms within a group (e.g. It’s our collective responsibility to…; we stand together in the fight…; We must unite to…). 2. The existence of mutual obligations or shared rights and duties. 3. Indication that the group must be integrated or even assimilated. 4. Possibly, the identification of an outgroup that contrasts with the in-group. Emphasis on unity and shared goals/duties/rights among the whole group. Compassionate solidarity Definition. Compassionate solidarity is primarily about recognising and supporting individuals or subgroups who are marginalised, disadvantaged, or vulnerable. It extends assistance to these individuals or subgroups, emphasising their need for protection or help. Characteristics. To identify compassionate solidarity, look for: 1. Acknowledgement of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g. We must pay special attention to the needs…; We need to recognize the difficulties…). 2. Emphasis on the need for their protection, assistance, or special consideration due to their marginalised or disadvantaged status (e.g. It is our duty to support…). Emphasis on extra care, support, and protection for those most vulnerable (both in-group and out-group). Exchange-based solidarity Definition. Exchange-based solidarity is linked to the principle of the role of exchange between partners with complementary qualities. Exchange-based solidarity is coded when a speaker refers to the usefulness of ‘exchange partners’ in terms of their actual or future contributions or willingness to contribute. These exchange partners can be rewarded or stimulated, and may also be required to contribute more to receive support. Characteristics. To identify exchange-based solidarity, look for references to a group: 1. That was actively doing, is doing and / or will be doing something useful. 2. Should either be rewarded for past, present or future behavior or that it should be supported to continue that useful behavior. 3. Does not receive enough and needs more support for its contribution (do not confuse it with compassionate solidarity). Empathic solidarity Definition. Empathic solidarity is a concept grounded in the intersubjective valuation of differences among individuals, recognising and respecting their unique characteristics. In contrast to compassionate solidarity, empathic solidarity is coded when a difference is not something that is bad or something that should disappear but is rather a strength or something positive. Empathic solidarity is coded when a speaker acknowledges diversity, difference, or unique characteristics as aspects that should be respected and taken into account. Characteristics. To identify empathic solidarity, look for: 1. Acknowledgement of a group’s diversity and that this diversity must be respected; 2. Recognition of the group’s unique and distinct characteristics; 3. Affirmation of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices. 4. Advocacy for the group’s freedom to develop themselves the way they want. 2. Anti-solidarity Group-based anti-solidarity Definition. The group-based anti-solidarity often points to an out-group characterised by pronounced differences (concerning interests and goals, values and norms, individual rights and responsibilities), and an expression of exclusion and lack of support for that group based on these differences. Characteristics. To identify group-based anti-solidarity, look for: 1. References to differing interests, goals, responsibilities, values, or norms between the groups and at the same time not willing to provide support based on the differences, sense of belonging to different groups. 2. A stronger emphasis on one’s own group’s interests (e.g. It’s every person for themselves…; We must stand up for our own interests…; We must not sacrifice our personal goals for…). 3. Resistance to assimilation or integration with other groups. Compassionate anti-solidarity Definition. Compassionate anti-solidarity is about disregarding and excluding individuals or subgroups either on the basis that they are perceived to be in a “good position” already, or are viewed as undeserving of help. It neglects these individuals or subgroups, downplaying their need for protection or help. Characteristics. To identify compassionate anti-solidarity, look for: 1. Denial of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g. We shouldn’t focus on the problems of…; We don’t need to pay attention to the struggles of…). 2. Minimising the need for their protection, assistance, or special consideration (e.g. It is not our responsibility to aid…) as they are already in a ?good position? or are in a ?bad position?, but still do not deserve help. Exchange-based anti-solidarity Definition. Exchange-based anti-solidarity focuses on the idea that some groups take more than they give. It brings attention to groups that seem to receive a lot of resources or support, but do not contribute much or cause problems. This perspective supports punishing such groups, giving them less help, or making them contribute more. Characteristics. To identify exchange-based anti-solidarity, look for references to a group: 1. Talk about a group that has received a lot, but has not given much back or has caused harm. 2. Suggestions that this group should be punished, or get less than they are currently getting. 3. Calls for the group to give or do more in return for help. Empathic anti-solidarity Definition. Empathic anti-solidarity refers to the cases when a speaker refers to a group indicating that this group is different from everyone else, but this difference (needs, characteristics, beliefs, opinions) should not be respected or recognised. The speaker could argue that this difference should not be acknowledged. The core logic behind this is: ?You’re not like me, and that’s a shame. I don’t understand or respect you or your differences?. Characteristics. To identify empathic anti-solidarity, look for: 1. Disregard for a group’s diversity, arguing that this diversity should not be respected. 2. Ignorance or dismissal of the group’s unique and distinct characteristics. 3. Denial of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices. 3. Mixed stance Definition. Mixed stance refers to instances where the speaker expresses both support and opposition towards the target group within the same discourse or text. It involves contradictions or conflicting statements regarding the speaker’s stance towards the target group. Characteristics. 1. Annotate instances where the speaker presents a mixed stance by expressing both support and opposition towards the target group in the same discourse or text. 2. Look for contradictions or inconsistencies in the level of support or opposition expressed towards the target group. 3. Note that the middle sentence has more weight. 4. None Definition. None refers to instances where the speaker neither expresses explicit support nor opposition towards the target group. The statement is unrelated to either solidarity or anti-solidarity or is unrelated to the group altogether. Characteristics. 1. Annotate instances where the speaker maintains a neutral position and does not express clear support or opposition towards the target group. 2. Look for statements that provide factual information, or refrain from taking a stance or expressing personal opinion. 3. Also label the instance as None if you think that it is unrelated to the target group.